### v1

# Qingdao 266237, China

Abstract
Multi-party private set union (MPSU) protocol enables $m$ $(m>2)$ parties, each holding a set, to collectively compute the union of their sets without revealing any additional information to other parties. There are two main categories of MPSU protocols: The first builds on public-key techniques. All existing works in this category involve a super-linear number of public-key operations, resulting in poor practical efficiency. The second builds on oblivious transfer and symmetric-key techniques. The only existing work in this category is proposed by Liu and Gao (ASIACRYPT 2023), which features the best concrete performance among all existing protocols, despite its super-linear computation and communication. Unfortunately, it does not achieve the standard semi-honest security, as it inherently relies on a non-collusion assumption, which is unlikely to hold in practice. Therefore, the problem of constructing a practical MPSU protocol based on oblivious transfer and symmetric-key techniques in standard semi-honest model remains open. Furthermore, there is no MPSU protocol achieving both linear computation and linear communication complexity, which leaves another unresolved problem. In this work, we resolve these two open problems.
  • We propose the first MPSU protocol based on oblivious transfer and symmetric-key techniques in the standard semi-honest model. This protocol is 4.99.3×4.9-9.3$\times$4.9 - 9.3 × faster than Liu and Gao in the LAN setting. Concretely, our protocol requires only $3.6$ seconds in online phase for 3 parties with sets of $2^{20}$ items each.
  • We propose the first MPSU protocol achieving both linear computation and linear communication complexity, based on public-key operations. This protocol has the lowest overall communication costs and shows a factor of 3.036.5×3.0-36.5$\times$3.0 - 36.5 × improvement in terms of overall communication compared to Liu and Gao.
We implement our protocols and conduct an extensive experiment to compare the performance of our protocols and the state-of-the-art. To the best of our knowledge, our implementation is the first correct and secure implementation of MPSU that reports on large-size experiments.
1 Introduction
Over the last decade, there has been growing interest in private set operation (PSO), which consists of private set intersection (PSI), private set union (PSU), and private computing on set intersection (PCSI), etc. Among these functionalities, PSI, especially two-party PSI [PSZ14, KKRT16, PRTY19, CM20, PRTY20, RS21, RR22], has made tremendous progress and become highly practical with extremely fast and cryptographically secure implementations. Meanwhile, multi-party PSI [KMP+17, NTY21, CDG+21, BNOP22] is also well-studied. In contrast, the advancement of PSU has been sluggish until recently, several works proposed efficient two-party PSU protocols [KRTW19, GMR+21, JSZ+22, ZCL+23, CZZ+24a]. However, multi-party PSU has still not been extensively studied. In this work, we focus on PSU in the multi-party setting.
Multi-party private set union (MPSU) enables $m$ $(m>2)$ mutually untrusted parties, each holding a private set of elements, to compute the union of their sets without revealing any additional information. MPSU and its variants have numerous applications, such as information security risk assessment [LV04], IP blacklist and vulnerability data aggregation [HLS+16], joint graph computation [BS05], distributed network monitoring [KS05], building block for private DB supporting full join [KRTW19], private ID [GMR+21] etc.
According to the underlying techniques, existing MPSU protocols can be mainly divided into two categories: The first category, denoted PK-MPSU, is primarily based on public-key techniques, and has been explored in a series of works [KS05, Fri07, VCE22, GNT23]. The drawbacks of these works are that each party has to perform a substantial number of public-key operations, leading to super-linear computation complexity and poor practical efficiency. The second category, denoted SK-MPSU, is primarily based on symmetric-key techniques, and has only one existing work [LG23] to date. This work exhibits much better performance than all prior works. However, it fails to achieve the standard semi-honest security due to its inherent reliance on a non-collusion assumption, assuming the party who obtains the union (we call it leader hereafter) not to collude with other parties. Furthermore, it has super-linear complexity as well. Motivated by the above, we raise the following two questions:
Can we construct a MPSU protocol based on oblivious transfer and symmetric-key operations, without any non-collusion assumptions? Can we construct a MPSU protocol with both linear computation and linear communication complexity [^1] In the context of MPSU, linear complexity means that the complexity for each party scales linearly with the total size of all parties’ sets. In this paper, we consider the balanced setting where each party holds sets of equal size, thus linear complexity means that the complexity for each party scales linearly with both the number of parties $m$ and the set size $n$. Meanwhile, following current conventions in the area of PSO, linear complexity only considers the online phase.?
1.1 Our Contribution
In this work, we resolve the above two open problems by first presenting a new primitive for MPSU, called batch secret-shared private membership test (batch ssPMT), then employing the batch ssPMT to build a SK-MPSU in the standard semi-honest model, and a PK-MPSU with linear computation and communication complexity. Our contributions are summarized as follows:
1.1.1 Efficient Batch ssPMT.
At the technical core of the state-of-the-art MPSU protocol [LG23] (hereafter referred to as LG protocol) is the multi-query secret-shared private membership test (mq-ssPMT), which dominates both computation and communication costs of LG protocol. In analogy of the relation between batched oblivious pseudorandom function (batch OPRF) [KKRT16] and multi-point oblivious pseudorandom function (multi-point OPRF) [PRTY19, CM20], we abstract a new functionality called batch secret-shared private membership test (batch ssPMT), which allows a fairly efficient construction and can be used to build an alternative to mq-ssPMT in the context of MPSU. Looking ahead, our batch ssPMT serves as a core building block in our two MPSU protocols, and significantly contributes to our speedup compared to LG protocol.
1.1.2 SK-MPSU in Standard Semi-Honest Model.
We generalize random OT (ROT) into multi-party setting, which we call multi-party secret-shared random oblivious transfer (mss-ROT). Based on batch ssPMT and mss-ROT, we propose the first SK-MPSU protocol in the standard semi-honest model. Compared to LG protocol, our SK-MPSU has superior online performance with a 4.99.3×4.9-9.3$\times$4.9 - 9.3 × improvement in the LAN setting.
1.1.3 PK-MPSU with Linear Complexity.
Based on batch ssPMT and multi-key rerandomizable public-key encryption (MKR-PKE) [GNT23], we propose the first MPSU protocol with both linear computation and communication. Our PK-MPSU has the lowest overall communication costs with a factor of 3.036.5×3.0-36.5$\times$3.0 - 36.5 × improvement compared to LG protocol. It also achieves a 1.85.4×1.8-5.4$\times$1.8 - 5.4 × speedup in terms of overall running time in the WAN setting. Along the way, we find that the PK-MPSU protocol of Gao et al. [GNT23] is insecure against arbitrary collusion and give a practical attack to demonstrate that it necessitates the same non-collusion assumption as LG protocol as well.
Figure 1 depicts the technical overview of our new MPSU framework. We will elaborate the details in Section 2.
Our SK-MPSUOur PK-MPSUbatch ssPMTbatch OPPRFssPEQTmss-ROTmulti-party secret-sharedshuffleMKR-PKE
Figure 1: Technical overview of our MPSU framework. The newly introduced primitives are marked with solid boxes. The existing primitives are marked with dashed boxes.
1.2 Related Works
We review the existing semi-honest MPSU protocols in the literature.
1.2.1 PK-MPSU.
Kisser and Song [KS05] introduced the first MPSU protocol, based on polynomial representations and additively homomorphic encryption (AHE). This protocol requires a substantial number of AHE operations and high-degree polynomial calculations, so it is completely impractical.
Frikken [Fri07] improved [KS05] by decreasing the polynomial degree. However, the number of AHE operations remains quadratic in the set size due to the necessity of performing multi-point evaluations on the encrypted polynomials.
Vos et al. [VCE22] proposed a MPSU protocol based on the bit-vector representations. The parties collectively compute the union by performing the private OR operations on the bit-vectors, realized by ElGamal encryption. The original version of this protocol is merely applicable for small universes. Even after leveraging divide-and-conquer approach, the protocol requires quadratic computation and communication complexity in the number of parties for the leader and shows poor concrete efficiency reported by [LG23].
Recently, Gao et al. [GNT23] proposed a MPSU protocol based on three newly introduced cryptographic tools, including membership Oblivious Transfer (mOT), conditional Oblivious Pseudorandom Function (cOPRF) and MKR-PKE. This protocol achieves near-linear complexity in the set size and linear complexity in the number of parties, and is the most advanced MPSU in terms of theoretical complexity. Unfortunately, their protocol turns out to be insecure against arbitrary collusion. We propose a practical attack to show that it requires the same non-collusion assumption as LG protocol (see Appendix 0.A for details).
1.2.2 SK-MPSU.
Recently, Liu and Gao [LG23] proposed a practical MPSU protocol based on oblivious transfer and symmetric-key operations. This protocol is several orders of magnitude faster than the prior works. For instance, when computing on datasets of $2^{10}$ element, it is 109×109$\times$109 × faster than [VCE22]. However, their protocol is not secure in the standard semi-honest model.
1.2.3 Other Related Works.
Blanton et al. [BA12] proposed a MPSU protocol based on oblivious sorting and generic MPC. Their work focuses on devising the circuit for MPSU. The heavy dependency on general MPC leads to inefficiency.
Table 1 provides a comprehensive theoretical comparison between existing MPSU protocols and our proposed protocols. Leader denotes the participant who obtains the union result. Client refers to the remaining participants.
Protocol Computation Communication Round Security
Leader Client Leader Client
 [KS05] $m^{2}n^{3}$ pub $m^{2}n^{3}$ pub $\lambda m^{3}n^{2}$ $\lambda m^{3}n^{2}$ $m$
 [Fri07] $mn^{2}$ pub $mn^{2}$ pub $\lambda mn$ $\lambda mn$ $m$
 [VCE22] $lm^{2}n$ pub $lmn$ pub $\lambda lm^{2}n$ $\lambda lmn$ $l$
 [BA12] $\sigma mn\log n+m^{2}$ sym $\sigma^{2}mn\log n+\sigma m^{2}$ $\log m$
[GNT23] $mn(\log n/\log\log n)$ pub $(\gamma+\lambda)mn(\log n/\log\log n)$ $\log\gamma+m$
[LG23] $(T+l+m)mn$ sym $(T+l)mn$ sym $(T+l)mn+lm^{2}n$ $(T+l)mn$ $\log(l-\log n)+m$
Our SK-MPSU $m^{2}n$ sym $m^{2}n$ sym $\gamma mn+lm^{2}n$ $(\gamma+l+m)mn$ $\log\gamma+m$
Our PK-MPSU $mn$ pub $mn$ pub $(\gamma+\lambda)mn$ $(\gamma+\lambda)mn$ $\log\gamma+m$
Table 1: Asymptotic communication (bits) and computation costs of MPSU protocols in the semi-honest setting. For the sake of comparison, we omit the Big $O$ notations and simplify the complexity by retaining only the dominant terms. We use to denote protocols in the standard semi-honest model and to denote protocols requiring non-collusion assumption. pub: public-key operations; sym: symmetric cryptographic operations. We ignore the offline phase cost in all SK-MPSU protocols and the symmetric-key operations in all PK-MPSU protocols. [KS05] and [Fri07] use Paillier while [VCE22], [GNT23] and our PK-MPSU use ElGamal. $n$ is the set size. $m$ is the number of participants. $\lambda$ and $\sigma$ are computational and statistical security parameter respectively. $T$ is the number of AND gate in a SKE decryption circuit in [LG23]. $l$ is the bit length of input elements. $\gamma$ is the output length of OPPRF. In the typical setting, $n\leq 2^{24}$, $m\leq 32$, $\lambda=128$, $\sigma=40$, $T\approx 600$, $l\leq 128$ , $\gamma\leq 64$.
2 Technical Overview
2.1 LG Protocol Revisit
We start by recalling the high-level idea of LG protocol. For the sake of simplicity, we focus here on the case of three parties $P_{1},P_{2},P_{3}$, whose inputs are $X_{1},X_{2},X_{3}$ respectively. We designate $P_{1}$ as the leader. Since $P_{1}$ already holds $X_{1}$, it needs to obtain the set difference $Y_{1}=(X_{1}\cup X_{2}\cup X_{3})\setminus X_{1}$ from $P_{2}$ and $P_{3}$.
Roughly speaking, their protocol enables $P_{2}$ to somehow secret-share $Y_{2}=X_{2}\setminus X_{1}$ and $P_{3}$ to somehow secret-share $Y_{3}=X_{3}\setminus(X_{1}\cup X_{2})$ among all parties. Since $\{Y_{2},Y_{3}\}$ is a partition of $Y_{1}$, each party holds a share of $Y_{1}$ in the order of $Y_{2},Y_{3}$ eventually, and any two parties cannot collude to obtain information of the last one’s inputs.
A naive approach to reconstructing $Y_{1}$ to $P_{1}$ is to let $P_{2}$ and $P_{3}$ send their shares to $P_{1}$ straightforwardly. However, in this way, $P_{1}$ can determine the party that each obtained element $x\in Y_{1}$ belongs. Their solution is to let the parties invoke multi-party secret-shared shuffle to randomly permute and re-share all shares. Multi-party secret-shared shuffle guarantees that any two parties have no knowledge of the permutation and all shares are refreshed after the invocation, hence the adversary is unaware of the correspondence between shares and the individual difference sets $Y_{2},Y_{3}$. Then $P_{2}$ and $P_{3}$ can send their shuffled shares to $P_{1}$, who reconstructs $Y_{1}$ and obtains the union by appending elements of $X_{1}$.
LG protocol utilizes two main ingredients: (1) The secret-shared private membership test (ssPMT) [CO18, ZMS+21], where the sender $\mathcal{S}$ inputs a set $X$, and the receiver $\mathcal{R}$ inputs an element $y$. If $y\in X$, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of 1, otherwise they receive secret shares of 0. Liu and Gao proposed a multi-query ssPMT (mq-ssPMT), which supports the receiver querying multiple elements’ memberships of the sender’s set simultaneously. Namely, $\mathcal{S}$ inputs $X$, and $\mathcal{R}$ inputs $y_{1},\cdots,y_{n}$. $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $n$, where if $y_{i}\in X$, the $i$th bit is 1, otherwise 0. (2) A two-choice-bit version of random oblivious transfer (ROT), where the sender $\mathcal{S}$ and the receiver $\mathcal{R}$ each holds a choice bit $e_{0},e_{1}$. $\mathcal{S}$ receives two random messages $r_{0},r_{1}$. If $e_{0}\oplus e_{1}=0$, $\mathcal{R}$ receives $r_{0}$, otherwise $r_{1}$ [^2] This special ROT is identical to the standard 1-out-of-2 ROT, where $e_{0}$ is determined by $\mathcal{S}$ to indicate whether it would swap the order of $r_{0}$ and $r_{1}$. We use the standard ROT to formally describe our protocol specifications in the later sections. Here, we boil down a number of steps to the two-choice-bit ROT for facilitating illustration.. The following is to elaborate that how to utilize these ingredients to realize the secret-sharing processes.
The process for $P_{2}$ to secret-share $Y_{2}$ is as follows: $P_{2}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$. For each item $x\in X_{2}$, $P_{2}$ and $P_{1}$ receive shares $e_{2,1}$ and $e_{1,2}$. If $x\in X_{1}$, $e_{2,1}\oplus e_{1,2}=1$, otherwise $e_{2,1}\oplus e_{1,2}=0$. Then $P_{2}$ acts as $\mathcal{S}$ and executes the two-choice-bit ROT with $P_{1}$. $P_{2}$ and $P_{1}$ each inputs $e_{2,1},e_{1,2}$. $P_{2}$ receives $r_{2,1}^{0},r_{2,1}^{1}$. If $e_{2,1}\oplus e_{1,2}=0$, $P_{1}$ receives $r_{2,1}^{0}$, otherwise $P_{1}$ receives $r_{2,1}^{1}$. $P_{1}$ sets the output as its share $s_{2,1}$. $P_{2}$ sets its share $s_{2,2}$ to be $r_{2,1}^{0}\oplus x\|\mathsf{H}(x)$ [^3] Suppose a lack of specific structure for set elements, then $P_{1}$ cannot distinguish a set element $x$ and a random value $r$ when it reconstructs the difference set. To address this, all parties append the hash value when secret sharing an element, i.e., sharing $x\|\mathsf{H}(x)$. It is provable that if the output length of $\mathsf{H}$ is sufficiently long (according to [LG23], at least $\sigma+\log(m-1)+\log n$ bits), the probability of existing $x$ that satisfies $x\|\mathsf{H}(x)=r$ is negligible. Therefore, $P_{1}$ can distinguish set elements from random values with overwhelming probability.. $P_{3}$ sets $s_{2,3}$ to 0. If $x\notin X_{1}$, $e_{2,1}\oplus e_{1,2}=0$, $s_{2,1}\oplus s_{2,2}\oplus s_{2,3}=r_{2,1}^{0}\oplus(r_{2,1}^{0}\oplus x\|\mathsf{H}(x))=x\|\mathsf{H}(x)$. Otherwise, $e_{2,1}\oplus e_{1,2}=1$, $s_{2,1}\oplus s_{2,2}\oplus s_{2,3}=r_{2,1}^{1}\oplus r_{2,1}^{0}\oplus x\|\mathsf{H}(x)$ is uniformly random. That is, $Y_{2}$ is secret-shared among all parties [^4] Since $X_{2}\setminus X_{1}$ merely contains the information of $X_{1}$ and $X_{2}$, it is sufficient to secret-share $X_{2}\setminus X_{1}$ among $P_{1}$ and $P_{2}$. Other difference sets are shared similarly., and the other elements in $X_{2}$ are masked by random values before being secret-share.
We proceed to the process for $P_{3}$ to secret-share $Y_{3}$: $P_{3}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$ and $P_{2}$ separately. For each $x\in X_{3}$, $P_{3}$ and $P_{1}$ receive shares $e_{3,1}$ and $e_{1,3}$, while $P_{3}$ and $P_{2}$ receive shares $e_{3,2}$ and $e_{2,3}$. Then $P_{3}$ acts as $\mathcal{S}$ and executes the two-choice-bit ROT with $P_{1}$ and $P_{2}$. In the ROT between $P_{3}$ and $P_{1}$, $P_{3}$ inputs $e_{3,1}$ and $P_{1}$ inputs $e_{1,3}$. $P_{3}$ receives $r_{3,1}^{0},r_{3,1}^{1}$. If $e_{3,1}\oplus e_{1,3}=0$, $P_{1}$ receives $r_{3,1}^{0}$, otherwise $P_{1}$ receives $r_{3,1}^{1}$. $P_{1}$ sets the output as its share $s_{3,1}$. In the ROT between $P_{3}$ and $P_{2}$, $P_{3}$ inputs $e_{3,2}$ and $P_{2}$ inputs $e_{2,3}$. $P_{3}$ receives $r_{3,2}^{0},r_{3,2}^{1}$. If $e_{3,2}\oplus e_{2,3}=0$, $P_{2}$ receives $r_{3,2}^{0}$, otherwise $P_{2}$ receives $r_{3,2}^{1}$. $P_{2}$ sets the output as its share $s_{3,2}$. $P_{3}$ sets its share $s_{3,3}$ to be $r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x)$. If $x\notin X_{1}$ and $x\notin X_{2}$, $e_{3,1}\oplus e_{1,3}=0$ and $e_{3,2}\oplus e_{2,3}=0$, $s_{3,1}\oplus s_{3,2}\oplus s_{3,3}=r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus(r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x))=x\|\mathsf{H}(x)$. Otherwise, there is at least one random value $r$ ($r_{3,1}^{1}$ or $r_{3,2}^{1}$) cannot be canceled out from the summation.
The LG protocol described above has two main drawbacks: First, as the heart of LG protocol, mq-ssPMT is given a heavy instantiation which is fed much computation task through expensive general MPC machinery and renders it the bottleneck of the entire protocol. Second, LG protocol fails to achieve security against arbitrary collusion. In the following two sections, we are devoted to improving the efficiency of mq-ssPMT and enhancing LG protocol into standard semi-honest security, respectively.
2.2 Efficient Batch ssPMT
To improve the efficiency of mq-ssPMT, we abstract a new functionality called batch ssPMT, which is essentially the batched version of single-query ssPMT. In the batch ssPMT functionality, the sender $\mathcal{S}$ inputs $n$ disjoint sets $X_{1},\cdots,X_{n}$, and the receiver $\mathcal{R}$ inputs $n$ elements $y_{1},\cdots,y_{n}$. $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $n$, whose $i$th bit is 1 if $y_{i}\in X_{i}$, otherwise 0. The relationship between batch ssPMT and mq-ssPMT is two-fold: In terms of functionality, batch ssPMT is to mq-ssPMT what batch OPRF is to multi-point OPRF, testing a batch of elements’ memberships across distinct sets rather than a common set; In terms of efficiency, batch ssPMT admits a more efficient construction, providing a superior alternative to mq-ssPMT in the context of MPSU.
We adopt the following two-step approach to build the batch ssPMT protocol: (1) $\mathcal{S}$ and $\mathcal{R}$ invoke batched oblivious programmable pseudorandom function (batch OPPRF) [KMP+17, PSTY19]. To elaborate, for the $i$th ssPMT instance ($i=1,\cdots,n$), $\mathcal{S}$ choose a random $s_{i}$, and creates a set $E_{i}$ comprising key-value pairs with elements from $X_{i}$ as keys and $s_{i}$ as all elements’ corresponding values. Then $\mathcal{S}$ and $\mathcal{R}$ invoke batch OPPRF of size $n$, where in the $i$th instance of OPPRF, $\mathcal{S}$ inputs $E_{i}$, while $\mathcal{R}$ inputs $y_{i}$ and receives $t_{i}$. According to the definition of OPPRF (cf. Section 3.5), if $y_{i}\in X_{i}$, $t_{i}$ equals $s_{i}$, otherwise $t_{i}$ is pseudorandom to $s_{i}$. If $y_{i}\notin X_{i}$, the probability of $t_{i}=s_{i}$ is $2^{-\gamma}$, where $\gamma$ is the output length of OPPRF, i.e., the bit length of $s_{i}$ and $t_{i}$. To ensure that the probability of any $t_{i}\neq s_{i}$ occurring is less than $2^{-\sigma}$ ($\sigma$ is the statistical security parameter), we set $\gamma\geq\sigma+\log n$ so that $n\cdot 2^{-\gamma}\leq 2^{-\sigma}$. (2) $\mathcal{S}$ and $\mathcal{R}$ invoke $n$ instances of secret-shared private equality test (ssPEQT) [PSTY19, CGS22], where in the $i$th instance of ssPEQT, $\mathcal{S}$ inputs $s_{i}$ and $\mathcal{R}$ inputs $t_{i}$. If $s_{i}=t_{i}$, they receive secret shares of 1, otherwise 0. It’s easy to verify that if $y_{i}\in X_{i}$, $\mathcal{S}$ and $\mathcal{R}$ receive shares of 1, otherwise they receive shares of 0 with overwhelming probability.
Now we start to build an alternative to mq-ssPMT using our batch ssPMT. Recall that in mq-ssPMT, the sender $\mathcal{S}$’s input is a single set $X$. First, $\mathcal{S}$ and $\mathcal{R}$ preprocess their inputs through hashing to bin technique: $\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign its input elements $y_{1},\cdots,y_{n}$ to $B$ bins via Cuckoo hashing [PR04], which ensures that each bin accommodates at most one element. At the same time, $\mathcal{S}$ assigns each $x\in X$ to all bins determined by $h_{1}(x),h_{2}(x),h_{3}(x)$. Then the parties invoke $B$ instances of ssPMT, where in the $j$th instance, $\mathcal{S}$ inputs the subset $X_{j}\in X$ containing all elements mapped into its bin $j$ and $\mathcal{R}$ inputs the sole element mapped into its bin $j$. If some $y_{i}$ is mapped to bin $j$, and $y_{i}\in X$, then $\mathcal{S}$ certainly maps $y_{i}$ to bin $j$ (and other bins) as well, i.e., $y_{i}\in X_{j}$. Therefore, for each bin $j$ of $\mathcal{R}$, if the inside element $y_{i}$ is in $X$, we have $y_{i}\in X_{j}$, then $\mathcal{S}$ and $\mathcal{R}$ receive shares of 1 from the $i$th instance of batch ssPMT. Otherwise, we have $y_{i}\notin X_{j}$, then $\mathcal{S}$ and $\mathcal{R}$ receive shares of 0.
The above process in fact realizes a different functionality with mq-ssPMT. In the above construction, the query sequence of $\mathcal{R}$ is determined by the Cuckoo hash positions of its input elements. Since the Cuckoo hash positions depends on the whole input set of $\mathcal{R}$, and all shares is arranged in the order of the parties’ Cuckoo hash positions, the reconstruction stage may leak some information about the parties’ input sets to $P_{1}$ and hence cannot be simulated. Therefore, it is crucial to eliminate the dependence of shares’ order on the parties’ Cuckoo hash positions before the reconstruction, while fortunately, it has been realized by the execution of multi-party secret-shared shuffle as we discussed before. For the difference in efficiency, refer to Section 7.1.
2.3 SK-MPSU from Batch ssPMT and mss-ROT
By replacing mq-ssPMT with the construction using batch ssPMT plus hash to bin technique as previous, we can improve the performance of LG protocol significantly. However, the improved protocol still relies on the non-collusion assumption. So it remains to show that how to eliminate the non-collusion assumption to obtain a SK-MPSU protocol in the standard semi-honest model.
First, let us figure out that why LG protocol needs the non-collusion assumption. We still use the aforementioned three-party PSU protocol as illustration: After the secret sharing process of $P_{3}$, all parties hold the secret shares of
$\begin{cases}x\|\mathsf{H}(x)&x\in Y_{3}\\ r_{3,1}^{0}\oplus r_{3,1}^{1}\oplus x\|\mathsf{H}(x)&x\in Y_{1}\\ r_{3,2}^{0}\oplus r_{3,2}^{1}\oplus x\|\mathsf{H}(x)&x\in Y_{2}\\ r_{3,1}^{1}\oplus r_{3,2}^{1}\oplus r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x)&x\in X_{2}\cap X_{1}\end{cases}$
In the final stage, if $x\in Y_{3}$, $P_{1}$ reconstructs the secret $x\|\mathsf{H}(x)$. Otherwise, the secret is uniformly at random in the $P_{1}$’s view so that $P_{1}$ gains no information about $x$. Nevertheless, this is only true when $P_{1}$ does not collude with $P_{3}$. Since there is no randomness in $P_{3}$’s view (In the two invocations of ROT, $P_{3}$ receives $r_{3,1}^{0},r_{3,1}^{1}$ and $r_{3,2}^{0},r_{3,2}^{1}$), for each $x\in X_{3}$, the coalition of $P_{1}$ and $P_{3}$ can easily check the reconstructed values, and distinguish the cases that $x\in Y_{1}$, $x\in Y_{2}$, and $x\in X_{2}\cap X_{1}$, which reveals the information of $P_{2}$’s inputs.
To resolve the above security problem, we generalize ROT into multi-party setting, which we call multi-party secret-shared random oblivious transfer (mss-ROT). Suppose there are $d<m$ involved parties and two of them, denoted $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$, possessing choice bits $b_{0}$ and $b_{1}$ respectively. The mss-ROT functionality gives each involved party $P_{i}$ ($i\in\{1,\cdots,d\}$) two outputs $r_{i}$ and $\Delta_{i}$, such that if $b_{0}\oplus b_{1}=0$, $r_{1}\oplus\cdots\oplus r_{d}=0$. Otherwise, $r_{1}\oplus\cdots\oplus r_{d}=\Delta_{1}\oplus\cdots\oplus\Delta_{d}$. Note that if $b_{0}\oplus b_{1}=1$, the parties share a value which is uniformly random in the view of a coalition of any $d-1$ parties.
Equipped with this new primitive, the above protocol can be amended to achieve security under arbitrary collusion: After $P_{3}$ and $P_{1}$ receiving shares $e_{3,1}$ and $e_{1,3}$, while $P_{3}$ and $P_{2}$ receiving shares $e_{3,2}$ and $e_{2,3}$, $P_{1},P_{2},P_{3}$ invoke mss-ROT twice. In the first invocation, $P_{3}$ and $P_{1}$ act as $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ holding as choice bits $e_{3,1}$ and $e_{1,3}$. $P_{1},P_{2},P_{3}$ receives $r_{1,31},r_{2,31},r_{3,31}$ and $\Delta_{1,31},\Delta_{2,31},\Delta_{3,31}$ separately. If $e_{3,1}\oplus e_{1,3}=0$, $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=0$, otherwise $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=\Delta_{1,31}\oplus\Delta_{2,31}\oplus\Delta_{3,31}$. Likewise, in the second invocation, $P_{3}$ and $P_{2}$ hold as choice bits $e_{3,2}$ and $e_{2,3}$. $P_{1},P_{2},P_{3}$ receives $r_{1,32},r_{2,32},r_{3,32}$ and $\Delta_{1,32},\Delta_{2,32},\Delta_{3,32}$ separately. If $e_{3,2}\oplus e_{2,3}=0$, $r_{1,32}\oplus r_{2,32}\oplus r_{3,32}=0$, otherwise $r_{1,32}\oplus r_{2,32}\oplus r_{3,32}=\Delta_{1,32}\oplus\Delta_{2,32}\oplus\Delta_{3,32}$. Finally, $P_{1}$ sets its share $sh_{1}$ to be $r_{1,31}^{0}\oplus r_{1,32}^{0}$. $P_{2}$ sets its share $sh_{2}$ to be $r_{2,31}^{0}\oplus r_{2,32}^{0}$. $P_{3}$ sets its share $sh_{3}$ to be $r_{3,31}^{0}\oplus r_{3,32}^{0}\oplus x\|\mathsf{H}(x)$.
We conclude that all parties hold the secret shares of
$\begin{cases}x\|\mathsf{H}(x)&x\in Y_{3}\\ \Delta_{1,31}\oplus\Delta_{2,31}\oplus\Delta_{3,31}\oplus x\|\mathsf{H}(x)&x\in Y_{1}\\ \Delta_{1,32}\oplus\Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in Y_{2}\\ \Delta_{1,31}\oplus\Delta_{2,31}\oplus\Delta_{3,31}\oplus\Delta_{1,32}\oplus\Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in X_{2}\cap X_{1}\end{cases}$
It is immediate that if $x\notin Y_{3}$, the secret-shared value is uniform and independent of $x$ in the view of a coalition of any two parties. Hence, the reconstruction reveals no additional information to $P_{1}$ even if it colludes with $P_{2}$ or $P_{3}$.
2.4 PK-MPSU from Batch ssPMT and MKR-PKE
As previously mentioned, there are no existing MPSU works that achieve both linear computation and communication complexity. Thereinto, LG protocol and our SK-MPSU fail to achieve linear complexity because in the secret-sharing based framework, each party holds shares of $(m-1)n$ elements and in the final reconstruction stage, all parties except $P_{1}$ send their shares to $P_{1}$, then $P_{1}$ reconstruct $(m-1)n$ secrets to pick up all elements comprising the union. It is clear that even in this single stage, the computation and communication complexity of $P_{1}$ deviate from the linear criteria. So we have to turn to another approach.
We notice that the most advanced work in terms of asymptotic complexity is [GNT23], achieving near-linear complexity. We start by analyzing [GNT23]: The first phase enables $P_{1}$ to somehow acquire encrypted $X_{i}\setminus(X_{1}\cup\cdots\cup X_{i-1})$ for $2\leq i\leq m$, interspersed with encrypted dummy messages filling positions of the duplicate elements. Note that if $P_{1}$ decrypts these ciphertexts by itself, then it can establish a mapping associating each element $x\in X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$ with the party to whom $x$ belongs. To address this problem, they introduce a PKE variant, MKR-PKE, with several tailored properties, enabling all parties to conduct a collaborative decryption and shuffle procedure.
The second phase, which is the aforemetioned collaborative decryption and shuffle procedure, works as follows: $P_{1}$ sends the ciphertexts to $P_{2}$ after permuting them using a random permutation $\pi_{1}$. $P_{2}$ performs partial decryption on the received ciphertexts before rerandomization, and permutes them using a random permutation $\pi_{2}$. Then it forwards the permuted partially decrypted ciphertexts to $P_{3}$. This iterative process continues until the last party, $P_{m}$, who returns its permuted partially decrypted ciphertexts to $P_{1}$. $P_{1}$ decrypts the ultimate ciphertexts, filters out the dummy elements, retains the desired set $X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$, and appends the elements in $X_{1}$, to obtain the union.
We identify that their non-linear complexity stems from their construction of the first phase. Despite that they also use hashing to bin technique, their construction does not support amortizing the batching cost. In the case $B=O(n)$ bins are considered, each party has to execute mOT and cOPRF of size $O(\log\log n)$ per bin, resulting in overall $O(n\log\log n)$ computation and communication complexity. Besides, their insecurity against arbitrary collusion roots in the construction of the first phase (cf. Appendix 0.A). In contrast, the second phase already has linear computation and communication complexity, as well as security against arbitrary collusion. [^5] As we have said, the second phase addresses the same problem solved by the multi-party secret-shared shuffle, meanwhile there currently exists no construction for multi-party secret-shared shuffle that achieves total linear complexity. This is the root why our PK-MPSU protocol can achieve total linear complexity while our SK-MPSU protocol cannot. The key distinction lies in our PK-MPSU protocol’s departure from the secret-shared paradigm, opting instead for a PKE-based approach to replace the non-linear component. Therefore, the task of devising linear-complexity MPSU reduces to devising a linear-complexity and secure construction against arbitrary collusion for the first phase.
Our first attempt is as follows: For $2\leq i\leq m$, each $P_{i}$ and $P_{1}$ execute batch ssPMT (after preprocessing their inputs using hashing to bin, we ignore this for the sake of simplicity in the later explanation) and receive secret shares as the choice bits for the subsequent execution of two-choice-bit OT [^6] The difference of functionality between two-choice-bit OT and two-choice-bit ROT is that $\mathcal{S}$ inputs two messages $m_{0},m_{1}$ instead of obtaining two uniform messages., where for each $x\in X_{i}$, $P_{i}$ acts as $\mathcal{S}$ and inputs the encrypted $x$ and an encrypted dummy message. As a result, if $x$ belongs to the difference set $X_{i}\setminus X_{1}$, $P_{1}$ receives the encrypted $x$; otherwise, it receives the encrypted dummy message. Nevertheless, the goal of the first phase is to obtain the encrypted difference set $X_{i}\setminus(X_{1}\cup\cdots\cup X_{i-1})$, which is the subset of $X_{i}\setminus X_{1}$.
In order to attain the above goal, our next strategy is to let each $P_{i}$ engage in a procedure with each $P_{j}$ ($1<j<i$) before executing the batch ssPMT and two-choice-bit OT with $P_{1}$, so that $P_{j}$ can “obliviously” replace the encrypted items in the intersection $X_{j}\cap X_{i}$ with encrypted dummy messages. The procedure unfolds as follows.
$P_{i}$ plays the role of $\mathcal{R}$ and executes the batch ssPMT with $P_{j}$. For each $x\in X_{i}$, $P_{i}$ and $P_{j}$ receive shares $e_{i,j}$ and $e_{j,i}$. $P_{i}$ initializes a message $m_{0}$ to the encrypted $x$ and $m_{1}$ to an encrypted dummy message, then executes two-choice-bit OT with $P_{j}$. $P_{i}$ and $P_{j}$ input $e_{i,j}$ and $e_{j,i}$ as their choice bits. $P_{i}$ acts as $\mathcal{S}$ and inputs $m_{0}$ and $m_{1}$. $P_{j}$ acts as $\mathcal{R}$ and receives $\mathsf{ct}=m_{e_{i,j}\oplus e_{j,i}}$. If $x\notin X_{j}$, $\mathsf{ct}$ is the ciphertext of $x$; otherwise, the ciphertext of dummy message. $P_{j}$ rerandomizes $\mathsf{ct}$ to $\mathsf{ct}^{\prime}$ and returns $\mathsf{ct}^{\prime}$ to $P_{i}$. $P_{i}$ rerandomizes $\mathsf{ct}^{\prime}$, then updates $m_{0}$ to $\mathsf{ct}^{\prime}$ and $m_{1}$ to the rerandomized $m_{1}$ before repeating the above procedure with the next party $P_{j+1}$. The final values of $m_{0}$ and $m_{1}$ shall be used during the subsequent invocation of two-choice-bit OT with $P_{1}$.
3 Preliminaries
3.1 Notation
[…]
3.2 Multi-party Private Set Union
The ideal functionality of MPSU is formalized in Figure 2.
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$, where $P_{\textnormal{ld}}$ is the leader, $\textnormal{ld}\in[m]$. Size $n$ of input sets. The bit length $l$ of set elements.
  • Functionality. On input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$ from $P_{i}$, give $\bigcup_{i=1}^{m}X_{i}$ output to $P_{\textnormal{ld}}$.
Figure 2: Multi-party Private Set Union Functionality $\mathcal{F}_{\mathsf{mpsu}}$
3.3 Batch Oblivious Pseudorandom Function
Oblivious pseudorandom function (OPRF) [FIPR05] is a central primitive in the area of PSO. Kolesnikov et al. [KKRT16] introduced batched OPRF, which provides a batch of OPRF instances in the following way. In the $i$th instance, the receiver $\mathcal{R}$ has an input $x_{i}$; the sender $\mathcal{S}$ learns a PRF key $k_{i}$ and $\mathcal{R}$ learns $\mathsf{PRF}(k_{i},x_{i})$. Note that $\mathcal{R}$ evaluates the PRF on only one point per key. The batch OPRF functionality is described formally in Figure 3.
  • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of set elements. Some PRF family $\mathsf{PRF}:\{0,1\}^{*}\times\{0,1\}^{l}\to\{0,1\}^{\gamma}$.
  • Functionality. On input $\vec{x}\subseteq(\{0,1\}^{l})^{B}$ from $\mathcal{R}$,
    • For each $i\in[B]$, choose uniform PRF key $k_{i}$.
    • For each $i\in[B]$, define $f_{i}=\mathsf{PRF}(k_{i},x_{i})$.
    • Give vector $\vec{k}=(k_{1},\cdots,k_{B})$ to $\mathcal{S}$ and vector $\vec{f}=(f_{1},\cdots,f_{B})$ to $\mathcal{R}$.
Figure 3: Batch OPRF Functionality $\mathcal{F}_{\mathsf{bOPRF}}$
3.4 Oblivious Key-Value Stores
A key-value store [PRTY20, GPR+21, RR22, BPSY23] is a data structure that compactly represents a desired mapping from a set of keys to corresponding values.
Definition 1
A key-value store is parameterized by a set $\mathcal{K}$ of keys, a set $\mathcal{V}$ of values, and a set of functions $\mathsf{H}$, and consists of two algorithms:
  • $\mathsf{Encode_{H}}$ takes as input a set of key-value pairs $\{(k_{i},v_{i})|i\in[n]\}$ and outputs an object $D$ (or, with statistically small probability, an error indicator $\perp$).
  • $\mathsf{Decode_{H}}$ takes as input an object $D$, a key $k$, and outputs a value $v$.
Correctness. For all $A\subseteq\mathcal{K}\times\mathcal{V}$ with distinct keys:
$(k,v)\in A\ \ and\ \perp\neq D\leftarrow\mathsf{Encode_{H}}(A)\Longrightarrow\mathsf{Decode_{H}}(D,k)=v$
Obliviousness. For all distinct $\{k_{1}^{0},\cdots,k_{n}^{0}\}$ and all distinct $\{k_{1}^{1},\cdots,k_{n}^{1}\}$, if $\mathsf{Encode_{H}}$ does not output $\perp$ for $\{k_{1}^{0},\cdots,k_{n}^{0}\}$ or $\{k_{1}^{1},\cdots,k_{n}^{1}\}$, then the distribution of $D_{0}\leftarrow\mathsf{Encode_{H}}(\{(k_{1}^{0},v_{1}),\cdots,(k_{n}^{0},v_{n})\})$ is computationally indistinguishable to $D_{1}\leftarrow\mathsf{Encode_{H}}(\{(k_{1}^{1},v_{1}),\cdots,(k_{n}^{1},v_{n})\})$, where $v_{i}\leftarrow\mathcal{V}$ for $i\in[n]$.
A KVS is an Oblivious KVS (OKVS) if it satisfies the obliviousness property. we also require an additional randomness property [ZCL+23] from the OKVS.
Randomness. For any $A=\{(k_{1},v_{1}),\cdots,(k_{n},v_{n})\}$ and $k^{*}\notin\{k_{1},\cdots,k_{n}\}$, the output of $\mathsf{Decode_{H}}(D,k^{*})$ is statistically indistinguishable to that of uniform distribution over $\mathcal{V}$, where $D\leftarrow\mathsf{Encode_{H}}(A)$.
3.5 Batch Oblivious Programmable Pseudorandom Function
Oblivious programmable pseudorandom function (OPPRF) [KMP+17, PSTY19, CGS22, RS21, RR22] is an extension of OPRF, which lets $\mathcal{S}$ program a PRF $F$ so that it has specific random outputs for some specific inputs and pseudorandom outputs for all other inputs. This kind of PRF with the additional property that on a certain programmed set of inputs the function outputs programmed values is called programmable PRF (PPRF) [PSTY19]. $\mathcal{R}$ evaluates the OPPRF with no knowledge of whether it learns a programmed output of $F$ or just a pseudorandom value. The batch OPPRF functionality is given in Figure 4.
  • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of keys. The bit length $\gamma$ of values. An OKVS scheme $(\mathsf{Encode_{H}},\mathsf{Decode_{H}})$.
  • Sender’s inputs. $\mathcal{S}$ inputs $B$ sets of key-value pairs including:
    • Disjoint key sets $K_{1},\cdots,K_{B}$.
    • The value sets $V_{1},\cdots,V_{B}$, where $\lvert K_{i}\rvert=\lvert V_{i}\rvert$ for every $i\in[B]$, and $V_{i}(j)\in\{0,1\}^{\gamma}$ for every $i\in[B]$ and $1\leq j\leq\lvert K_{i}\rvert$.
  • Receiver’s inputs. $\mathcal{R}$ inputs $B$ queries $\vec{x}\subseteq(\{0,1\}^{l})^{B}$.
  • Functionality: On input $(K_{1},\cdots,K_{B})$ and $(V_{1},\cdots,V_{B})$ from $\mathcal{S}$ and $\vec{x}\subseteq(\{0,1\}^{l})^{B}$ from $\mathcal{R}$,
    • Choose uniformly random and independent PPRF key $k_{i}$, for each $i\in[B]$;
    • Sample a PPRF $F:\{0,1\}^{*}\times\{0,1\}^{l}\to\{0,1\}^{\gamma}$ such that $F(k_{i},K_{i}(j))=V_{i}(j)$ for $i\in[B],1\leq j\leq\lvert K_{i}\rvert$;
    • Define $f_{i}=F(k_{i},x_{i})$, for each $i\in[B]$;
    • Give vector $\vec{f}=(f_{1},\cdots,f_{B})$ to $\mathcal{R}$.
Figure 4: Batch OPPRF Functionality $\mathcal{F}_{\mathsf{bOPPRF}}$
3.6 Hashing to Bin
The hashing to bin technique was introduced by Pinkas et al. [PSZ14, PSSZ15], which is originally applied to construct two-party PSI protocol. At the high level, the receiver $\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign its items to $B$ bins via Cuckoo hashing [PR04], so that each bin has at most one item. The hashing process uses eviction and the choice of which of the bins is used depends on the entire set. Using the same hash functions and simple hashing, sender $\mathcal{S}$ assigns each of its items $x$ to all of the bins $h_{1}(x),h_{2}(x),h_{3}(x)$, so that if the item $x$ is also in $\mathcal{R}$’s set and is mapped into the bin $b\in\{h_{1}(x),h_{2}(x),h_{3}(x)\}$ by Cuckoo hashing, then the bin $b$ of $\mathcal{S}$’s simple hash table certainly contains $x$.
We write simple hashing with the following notation:
$\mathcal{T}^{1},\cdots,\mathcal{T}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X)$
This expression means to hash the items of $X$ into $B$ bins using simple hashing with hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$. The output is the simple hash table denoted by $\mathcal{T}^{1},\cdots,\mathcal{T}^{B}$, where for each $x\in X$ we have $\mathcal{T}^{h_{i}(x)}=\{x\|i|i=1,2,3\}$. [^7] Appending the index of the hash function is helpful for dealing with edge cases like $h_{1}(x)=h_{2}(x)=i$, which happen with non-negligible probability. Without appending $i$, the OKVS would contain the identical value $F(k_{i},x)$, which leaks the fact that a collision $h_{1}(x)=h_{2}(x)$ occurred. Such an event is input-dependent so cannot be simulated.
We write Cuckoo hashing with the following notation:
$\mathcal{C}^{1},\cdots,\mathcal{C}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X)$
This expression means to hash the items of $X$ into $B$ bins using Cuckoo hashing on hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$. The output is the Cuckoo hash table denoted by $\mathcal{C}^{1},\cdots,\mathcal{C}^{B}$, where for each $x\in X$ there is some $i\in\{1,2,3\}$ such that $\mathcal{C}^{h_{i}(x)}=x\|i$. Some Cuckoo hash positions do not matter, corresponding to empty bins. We use these symbols throughout the subsequent sections.
3.7 Secret-Shared Private Equality Test
Secret-shared private equality test protocol (ssPEQT) [PSTY19, CGS22] can be viewed as an extreme case of ssPMT when the sender $\mathcal{S}$’s input set has only one item. In Figure 5 we formally define this functionality.
  • Parameters. Two parties $P_{1},P_{2}$. The bit length $\gamma$ of inputs.
  • Functionality. On input $x$ from $P_{1}$ and input $y$ from $P_{2}$, sample two random bits $a,b$ under the constraint that $a\oplus b=\bigwedge_{i=1}^{\gamma}(1\oplus x[i]\oplus y[i])$. Namely, if $x=y$, $a\oplus b=1$. Otherwise $a\oplus b=0$. Give $a$ to $P_{1}$ and $b$ to $P_{2}$.
Figure 5: Secret-Shared Private Equality Test Functionality $\mathcal{F}_{\mathsf{ssPEQT}}$
3.8 Random Oblivious Transfer
Oblivious transfer (OT) [Rab05] is a foundational primitive in MPC, the functionality of 1-out-of-2 random OT (ROT) is given in Figure 6.
  • Parameters. Sender $\mathcal{S}$, Receiver $\mathcal{R}$. The message length $l$.
  • Functionality. On input $b\in\{0,1\}$ from $\mathcal{R}$, sample $r_{0},r_{1}\leftarrow\{0,1\}^{l}$. Give $(r_{0},r_{1})$ to $\mathcal{S}$ and give $r_{b}$ to $\mathcal{R}$.
Figure 6: 1-out-of-2 Random OT Functionality $\mathcal{F}_{\mathsf{rot}}$
3.9 Multi-Party Secret-Shared Shuffle
Multi-party secret-shared shuffle has the capability to permute the share vectors of all parties in a random permutation unknown to any coalition of $m-1$ parties. Then it refreshs all shares. The functionality is given in Figure 7.
Eskandarian et al. [EB22] propose an online-efficient protocol by extending [CGP20] to the multi-party setting. In the offline phase, each party generates a random permutation and a set of correlated vectors called share correlation. In the online phase, each party permutes and refreshes the share vectors efficiently using share correlation. We give the functionality of share correlation and details of the online protocol in Appendix 0.C.
  • Parameters. $m$ parties $P_{1},\cdots P_{m}$. The dimension of vector $n$. The item length $l$.
  • Functionality. On input $\vec{x}_{i}={(x_{i}^{1},\cdots,x_{i}^{n})}$ from each $P_{i}$, sample a random permutation $\pi:[n]\to[n]$. For $1\leq i\leq m$, sample $\vec{{x}_{i}^{\prime}}\leftarrow(\{0,1\}^{l})^{n}$ satisfying $\bigoplus_{i=1}^{m}\vec{{x}_{i}^{\prime}}=\pi(\bigoplus_{i=1}^{m}\vec{{x}_{i}})$. Give $\vec{{x}_{i}^{\prime}}$ to $P_{i}$.
Figure 7: Multi-Party Secret-Shared Shuffle Functionality $\mathcal{F}_{\mathsf{ms}}$
3.10 Multi-Key Rerandomizable Public-Key Encryption
Gao et al. [GNT23] introduce the concept of multi-key rerandomizable public-key encryption (MKR-PKE), which is a variant of PKE with several additional properties. Let $\mathcal{SK}$ denote the space of secret keys which forms an abelian group under the operation $+$, $\mathcal{PK}$ the space of public keys which forms an abelian group under the operation $\cdot$, $\mathcal{M}$ the space of plaintexts, and $\mathcal{C}$ the space of ciphertexts. MKR-PKE is a tuple of PPT algorithms $(\mathsf{Gen},\mathsf{Enc},\mathsf{ParDec},\mathsf{Dec},\mathsf{ReRand})$ such that:
  • The key-generation algorithm $\mathsf{Gen}$ takes as input a security parameter $1^{\lambda}$ and outputs a pair of keys $(pk,sk)\in\mathcal{PK}\times\mathcal{SK}$.
  • The encryption algorithm $\mathsf{Enc}$ takes as input a public key $pk\in\mathcal{PK}$ and a plaintext message $x\in\mathcal{M}$, and outputs a ciphertext $\mathsf{ct}\in\mathcal{C}$.
  • The partial decryption algorithm $\mathsf{ParDec}$ takes as input a secret key share $sk\in\mathcal{SK}$ and a ciphertext $\mathsf{ct}\in\mathcal{C}$, and outputs another ciphertext $\mathsf{ct}^{\prime}\in\mathcal{C}$.
  • The decryption algorithm $\mathsf{Dec}$ takes as input a secret key $sk\in\mathcal{SK}$ and a ciphertext $\mathsf{ct}\in\mathcal{C}$, outputs a message $x\in\mathcal{M}$ or an error symbol $\perp$.
  • The rerandomization algorithm $\mathsf{ReRand}$ takes as input a public key $pk\in\mathcal{PK}$ and a ciphertext $\mathsf{ct}\in\mathcal{C}$, outputs another ciphertext $\mathsf{ct}^{\prime}\in\mathcal{C}$.
MKR-PKE requires the following additional properties besides those of PKE with indistinguishable multiple encryptions. (cf. Appendix 0.B):
3.10.1 Partially Decryptable
For any two pairs of keys $(sk_{1},pk_{1})\leftarrow\mathsf{Gen}(1^{\lambda})$ and $(sk_{2},pk_{2})\leftarrow\mathsf{Gen}(1^{\lambda})$ and any $x\in\mathcal{M}$, it holds that
$\mathsf{ParDec}(sk_{1},\mathsf{Enc}(pk_{1}\cdot pk_{2},x))=\mathsf{Enc}(pk_{2},x)$
3.10.2 Rerandomizable
For any $pk\in\mathcal{PK}$ and any $x\in\mathcal{M}$, it holds that
[…]
[GNT23] uses elliptic curve based ElGamal encryption [ElG85] to instantiate MKR-PKE. The concrete EC MKR-PKE is described in Appendix 0.D.
Remark 1
We note that only a few of elliptic curves support efficient encoding from bit-strings to EC points. Therefore, the plaintext space of EC MKR-PKE is generally restricted to EC points to guarantee rerandomizable property.
4 Batch Secret-Shared Private Membership Test
The batch secret-shared private membership test (batch ssPMT) is a central building block in both our SK-MPSU and PK-MPSU protocols. In this section, we formally introduce this functionality and provide a construction built by the aforementioned primitives.
The batch ssPMT is a two-party protocol implementing a batch of ssPMT instances between a sender and a receiver. Assuming a batch size of $B$, the sender $\mathcal{S}$ inputs $B$ sets $X_{1},\cdots,X_{B}$, while the receiver $\mathcal{R}$ inputs $B$ elements $x_{1},\cdots,x_{B}$. Consequently, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $B$, whose $i$th bit is 1 if $x_{i}\in X_{i}$, otherwise 0. The batch ssPMT functionality is presented in Figure 8 and the construction is given in Figure 9.
  • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of set elements. The output length $\gamma$ of OPPRF.
  • Inputs. $\mathcal{S}$ inputs $B$ disjoint sets $X_{1},\cdots,X_{B}$ and $\mathcal{R}$ inputs $\vec{x}\subseteq(\{0,1\}^{l})^{B}$.
  • Functionality. On inputs $X_{1},\cdots,X_{B}$ from $\mathcal{S}$ and input $\vec{x}$ from $\mathcal{R}$, for each $i\in[B]$, sample two random bits $e_{S}^{i},e_{R}^{i}$ under the constraint that if $x_{i}\in X_{i},e_{S}^{i}\oplus e_{R}^{i}=1$, otherwise $e_{S}^{i}\oplus e_{R}^{i}=0$. Give $\vec{e}_{S}=(e_{S}^{1},\cdots,e_{S}^{B})$ to $\mathcal{S}$ and $\vec{e}_{R}=(e_{R}^{1},\cdots,e_{R}^{B})$ to $\mathcal{R}$.
Figure 8: Batch ssPMT Functionality $\mathcal{F}_{\mathsf{bssPMT}}$
  • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of set elements. The output length $\gamma$ of OPPRF.
  • Inputs. $\mathcal{S}$ inputs $B$ disjoint sets $X_{1},\cdots,X_{B}$ and $\mathcal{R}$ inputs $\vec{x}\subseteq(\{0,1\}^{l})^{B}$.
  • Protocol.
    1. 1.
      For each $i\in[B]$, $\mathcal{S}$ chooses random $s_{i}\leftarrow\{0,1\}^{\gamma}$ and computes a multiset $S_{i}$ comprising $\lvert X_{i}\rvert$ repeated elements that all equal to $s_{i}$.
    2. 2.
      The parties invoke $\mathcal{F}_{\mathsf{bOPPRF}}$ of batch size $B$. $\mathcal{S}$ acts as sender and inputs $X_{1},\cdots,X_{B}$ as key sets and $S_{1},\cdots,S_{B}$ as value sets. $\mathcal{R}$ acts as receiver with input $\vec{x}$, and receives a vector $\vec{t}=(t_{1},\cdots,t_{B})$.
    3. 3.
      The parties invoke $B$ instances of $\mathcal{F}_{\mathsf{ssPEQT}}$, where in the $i$th instance $\mathcal{S}$ inputs $s_{i}$ and $\mathcal{R}$ inputs $t_{i}$. In the end, $\mathcal{S}$ receives $e_{S}^{i}\in\{0,1\}$ and $\mathcal{R}$ receives $e_{R}^{i}\in\{0,1\}$.
Figure 9: Batch ssPMT $\Pi_{\mathsf{bssPMT}}$
  • Correctness. According to the functionality of batch OPPRF, if $x_{i}\in X_{i}$, then $\mathcal{R}$ receives $t_{i}=s_{i}$. Subsequently, in the $i$th instance of ssPEQT, they input $s_{i}$ and $t_{i}$ which are equal, hence their outputs satisfy $e_{S}^{i}\oplus e_{R}^{i}=1$. Conversely, if $x_{i}\notin X_{i}$, then $\mathcal{R}$ receives a pseudorandom value $t_{i}$. The probability of $t_{i}=s_{i}$ equals $2^{-\gamma}$ and any $t_{i}=s_{i},i\in[B]$ occurs with a probability of $B\cdot 2^{-\gamma}$. As we set $\gamma\geq\sigma+\log B$, so $B\cdot 2^{-\gamma}\leq 2^{-\sigma}$, which means the probability of any $t_{i}=s_{i}$ occurring is negligible. After the invocation of ssPEQT, we conclude that if $x_{i}\notin X_{i}$, $e_{S}^{i}$ and $e_{R}^{i}$ satisfy $e_{S}^{i}\oplus e_{R}^{i}=0$ with overwhelming probability.
Theorem 4.1
Protocol $\Pi_{\mathsf{bssPMT}}$ securely realizes $\mathcal{F}_{\mathsf{bssPMT}}$ in the $(\mathcal{F}_{\mathsf{bOPRF}},\mathcal{F}_{\mathsf{ssPEQT}})$-hybrid model.
The security of the protocol follows immediately from the security of the batch OPPRF and the ssPEQT functionalities.
  • Complexity Analysis. We set $B=O(n)$ to be consistent with our MPSU protocols. One thing to clarify is that the asymptotic complexity, to a great extent, depends on the choices of the instantiation of batch OPRF and the generation method of Beaver Triples [Bea91] in ssPEQT. In consideration of concrete efficiency, we employ subfield vector oblivious linear evaluation (subfield-VOLE) [BCG+19a, BCG+19b, RRT23] for both options. A comprehensive complexity analysis is in Appendix 0.F.1. We summarize the total costs as below:
    • Offline phase. The computation complexity of each party is $O(\gamma n\log n)$. The communication complexity of each party is $O(t\lambda\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(n)$. The communication complexity of each party is $O(\gamma n)$. The round complexity is $O(\log\gamma)$.
5 MPSU from Symmetric-Key Techniques
In this section, we introduce a new primitive called multi-party secret-shared random oblivious transfer (mss-ROT), then we utilize its two-choice-bit version to construct a SK-MPSU based on oblivious transfer and symmetric-key operations in the standard semi-honest model.
5.1 Multi-Party Secret-Shared Random Oblivious Transfer
The abstraction of the general version of mss-ROT is inspired by the generation of Beaver Triples in multi-party setting. In the two-party GMW protocol, a Beaver Triple is produced by two role-switching executions of 1-bit ROT [ALSZ13]. As a result, $P_{1}$ holds $a_{1},b_{1},c_{1}$ and $P_{2}$ holds $a_{2},b_{2},c_{2}$, such that $(a_{1}\oplus a_{2})\cdot(b_{1}\oplus b_{2})=c_{1}\oplus c_{2}$. In the multi-party setting, the functionality of Beaver Triple is extended to let each party $P_{i}$ hold $a_{i},b_{i},c_{i}$ with the correlation $(\bigoplus_{i=1}^{m}a_{i})\cdot(\bigoplus_{i=1}^{m}b_{i})=\bigoplus_{i=1}^{m}c_{i}$. Given that $(\bigoplus_{i=1}^{m}a_{i})\cdot(\bigoplus_{i=1}^{m}b_{i})$ can be written as $(2-m)(\bigoplus_{i=1}^{m}a_{i}\cdot b_{i})+\bigoplus_{1\leq i\leq j\leq m}(a_{i}+a_{j})\cdot(b_{i}+b_{j})$, multi-party Beaver Triple reduces to pairwise two-party Beaver Triples. In consideration of the simple fact that the functionality of ROT can be interpreted as $r_{0}\oplus r_{b}=b\cdot\Delta$, where $\Delta=r_{0}\oplus r_{1}$, we follow the approach to producing multi-party Beaver Triples to extend ROT into multi-party setting, which is to let each party $P_{i}$ to individually input a choice bit $b_{i}$ and receive the output shares $r_{i}$ and $\Delta_{i}$ such that $r_{1}\oplus\cdots\oplus r_{m}=(b_{1}\oplus\cdots\oplus b_{m})\cdot(\Delta_{1}\oplus\cdots\oplus\Delta_{m})$.
In the context of MPSU, we only need two-choice-bit version of mss-ROT, which allows a more efficient construction with less pairwise invocations of two-party ROT compared to the general version. To elaborate, the two-choice-bit version of mss-ROT (for simplicity, we call it mss-ROT for short hereafter) allows two parties $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ to hold their choice bits $b_{0}$ and $b_{1}$ s.t. $r_{1}\oplus\cdots\oplus r_{m}=(b_{0}\oplus b_{1})\cdot(\Delta_{1}\oplus\cdots\oplus\Delta_{m})$. We give the formal functionality in Figure 10 and the detailed construction in Figure 11.
Theorem 5.1
Protocol $\Pi_{\mathsf{mss\text{-}rot}}$ securely implements $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ in the presence of any semi-honest adversary corrupting $t<m$ parties in the $\mathcal{F}_{\mathsf{rot}}$-hybrid model.
It is easy to see that our construction essentially boils down to performing ROT pairwise. As one of the benefits, we can utilize the derandomization technique [Bea91] to bring most tasks forward to the offline phase. And the correctness and security of the mss-ROT protocol stems from the correctness and security of ROT. For the complete proof, refer to Appendix 0.E.1.
5.2 Construction of Our SK-MPSU
We now turn our attention to construct a SK-MPSU. The construction of our SK-MPSU follows the high-level ideas we introduced in the technical overview and is formally presented in Figure 12.
Theorem 5.2
Protocol $\Pi_{\mathsf{SK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{mss\text{-}rot}},\mathcal{F}_{\mathsf{ms}})$-hybrid model, where $P_{1}=P_{\textnormal{ld}}$.
The security proof of Theorem 5.2 is deferred to Appendix 0.E.2.
  • Complexity Analysis. We provide a comprehensive complexity analysis for our SK-MPSU protocol in Appendix 0.F.2. The total costs are summarized below:
    • Offline phase. The offline computation complexity per party is $O(\gamma mn\log n+m^{2}n(\log m+\log n))$. The offline communication complexity per party is $O(t\lambda m\log(\gamma n/t)+t\lambda m^{2}\log(n/t)+\lambda m^{2}n(\log m+\log n))$. The offline round complexity is $O(1)$.
    • Online phase. The online computation complexity per party is $O(m^{2}n)$. The online communication complexity of $P_{1}$ is $O(\gamma mn+(l+\kappa)m^{2}n)$. The online communication complexity of $P_{j}$ is $O(\gamma mn+m^{2}n+(l+\kappa)mn)$. The online round complexity is $O(\log\gamma+m)$.
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$, where $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ provide inputs as shares of the choice bit, $\text{ch}_{0},\text{ch}_{1}\in[m]$. The message length $l$.
  • Functionality. On input $b_{0}\in\{0,1\}$ from $P_{\text{ch}_{0}}$ and $b_{1}\in\{0,1\}$ from $P_{\text{ch}_{1}}$,
    • Sample $r_{2},\cdots r_{m},\Delta_{1},\Delta_{2},\cdots,\Delta_{m}\leftarrow\{0,1\}^{l}$ and give $(r_{j},\Delta_{j})$ to $P_{j}$ for $2\leq j\leq m$.
    • If $b_{0}\oplus b_{1}=0$, compute $r_{1}=\bigoplus_{j=2}^{m}r_{j}$, else compute $r_{1}=\Delta_{1}\oplus(\bigoplus_{j=2}^{m}(r_{j}\oplus\Delta_{j}))$. Give $r_{1}$ to $P_{1}$.
Figure 10: Multi-Party Secret-Shared Random OT Functionality $\mathcal{F}_{\mathsf{mss\text{-}rot}}$
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$. where $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ provide inputs as shares of the choice bit, $\text{ch}_{0},\text{ch}_{1}\in[m]$. We use $J$ to denote the set of indices for the remaining parties who do not provide a choice-bit input, i.e., $J=[m]\setminus\{\text{ch}_{0},\text{ch}_{1}\}$. The message length $l$.
  • Inputs. $P_{\text{ch}_{0}}$ has input $b_{0}\in\{0,1\}$ and $P_{\text{ch}_{1}}$ has input $b_{1}\in\{0,1\}$.
  • Protocol.
    1. 1.
      $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ invoke $\mathcal{F}_{\mathsf{rot}}$ twice: First, $P_{\text{ch}_{0}}$ acts as receiver with input $b_{0}$ and $P_{\text{ch}_{1}}$ acts as sender. $P_{\text{ch}_{0}}$ receives $r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{b_{0}}\in\{0,1\}^{l}$. $P_{\text{ch}_{1}}$ receives $r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{0},r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{1}\in\{0,1\}^{l}$; Second, $P_{\text{ch}_{1}}$ acts as receiver with input $b_{1}$ and $P_{\text{ch}_{0}}$ acts as sender. $P_{\text{ch}_{1}}$ receives $r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{b_{1}}\in\{0,1\}^{l}$. $P_{\text{ch}_{0}}$ receives $r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{0},r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{1}\in\{0,1\}^{l}$.
    2. 2.
      For $j\in J$: $P_{\text{ch}_{0}}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{\text{ch}_{0}}$ acts as receiver with input $b_{0}$ and $P_{j}$ as sender without input. $P_{\text{ch}_{0}}$ receives $r_{j,{\text{ch}_{0}}}^{b_{0}}\in\{0,1\}^{l}$. $P_{j}$ receives $r_{j,{\text{ch}_{0}}}^{0},r_{j,{\text{ch}_{0}}}^{1}\in\{0,1\}^{l}$.
    3. 3.
      For $j\in J$: $P_{\text{ch}_{1}}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{\text{ch}_{1}}$ acts as receiver with input $b_{1}$ and $P_{j}$ as sender. $P_{\text{ch}_{1}}$ receives $r_{j,{\text{ch}_{1}}}^{b_{1}}\in\{0,1\}^{l}$. $P_{j}$ receives $r_{j,{\text{ch}_{1}}}^{0},r_{j,{\text{ch}_{1}}}^{1}\in\{0,1\}^{l}$.
    4. 4.
      For $j\in J$: $P_{j}$ samples $\Delta_{j}\leftarrow\{0,1\}^{l}$ and computes $r_{j}=r_{j,{\text{ch}_{0}}}^{0}\oplus r_{j,{\text{ch}_{1}}}^{0}$. $P_{j}$ sends $u_{j,{\text{ch}_{0}}}=\Delta_{j}\oplus r_{j,{\text{ch}_{0}}}^{0}\oplus r_{j,{\text{ch}_{0}}}^{1}$ to $P_{\text{ch}_{0}}$, and $u_{j,{\text{ch}_{1}}}=\Delta_{j}\oplus r_{j,{\text{ch}_{1}}}^{0}\oplus r_{j,{\text{ch}_{1}}}^{1}$ to $P_{\text{ch}_{1}}$, then outputs $(r_{j},\Delta_{j})$.
    5. 5.
      $P_{\text{ch}_{0}}$ computes $\Delta_{\text{ch}_{0}}=r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{0}\oplus r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{1}$ and $r_{\text{ch}_{0}}=\bigoplus_{j\in J}(r_{j,{\text{ch}_{0}}}^{b_{0}}\oplus b_{0}\cdot u_{j,{\text{ch}_{0}}})\oplus r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{0}\oplus r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{b_{0}}\oplus b_{0}\cdot\Delta_{\text{ch}_{0}}$ ($\cdot$ denotes bitwise-AND between the repetition code of $b_{0}$ and $u_{j,{\text{ch}_{0}}}$, which are both strings of length $l$. Similarly hereinafter), then outputs $(r_{\text{ch}_{0}},\Delta_{\text{ch}_{0}})$. $P_{\text{ch}_{1}}$ computes $\Delta_{\text{ch}_{1}}=r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{0}\oplus r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{1}$ and $r_{\text{ch}_{1}}=\bigoplus_{j\in J}(r_{j,{\text{ch}_{1}}}^{b_{1}}\oplus b_{1}\cdot u_{j,{\text{ch}_{1}}})\oplus r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{0}\oplus r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{b_{1}}\oplus b_{1}\cdot\Delta_{\text{ch}_{1}}$, then outputs $(r_{\text{ch}_{1}},\Delta_{\text{ch}_{1}})$.
Figure 11: Multi-Party Secret-Shared Random OT $\Pi_{\mathsf{mss\text{-}rot}}$
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$. Size $n$ of input sets. The bit length $l$ of set elements. Cuckoo hashing parameters: hash functions $h_{1},h_{2},h_{3}$ and number of bins $B$. An OKVS scheme $(\mathsf{Encode_{H}},\mathsf{Decode_{H}})$. A collision-resisitant hash function $\mathsf{H}(x):\{0,1\}^{l}\to\{0,1\}^{\kappa}$.
  • Inputs. Each party $P_{i}$ has input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$.
  • Protocol.
    1. 1.
      Hashing to bin. For $1\leq j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
    2. 2.
      Batch secret-shared private membership test. For $1\leq i<j\leq m$: $P_{i}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{bssPMT}}$ of batch size $B$, where $P_{i}$ acts as sender with inputs $\mathcal{T}_{i}^{1},\cdots,\mathcal{T}_{i}^{B}$ and $P_{j}$ acts as receiver with inputs $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}$. For the instance $b\in[B]$, $P_{i}$ receives $e_{i,j}^{b}\in\{0,1\}$, and $P_{j}$ receives $e_{j,i}^{b}\in\{0,1\}$.
    3. 3.
      Multi-party secret-shared random oblivious transfers. For $1\leq i<j\leq m,1\leq b\leq B$: $P_{1},\cdots,P_{j}$ invoke $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ where $P_{i}$ acts as $P_{\text{ch}_{0}}$ with input $e_{i,j}^{b}$ and $P_{j}$ acts as $P_{\text{ch}_{1}}$ with input $e_{j,i}^{b}$. For $1\leq d<j$, $P_{d}$ receives $r_{d,ji}^{b},\Delta_{d,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{d,j}^{b}=\bigoplus_{i=1}^{j-1}r_{d,ji}^{b}$. $P_{j}$ receives $r_{j,ji}^{b},\Delta_{j,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{j,j}^{b}=\bigoplus_{i=1}^{j-1}r_{j,ji}^{b}\oplus(\mathsf{Elem}(\mathcal{C}_{j}^{b})\|\mathsf{H}(\mathsf{Elem}(\mathcal{C}_{j}^{b})))$ if $\mathcal{C}_{j}^{b}$ is not corresponding to an empty bin, otherwise chooses $u_{j,j}^{b}$ at random. $\mathsf{Elem}(\mathcal{C}_{j}^{b})$ denotes the element in $\mathcal{C}_{j}^{b}$.
    4. 4.
      Multi-party secret-shared shuffle.
      1. (a)
        For $1\leq i\leq m$, each party $P_{i}$ computes $\vec{sh}_{i}\in(\{0,1\}^{l+\kappa})^{(m-1)B}$ as follows: For $max(2,i)\leq j\leq m,1\leq b\leq B$, $sh_{i,(j-2)B+b}=u_{i,j}^{b}$. Set all other positions to 00.
      2. (b)
        For $1\leq i\leq m$, all parties $P_{i}$ invoke $\mathcal{F}_{\mathsf{ms}}$ with input $\vec{sh}_{i}$. $P_{i}$ receives $\vec{sh}_{i}^{\prime}$.
    5. 5.
      Output reconstruction. For $2\leq j\leq m$, $P_{j}$ sends $\vec{sh}_{j}^{\prime}$ to $P_{1}$. $P_{1}$ recovers $\vec{v}=\bigoplus_{i=1}^{m}\vec{sh}_{i}^{\prime}$ and sets $Y=\emptyset$. For $1\leq i\leq(m-1)B$, if $v^{\prime}_{i}=x\|\mathsf{H}(x)$ holds for some $x\in\{0,1\}^{l}$, adds $x$ to $Y$. Outputs $X_{1}\cup Y$.
Figure 12: Our SK-MPSU $\Pi_{\mathsf{SK\text{-}MPSU}}$
6 MPSU from Public-Key Techniques
In this section, we describe how to construct a PK-MPSU achieving both linear computation and linear communication complexity. The construction of our PK-MPSU is formally presented in Figure 13.
Theorem 6.1
Protocol $\Pi_{\mathsf{PK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{rot}})$-hybrid model, assuming the security of MKR-PKE scheme, where $P_{1}=P_{\textnormal{ld}}$.
The security of the above protocol is based on security of the underlying building blocks, batch ssPMT and ROT, along with the rerandomizable property and indistinguishable multiple encryptions of MKR-PKE. For the complete proof, refer to Appendix 0.E.3.
  • Complexity Analysis. We provide a comprehensive complexity analysis for our PK-MPSU protocol in Appendix 0.F.3. The total costs are summarized below:
    • Offline phase. The offline computation complexity per party is $O(\gamma mn\log n)$. The offline communication complexity per party is $O(t\lambda m\log(\gamma n/t))$. The offline round complexity is $O(1)$.
    • Online phase. The online computation complexity per party is $O(mn)$ symmetric-key operations and $O(mn)$ public-key operations. The onlne communication complexity per party is $O((\gamma+\lambda)mn)$. The online round complexity is $O(\log\gamma+m)$.
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$. Size $n$ of input sets. The bit length $l$ of set elements. Cuckoo hashing parameters: hash functions $h_{1},h_{2},h_{3}$ and number of bins $B$. An OKVS scheme $(\mathsf{Encode_{H}},\mathsf{Decode_{H}})$. A MKR-PKE scheme $\mathcal{E}=(\mathsf{Gen},\mathsf{Enc},\mathsf{ParDec},\mathsf{Dec},\mathsf{ReRand})$.
  • Inputs. Each party $P_{i}$ has input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$.
  • Protocol. Each party $P_{i}$ runs $(pk_{i},sk_{i})\leftarrow\mathsf{Gen}(1^{\lambda})$, and distributes its public key $pk_{i}$ to other parties. Define $sk=sk_{1}+\cdots+sk_{m}$ and each party can compute its associated public key $pk=\prod_{i=1}^{m}pk_{i}$.
    1. 1.
      Hashing to bin. For $1\leq j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
    2. 2.
      Batch secret-shared private membership test. For $1\leq i<j\leq m$: $P_{i}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{bssPMT}}$ of batch size $B$, where $P_{i}$ acts as sender with inputs $\mathcal{T}_{i}^{1},\cdots,\mathcal{T}_{i}^{B}$ and $P_{j}$ acts as receiver with inputs $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}$. For the instance $b\in[B]$, $P_{i}$ receives $e_{i,j}^{b}\in\{0,1\}$, and $P_{j}$ receives $e_{j,i}^{b}\in\{0,1\}$.
    3. 3.
      Random oblivious transfers and messages rerandomization.
      1. (a)
        For $2\leq j\leq m,1\leq b\leq B$: $P_{j}$ defines a vector $\vec{c}_{j}$ and sets $c_{j}^{b}=\mathsf{Enc}(pk,\mathsf{Elem}(\mathcal{C}_{j}^{b}))$. $\mathsf{Elem}(\mathcal{C}_{j}^{b})$ denotes the element in $\mathcal{C}_{j}^{b}$.
        For $2\leq i<j$: $P_{i}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{i}$ acts as receiver with input $e_{i,j}^{b}$ and $P_{j}$ acts as sender. $P_{i}$ receives $r_{j,i,e_{i,j}^{b}}^{b}\in\{0,1\}^{\lambda}$. $P_{j}$ receives $r_{j,i,0}^{b},r_{j,i,1}^{b}\in\{0,1\}^{\lambda}$. $P_{j}$ computes $u_{j,i,e_{j,i}^{b}}^{b}=r_{j,i,e_{j,i}^{b}}^{b}\oplus c_{j}^{b}$, $u_{j,i,e_{j,i}^{b}\oplus 1}^{b}=r_{j,e_{j,i}^{b}\oplus 1}^{b}\oplus\mathsf{Enc}(pk,\perp)$, then sends $u_{j,i,0}^{b},u_{j,i,1}^{b}$ to $P_{i}$.
        $P_{i}$ defines $v_{j,i}^{b}=u_{j,i,e_{i,j}^{b}}^{b}\oplus r_{i,j}^{b}$ and sends ${v}_{j,i}^{\prime b}=\mathsf{ReRand}(pk,v_{j,i}^{b})$ to $P_{j}$. $P_{j}$ updates $c_{j}^{b}=\mathsf{ReRand}(pk,{v}_{j,i}^{\prime b})$.
      1. (b)
        For $2\leq j\leq m,1\leq b\leq B$:
        $P_{1}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{1}$ acts as receiver with input $e_{1,j}^{b}$ and $P_{j}$ acts as sender. $P_{1}$ receives $r_{1,j}^{b}=r_{j,1,e_{1,j}^{b}}^{b}\in\{0,1\}^{\lambda}$. $P_{j}$ receives $r_{j,1,0}^{b},r_{j,1,1}^{b}\in\{0,1\}^{\lambda}$. $P_{j}$ computes $u_{j,1,e_{j,1}^{b}}^{b}=r_{j,1,e_{j,1}^{b}}^{b}\oplus c_{j}^{b}$, $u_{j,1,e_{j,1}^{b}\oplus 1}^{b}=r_{j,1,e_{j,1}^{b}\oplus 1}^{b}\oplus\mathsf{Enc}(pk,\perp)$, then sends $u_{j,1,0}^{b},u_{j,1,1}^{b}$ to $P_{1}$.
        $P_{1}$ defines $\vec{\mathsf{ct}}_{1}^{\prime}\in(\{0,1\}^{\lambda})^{(m-1)B}$, and sets ${\mathsf{ct}}_{1}^{\prime(j-2)B+b}=\mathsf{ReRand}(pk,u_{j,1,e_{1,j}^{b}}^{b}\oplus r_{1,j}^{b})$.
    4. 4.
      Messages decryptions and shufflings. $P_{1}$ samples $\pi_{1}:[(m-1)B]\to[(m-1)B]$ and computes $\vec{\mathsf{ct}}_{1}^{\prime\prime}=\pi_{1}(\vec{{\mathsf{ct}}_{1}^{\prime}})$. $P_{1}$ sends $\vec{\mathsf{ct}}_{1}^{\prime\prime}$ to $P_{2}$.
      1. (a)
        For $2\leq j\leq m,1\leq i\leq(m-1)B$: $P_{j}$ computes ${\mathsf{ct}}_{j}^{i}=\mathsf{ParDec}(sk_{j},{\mathsf{ct}}_{j-1}^{\prime\prime i})$, $pk_{A_{j}}=pk_{1}\cdot\prod_{d=j+1}^{m}pk_{d}$, and ${\mathsf{ct}}_{j}^{\prime i}=\mathsf{ReRand}(pk_{A_{j}},{\mathsf{ct}}_{j}^{i})$. Then it samples $\pi_{j}:[(m-1)B]\to[(m-1)B]$ and computes $\vec{\mathsf{ct}}_{j}^{\prime\prime}={\pi_{j}}(\vec{\mathsf{ct}}_{j}^{\prime})$. If $j\neq m$, $P_{j}$ sends $\vec{\mathsf{ct}}_{j}^{\prime\prime}$ to $P_{j+1}$; else, $P_{m}$ sends $\vec{\mathsf{ct}}_{m}^{\prime\prime}$ to $P_{1}$.
      2. (b)
        For $1\leq i\leq(m-1)B$: $P_{1}$ computes $\mathsf{pt}_{i}=\mathsf{Dec}(sk_{1},{\mathsf{ct}}_{m}^{\prime\prime i})$.
    5. 5.
      Output reconstruction. $P_{1}$ sets $Y=\emptyset$. For $1\leq i\leq(m-1)B$, if $\mathsf{pt}_{i}\neq\perp$, it updates $Y=Y\cup\{\mathsf{pt}_{i}\}$. $P_{1}$ outputs $Y$.
Figure 13: Our PK-MPSU $\Pi_{\mathsf{PK\text{-}MPSU}}$
Remark 2
When instantiating our PK-MPSU framework with EC MKR-PKE, the element space has to be set as EC points accordingly, which may limit its usage in a wide range of applications. We argue that the resulting PK-MPSU protocol is still useful in scenarios that the element space is exactly EC points. We demonstrate this by building the first multi-party private ID protocol (as described in Appendix 0.G), in which $\Pi_{\mathsf{PK\text{-}MPSU}}$ for EC point elements is used as a core building block.
7 Theoretical Comparison
In this section, we compare the construction of mq-ssPMT in [LG23] (mq-ssPMT for short) and our alternative construction built on batch ssPMT (batch ssPMT for short) theoretically. Then we compare our two protocols with the recent works [LG23] (which represents the MPSU protocol with the best concrete performance) and [GNT23] (which represents the MPSU protocol with the best asymptotic complexity), respectively. We emphasize that both of the two protocols are not in the standard semi-honest model.
7.1 Comparison Between mq-ssPMT and Batch ssPMT
Liu and Gao adapt the SKE-based mq-RPMT in [ZCL+23] to construct mq-ssPMT. The most expensive part of their construction is the secret-shared oblivious decryption-then-matching (ssVODM) [ZCL+23] protocol, which is to implement a decryption circuit and a comparison circuit by the GMW protocol [GMW87]. In total, the ssVODM protocol requires $(T+l-\log n-1)n$ AND gates, where $l$ is the bit length of set elements, and $T$ is the number of AND gates in the SKE decryption circuit and is set to be considerably large ($\approx 600$) according to their paper.
The costs of the batch ssPMT of size $B=1.27n$ [^8] We use stash-less Cuckoo hashing [PSTY19] with 3 hash functions, where $B=1.27n$. consist of the costs of batch OPPRF of size $1.27n$ and the costs of $1.27n$ instances of ssPEQT, where the batch OPPRF can be instantiated by extremely fast specialized protocol and the ssPEQT is implemented by the GMW protocol. Moreover, the state-of-the-art batch OPPRF construction [CGS22, RS21, RR22] can achieve linear computation and communication complexity with respect to $n$ In the ssPEQT, the parties engage in $1.27(\gamma-1)n$ AND gates, where $\gamma$ is the output length of OPPRF. Note that in the typical setting where $n\leq 2^{24},l\leq 128,\gamma\leq 64$, we have $(T+l-\log n-1)n\gg 1.27(\gamma-1)n$, which means that the number of AND gates desired involved in batch ssPMT is far smaller than mq-ssPMT. Therefore, the construction built on batch ssPMT greatly reduces the dependency on general 2PC and significantly decreases both computation and communication complexity by a considerable factor, compared to the mq-ssPMT (For a finer-grained asymptotic analysis of mq-ssPMT and batch ssPMT, refer to Appendix 0.F.4 and Appendix 0.F.1).
7.2 Comparison Between LG Protocol and Our SK-MPSU
We provide a comprehensive complexity analysis for the state-of-the-art LG protocol in Appendix 0.F.4. In Table 2, we present a comparison of the theoretical computation and communication complexity for each party in both offline and online phases between LG protocol and our SK-MPSU.
Computation Communication
Offline Online Offline Online
Leader Client Leader Client
 [LG23] $(T+l+m)mn\log n$ $(T+l+m)mn$ $(T+l)mn$ $\lambda m^{2}n\log n$ $(T+l)mn+lm^{2}n$ $(T+l)mn$
Ours $(\gamma+m)mn\log n$ $m^{2}n$ $\lambda m^{2}n\log n$ $\gamma mn+lm^{2}n$ $(\gamma+l+m)mn$
Table 2: Asymptotic communication (bits) and computation costs of LG protocol and our SK-MPSU in the offline and online phases. $n$ is the set size. $m$ is the number of participants. $\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$. $T$ is the number of AND gate in an SKE decryption circuit, $T\approx 600$. $l$ is the bit length of input elements. $\gamma$ is the output length of OPPRF. $t$ is the noise weight in dual LPN, $t\approx 128$.
As depicted in Table 2, our offline communication complexity is comparable to theirs, our offline computation complexity and online computation and communication complexity of the leader are superior to theirs. Nevertheless, our online computation and communication complexity of client seem to be surpassed by theirs, as ours are quadratic in the number of parties $m$, while their is linear. We argue that it is because that in order to achieve security against arbitrary collusion, our protocol introduces extra overhead (from mss-ROT). Even so, for most applications when $m$ is moderate, and our online computation and communication costs still outperform theirs. This shall give the credit to our speedup by replacing their mq-ssPMT with batch ssPMT.
7.3 Comparison Between [GNT23] and Our PK-MPSU
In Table 3, we present a comparison of the theoretical computation and communication complexity for each party in both offline and online phases between  [GNT23] and our PK-MPSU.
Computation Communication
Offline Online Offline Online
 [GNT23] $\gamma mn\log n(\log n/\log\log n)$ $mn(\log n/\log\log n)$ $t\lambda m\log n(\log n/\log\log n)$ $(\gamma+\lambda)mn(\log n/\log\log n)$
Ours $\gamma mn\log n$ $mn$ $t\lambda m\log n$ $(\gamma+\lambda)mn$
Table 3: Asymptotic communication (bits) and computation costs of [GNT23] and our PK-MPSU protocol in the offline and online phases. In the offline phase, the computation is composed of symmetric-key operations; In the online phase, the computation is composed of public-key operations since we ignore symmetric-key operations. $n$ is the set size. $m$ is the number of participants. $\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$. $l$ is the bit length of input elements. $\gamma$ is the output length of OPPRF. $t$ is the noise weight in dual LPN, $t\approx 128$.
We conclude that the complexity of our PK-MPSU surpasses theirs in all respects, including the computation and communication complexity of the leader and clients in the offline and online phases, as depicted in the Table 3.
8 Performance Evaluation
In this section, we provide implementation details and experimental results for our SK-MPSU and PK-MPSU protocols. Most previous works [KS05, Fri07, BA12, GNT23] lack open-source codes. The work of [VCE22] shows fairly poor performance in comparison with the state-of-the-art LG protocol. Therefore, we only compare our works with LG protocol whose complete implementation is available on https://github.com/lx-1234/MPSU.
In the implementation of LG protocol, they designate one party to generate share correlations in the offline phase and store them as local files, so that other parties can read these files and consume these share correlations in the online phase [^9] Refer to the function ShareCorrelation::generate() in ShareCorrelationGen.cpp of the repository https://github.com/lx-1234/MPSU.git. This implementation is not faithful to the protocol specifications, which is to let the parties pairwise execute share translation protocols [CGP20]. It has two pitfalls: (1) Their code does not support distributed execution. (2) It would lead to serious information leakage. Concretely, if it is the leader who generates share correlations, then it can learn the party to whom each element in the union belongs (which is the same security problem addressed by the execution of multi-party secret shuffle as we mentioned in the technical overview). If it is a client who generates share correlations, then it can learn the union. Moreover, their implementation gives incorrect results in several test cases when the set size $n$ is quite large (cf. Table 4). As a result, their code cannot be considered as a correct or secure implementation for MPSU.
To conduct a fair comparison, we replace their implementation of share correlation generation with our correct one.
8.1 Experimental Setup
We run LG and our two protocols on Ubuntu 22.04 with a single Intel i7-13700 2.10 GHz CPU (16 physical cores) and 64 GB RAM. We emulate the two network connections using Linux tc command. In the LAN setting, the bandwidth is set to be 10 Gbps with 0.1 ms RTT latency. In the WAN setting, the bandwidth is set to be 400 Mbps with 80 ms RTT latency. We record the running time as the maximal time from protocol begin to end, including messages transmission time. We compute the communication costs of the leader as the sum of the data it sent and received. For a fair comparison, we stick to the following setting for all protocols being evaluated:
  • We set the computational security parameter $\kappa=128$ and the statistical security parameter $\lambda=40$.
  • We test the balanced scenario by setting all $m$ input sets to be of equal size. In LG protocol and our SK-MPSU, each party holds $n$ 64-bit strings. In our PK-MPSU, we assume that each party holds $n$ elements encoded as elliptic curve points (We provide two versions of the implementation of our PK-MPSU, including encoding elements in uncompressed form and in compressed form respectively).
  • For each party, we use $m-1$ threads to interact with all other parties simultaneously and 4 threads to perform share correlation generation (in LG protocol and our SK-MPSU), Beaver Triple generation (in all three), parallel SKE encryption (in LG protocol), ciphertext rerandomization and partial decryption (in our PK-MPSU).
8.2 Implementation Details
Our protocols are written in C++, and we use the following libraries in our implementation.
8.3 Experimental Results
We conduct an extensive experiment for the numbers of parties $\{3,4,5,7,9,10\}$ and a wide range of set sizes $\{2^{6},2^{8},2^{10},2^{12},2^{14},2^{16},2^{18},2^{20}\}$ in the LAN and WAN settings. We compare the performance of the protocols from four dimensions: online and total running time, and online and total communication costs. The results of online and total running time are depicted in Table 4. The results of online and total communication costs are depicted in Table 5.
Sett. $m$ Protocol Set size $n$
Online time (seconds) Total time (seconds)
$2^{6}$ $2^{8}$ $2^{10}$ $2^{12}$ $2^{14}$ $2^{16}$ $2^{18}$ $2^{20}$ $2^{6}$ $2^{8}$ $2^{10}$ $2^{12}$ $2^{14}$ $2^{16}$ $2^{18}$ $2^{20}$
LAN 3 LG23 0.030 0.038 0.056 0.106 0.286 1.024 3.899 $\times$ 0.631 0.660 0.822 1.188 4.709 22.62 95.98 $\times$
Our SK 0.004 0.005 0.006 0.012 0.033 0.138 0.749 3.641 0.232 0.254 0.282 0.344 0.824 3.042 14.32 119.2
Our PK 0.061 0.134 0.406 1.466 5.452 22.29 87.23 357.3 0.086 0.163 0.440 1.543 5.753 23.55 92.74 389.6
Our PK 0.083 0.200 0.562 2.039 8.030 31.53 127.4 514.3 0.108 0.229 0.596 2.116 8.331 32.79 133.0 546.6
4 LG23 0.050 0.055 0.063 0.117 0.336 1.274 5.572 $\times$ 0.948 0.970 1.098 1.744 6.939 32.12 135.4 $\times$
Our SK 0.006 0.007 0.008 0.017 0.049 0.202 1.043 4.905 0.442 0.458 0.489 0.665 1.700 7.167 36.16 344.71
Our PK 0.125 0.243 0.734 2.642 10.35 40.83 166.7 -- 0.161 0.280 0.784 2.760 11.00 43.51 178.3 --
Our PK 0.160 0.346 1.033 3.936 15.50 61.19 246.3 -- 0.196 0.383 1.083 4.054 16.16 63.87 257.9 --
5 LG23 0.064 0.066 0.077 0.133 0.383 1.509 7.613 -- 1.325 1.355 1.419 2.680 10.82 48.95 213.6 --
Our SK 0.009 0.010 0.013 0.021 0.061 0.290 1.478 -- 0.677 0.708 0.760 1.114 3.260 13.10 62.35 --
Our PK 0.170 0.401 1.229 4.191 16.41 64.84 262.5 -- 0.211 0.443 1.273 4.348 17.13 67.89 276.3 --
Our PK 0.192 0.509 1.611 6.226 24.62 98.04 400.0 -- 0.233 0.551 1.655 6.383 25.33 101.1 413.8 --
7 LG23 0.099 0.111 0.118 0.233 0.676 3.684 14.45 -- 2.356 2.515 2.752 5.247 23.77 103.6 452.1 --
Our SK 0.014 0.017 0.021 0.040 0.109 0.545 2.701 -- 1.312 1.357 1.510 2.336 8.368 37.17 146.0 --
Our PK 0.273 0.649 2.225 8.325 31.62 125.7 521.2 -- 0.338 0.715 2.314 8.695 33.02 131.7 552.7 --
Our PK 0.339 1.031 3.230 12.25 48.71 195.2 798.1 -- 0.404 1.097 3.319 12.62 50.11 201.3 829.6 --
9 LG23 0.136 0.142 0.18 0.49 1.584 6.121 -- -- 4.169 4.251 4.716 9.418 44.19 191.6 -- --
Our SK 0.025 0.029 0.034 0.058 0.186 0.972 -- -- 2.183 2.272 2.557 5.085 18.93 80.01 -- --
Our PK 0.429 1.064 3.553 13.23 51.74 206.8 -- -- 0.495 1.133 3.660 13.79 54.37 217.9 -- --
Our PK 0.567 1.558 5.346 22.65 80.03 322.5 -- -- 0.633 1.627 5.453 23.21 82.66 333.6 -- --
10 LG23 0.161 0.167 0.215 0.531 2.123 8.301 -- -- 4.943 5.178 5.604 12.82 59.17 252.5 -- --
Our SK 0.03 0.033 0.042 0.073 0.234 1.219 -- -- 2.745 2.883 3.290 6.733 25.12 110.7 -- --
Our PK 0.498 1.338 4.348 16.35 63.54 260.6 -- -- 0.581 1.435 4.492 17.03 66.77 275.4 -- --
Our PK 0.677 1.913 6.552 25.11 98.94 404.3 -- -- 0.760 2.010 6.696 25.79 102.2 419.1 -- --
WAN 3 LG23 4.510 4.516 4.521 4.979 6.295 11.87 17.98 $\times$ 11.87 13.04 15.00 18.33 28.49 85.85 247.6 $\times$
Our SK 2.413 2.570 2.580 3.402 4.129 5.241 10.91 31.83 8.591 9.731 12.81 16.63 25.56 56.39 192.6 826.3
Our PK 4.559 4.952 5.787 7.481 13.91 37.62 131.90 505.93 5.300 5.698 6.694 8.420 15.22 39.86 138.5 538.4
Our PK 4.571 4.662 5.643 7.796 14.82 42.08 150.5 593.6 5.312 5.408 6.550 8.735 16.13 44.33 157.1 626.1
4 LG23 5.789 5.876 6.293 6.956 7.587 13.58 23.32 $\times$ 17.10 19.64 26.64 32.29 49.43 139.9 490.9 $\times$
Our SK 3.376 3.536 3.553 4.408 5.919 8.448 20.65 64.01 11.81 14.41 20.42 27.29 48.65 132.1 499.2 2283
Our PK 6.231 7.181 8.962 11.88 23.57 69.76 251.4 -- 7.374 8.327 10.27 13.22 25.31 73.35 264.1 --
Our PK 6.264 6.436 8.256 12.14 25.92 81.50 294.2 -- 7.407 7.582 9.565 13.48 27.66 85.08 306.9 --
5 LG23 7.556 7.794 7.902 9.102 10.63 15.18 34.82 -- 23.26 26.46 35.49 47.53 83.94 250.6 969.6 --
Our SK 4.501 4.583 5.217 5.955 7.936 12.21 32.14 -- 18.11 21.85 30.93 45.68 91.47 278.9 1112 --
Our PK 8.050 9.745 11.31 16.08 35.07 108.2 388.9 -- 9.593 11.31 13.03 17.84 37.24 112.5 404.4 --
Our PK 8.072 8.869 10.91 10.84 42.83 129.1 468.9 -- 9.615 10.43 12.63 18.85 45.00 133.3 484.4 --
7 LG23 9.722 10.44 11.25 11.67 12.71 22.73 68.51 -- 34.35 45.68 65.99 95.61 203.1 707.2 2881 --
Our SK 7.241 7.414 8.121 9.808 14.52 22.93 69.19 -- 35.65 46.73 62.48 99.11 234.0 817.7 3399 --
Our PK 11.44 14.77 17.55 25.72 61.13 201.6 745.7 -- 13.79 17.14 20.10 28.30 64.28 209.3 779.2 --
Our PK 11.54 14.11 15.89 29.37 70.84 249.3 921.8 -- 13.89 16.48 18.43 31.96 73.98 257.0 955.3 --
9 LG23 11.07 11.41 13.94 14.91 15.86 34.82 -- -- 55.36 74.52 103.4 166.2 413.7 1531 -- --
Our SK 10.62 10.81 11.75 13.88 21.07 39.73 -- -- 57.91 88.27 113.7 189.5 505.3 1881 -- --
Our PK 16.30 19.45 24.87 37.99 93.69 327.8 -- -- 19.45 22.64 28.21 41.33 97.85 341.3 -- --
Our PK 15.65 19.58 22.33 42.38 109.7 396.3 -- -- 18.80 22.76 25.68 45.73 113.8 409.9 -- --
10 LG23 12.81 12.99 16.29 16.83 17.50 45.20 -- -- 66.79 91.10 137.2 126.4 579.0 2205 -- --
Our SK 12.56 12.76 13.60 16.27 25.20 49.94 -- -- 70.02 113.8 140.6 250.8 693.0 2644 -- --
Our PK 19.48 22.71 28.51 43.45 112.5 392.0 -- -- 23.06 26.32 32.28 47.28 117.2 408.5 -- --
Our PK 18.28 21.67 26.72 50.03 134.4 487.3 -- -- 21.87 25.28 30.49 53.86 139.1 503.8 -- --
[…]
m Protocol Set Size n
Online communication (MB) Total communication (MB)
$2^{6}$ $2^{8}$ $2^{10}$ $2^{12}$ $2^{14}$ $2^{16}$ $2^{18}$ $2^{20}$ $2^{6}$ $2^{8}$ $2^{10}$ $2^{12}$ $2^{14}$ $2^{16}$ $2^{18}$ $2^{20}$
3 LG23 0.157 0.284 0.962 3.662 14.43 57.58 229.8 $\times$ 6.311 7.904 13.37 32.58 114.6 474.3 2054 $\times$
Our SK 0.031 0.111 0.426 1.687 6.778 27.83 112.5 455.0 2.541 3.582 8.529 30.90 132.7 588.7 2615 11533
Our PK 1.854 2.147 3.293 7.984 26.69 102.4 406.8 1634 2.123 2.489 3.711 8.474 27.30 103.0 408.8 1640
Our PK 1.815 1.983 2.655 5.418 16.36 60.78 239.3 959.5 2.084 2.325 3.073 5.908 16.92 61.41 241.3 966.1
4 LG23 0.242 0.449 1.536 5.868 23.15 92.38 368.6 $\times$ 9.591 12.44 23.89 67.25 253.7 1074 4677 $\times$
Our SK 0.060 0.218 0.848 3.376 13.57 55.44 223.9 904.4 3.943 6.399 17.01 67.12 294.7 1315 5851 25819
Our PK 2.781 3.212 4.939 11.98 40.04 153.6 610.2 -- 3.185 3.725 5.566 12.71 40.88 154.5 613.2 --
Our PK 2.723 2.975 3.983 8.130 24.54 91.20 359.0 -- 3.127 3.488 4.610 8.127 25.38 92.12 362.0 --
5 LG23 0.330 0.630 2.173 8.325 32.86 131.2 523.5 -- 13.12 17.75 36.48 110.7 435.2 1869 8172 --
Our SK 0.097 0.361 1.411 5.628 22.63 92.18 372.0 -- 5.511 10.22 30.59 125.3 551.2 2455 10875 --
Our PK 3.708 4.283 6.585 15.97 53.38 204.8 813.6 -- 4.247 4.966 7.421 16.95 54.51 206.0 817.5 --
Our PK 3.630 3.967 5.311 10.84 32.72 121.6 478.7 -- 4.169 4.650 6.147 11.82 33.84 122.8 482.6 --
7 LG23 0.519 1.038 3.635 13.99 55.29 220.8 881.3 -- 19.92 29.90 71.59 243.2 997.9 4327 18930 --
Our SK 0.198 0.750 2.956 11.82 47.55 193.1 778.6 -- 8.853 18.92 65.23 277.2 1233 5505 24402 --
Our PK 5.563 6.425 9.878 23.95 80.07 307.1 1220 -- 6.371 7.450 11.13 25.42 81.76 309.0 1226 --
Our PK 5.445 5.950 7.966 16.25 49.07 182.3 718.0 -- 6.253 6.975 9.220 17.72 50.76 184.2 723.9 --
9 LG23 0.723 1.509 5.347 20.65 81.72 326.3 -- -- 28.15 44.74 115.0 415.3 1742 7609 -- --
Our SK 0.333 1.278 5.060 20.27 81.54 330.5 -- -- 13.90 33.27 120.9 520.0 2310 10268 -- --
Our PK 7.417 8.566 13.17 31.94 106.8 409.5 -- -- 8.495 9.933 14.84 33.90 109.0 412.0 -- --
Our PK 7.26 7.933 10.62 21.67 65.43 243.1 -- -- 8.338 9.300 12.29 23.63 67.68 245.6 -- --
10 LG23 0.831 1.768 6.296 24.36 96.43 385.1 -- -- 32.32 54.34 148.7 549.4 2314 10081 -- --
Our SK 0.413 1.594 6.322 25.34 101.9 412.9 -- -- 16.12 40.87 151.2 656.4 2921 12989 -- --
Our PK 8.344 9.637 14.82 35.93 120.1 460.7 -- -- 9.556 11.18 16.70 38.13 122.6 463.5 -- --
Our PK 8.168 8.925 11.95 24.38 73.61 273.5 -- -- 9.380 10.46 13.83 26.59 76.14 276.3 -- --
[…]
As we can see in the tables, our protocols outperform LG protocol in almost all the case studies. Especially, our SK-MPSU protocol performs best in terms of online running time in LAN setting, online communication costs for a relatively small number of parties, and total running time for medium and large sets in LAN setting among the three protocols. We observe a 4.99.3×4.9-9.3$\times$4.9 - 9.3 × improvement in online running time and a 1.78.5×1.7-8.5$\times$1.7 - 8.5 × improvement in total running time in the LAN setting. It is worth mentioning that our SK-MPSU takes only $3.6$ seconds for 3 parties with sets of $2^{20}$ items each, and $4.9$ seconds for 4 parties with sets of $2^{20}$ items each in online phase, while the implementation of LG protocol fails to run correctly in these cases.
On the other hand, our PK-MPSU protocol performs best in terms of total running time in WAN setting and total communication costs among the three protocols. It shows a 3.036.5×3.0-36.5$\times$3.0 - 36.5 × improvement in overall communication, and a 1.85.4×1.8-5.4$\times$1.8 - 5.4 × improvement in overall running time in the WAN setting, compared to LG protocol.
For the cases where both the number of parties and the set size are relatively large and the network environment is bandwidth-constrained, LG protocol marginally surpasses ours in terms of online running time, but its speedup is only up to 1.4×1.4$\times$1.4 ×. This result aligns with our theoretical analysis in Section 7.2. Nevertheless, when each party’s set size increases to sufficiently large ($2^{20}$), their implementation gives wrong results.
Appendix 0.A Leakage Analysis of [GNT23]
Gao et al. [GNT23] claimed that their MPSU protocol is secure in the presence of arbitrary colluding participants. However, we find that their protocol fails to achieve this security, and also requires the same non-collusion assumption as LG protocol. First, we give a brief review of their protocol.
Apart from MKR-PKE, the protocol of [GNT23] utilizes three primary ingredients: 1) The multi-point OPRF [PRTY19, CM20, RS21, RR22], where the sender $\mathcal{S}$ inputs a PRF key $k$, and the receiver $\mathcal{R}$ inputs $n$ elements $\{x_{1},\cdots,x_{n}\}$ and receives the corresponding PRF evaluations $\{F_{k}(x_{1}),\cdots,F_{k}(x_{n})\}$. 2) The conditional oblivious pseudorandom function (cOPRF), an extension they develop on the multi-point OPRF [^17] In their paper, the definition of cOPRF is the single-point version. For ease of illustration, we use the multi-point version., where the sender $\mathcal{S}$ additionally inputs a set $Y$. If $x_{i}\notin Y$, $\mathcal{R}$ receives $F_{k}(x_{i})$, else $\mathcal{R}$ receives a random value sampled by $\mathcal{S}$ [^18] In the latest version of [GNT23], they have made a slight adjustment to the functionality of cOPRF. However, this adjustment does not change their insecurity. Their latest protocol is still vulnerable to our attack. 3) The membership Oblivious Transfer (mOT), where $\mathcal{S}$ inputs an element $y$ and two messages $u_{0},u_{1}$, while $\mathcal{R}$ inputs a set $X$ and receives $u$, one of $u_{0},u_{1}$. If $y\notin X$, $u=u_{0}$, else $u=u_{1}$.
To illustrate the insecurity of their protocol, we consider a four-party case where the leader $P_{1}$ possesses a set $X_{1}$ of items while the remaining parties $P_{2},P_{3},P_{4}$, each possesses a single item $X_{i}$ = $\{x_{i}\}$ ($i=2,3,4$). We assume that $x_{2},x_{3},x_{4}\notin X_{1}$ and $x_{2}\neq x_{3}=x_{4}$ [^19] In their paper, to illustrate the high-level idea of the protocol, they give a four-party case as well, assuming that $x_{2}=x_{3}\notin X_{1}$ and $x_{4}\in X_{1}$, see [GNT23]..
[…]
Figure 14: Illustration of the mPSU protocol in [GNT23] for $m=4$ parties. $P_{1}$ possesses a set $X_{1}=\{x_{1}^{1},\cdots,x_{1}^{n}\}$ of $n$ items while $P_{i}$ has input set of only one item $x_{i}$ for $i=2,3,4$. We assume that $x_{2},x_{3},x_{4}\notin X_{1}$ and $x_{2}\neq x_{3}=x_{4}$. $P_{1}$ possesses two set $S,E$ and updates them in each round. In the $i$-th round, $P_{1}$ and $P_{i}$ invoke mpOPRF and mOT in sequence. $F_{k}(X_{1})$ denotes the set of PRF values on $X_{1}$, i.e., $\{F_{k}(x_{1}^{1}),\cdots,F_{k}(x_{1}^{n})\}$. $P_{i}$ invokes cOPRF and mOT with $P_{i+1},\cdots,P_{m}$ concurrently. At the end of $i$-th round, $S$ contains the multi-key PRF values on $X_{1}\cup\cdots\cup X_{i+1}$ using keys $k_{2},\cdots,k_{i+1}$, $E$ contains the MKR-PKE ciphertexts of $X_{1}\cup\cdots\cup X_{i+1}\setminus X_{1}$. Finally, all parties cooperate in a Shuffle&Decrypt protocol to decrypt $E$ in a random permutation. We denote $F_{k_{3}}(F_{k_{2}}(\cdot))$ as $F_{k_{2},k_{3}}(\cdot)$, more precisely, it is in the form of $H(k_{3},H(k_{2},\cdot\|1)\|1)$ in the paper. We emphasize that the ciphertext $\mathsf{Enc}(pk,x_{3})$ in Round 1 is different from that in Round 2. The latter is the rerandomized ciphertext of the former. We write them using the same denotation as this does not affect our illustration of the the attack. Changes to sets $S$ and $E$ at the end of each round are marked in blue font.
We demonstrate the execution process of their MPSU in Figure 14 pertaining to the scenario we describe above. In light of space constraints, Figure 14 only covers Round 1 and Round 2 and it omits some invocations of cOPRF and mOT since the potential information leakage already occurs during Round 2 when $P_{1}$ and $P_{3}$ invoke the mOT. To elaborate, $P_{1}$ receives $F_{k_{2}}(x_{2})$ from the invocation of mOT with $P_{2}$ in Round 1 and $F_{k_{2},k_{3}}(x_{3})$ from the invocation of mOT with $P_{3}$ in Round 2. If $P_{1}$ and $P_{3}$ collude, $P_{1}$ can compute $F_{k_{2},k_{3}}(x_{2})$ using $P_{3}$’s key $k_{3}$ and learn the knowledge of $F_{k_{2},k_{3}}(x_{2})\neq F_{k_{2},k_{3}}(x_{3})$, which indicates $x_{2}\neq x_{3}$. Recall that $P_{1}$ is the leader to obtain the union $X_{1}\cup\{x_{2},x_{3}=x_{4}\}$ in the end. Therefore, the coalition of $P_{1}$ and $P_{3}$ can deduce that the unique element in the union, not present in their joint sets $X_{1}\cup\{x_{3}\}$ (i.e., $x_{2}$), belongs to $P_{2}$. Accordingly, they know that $P_{4}$ possesses the duplicate element with either $P_{1}$ or $P_{3}$. Note that in the ideal world, the coalition of $P_{1}$ and $P_{3}$ should not acquire knowledge of the source party of the foreign element in the union or the party possessing the duplicate element. Therefore, their protocol at least requires the same assumption as LG protocol, assuming that the leader $P_{1}$ is not in collusion with any other parties.
Appendix 0.B Public-Key Encryption
A public-key encryption (PKE) scheme is a tuple of PPT algorithms $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ such that:
  • The key-generation algorithm $\mathsf{Gen}$ takes as input the security parameter $1^{\lambda}$ and outputs a pair of keys $(pk,sk)\in\mathcal{PK}\times\mathcal{SK}$.
  • The encryption algorithm $\mathsf{Enc}$ takes as input a public key $pk$ and a plaintext $x\in\mathcal{M}$, and outputs a ciphertext $\mathsf{ct}$.
  • The decryption algorithm $\mathsf{Dec}$ takes as input a secret key $sk$ and a ciphertext $\mathsf{ct}$, and outputs a message $x$ or or an error symbol $\perp$.
  • Correctness. For any $(pk,sk)$ outputed by $\mathsf{Gen}(1^{\lambda})$, and any $x\in\mathcal{M}$, it holds that $\mathsf{Dec}(sk,(\mathsf{Enc}(pk,x)))=x$.
The IND-CPA security of PKE implies security for encryption of multiple messages whose definition is as follows:
Definition 2
A public-key encryption scheme $\mathcal{E}=(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ has indistinguishable multiple encryptions if for all PPT adversaries $\mathcal{A}$ s.t. any tuples $(m_{1},\cdots,m_{q})$ and $({m}_{1}^{\prime},\cdots,{m}_{q}^{\prime})$ chosen by $\mathcal{A}$ (where $q$ is polynomial in $\lambda$):
[…]
$\{\mathsf{Enc}(pk,{m}_{1}^{\prime}),\cdots,\mathsf{Enc}(pk,{m}_{q}^{\prime}):(pk,sk)\leftarrow\mathsf{Gen}(1^{\lambda})\}$
Appendix 0.C Construction of Multi-Party Secret-Shared Shuffle
The functionality of share correlation generated in the offline phase is depicted in Figure 15. The online details of multi-party secret-shared shuffle in [EB22] are given in Figure 16.
  • Parameters. $m$ parties $P_{1},\cdots P_{m}$. The dimension of vector $n$. The item length $l$.
  • Functionality. On input $\pi_{i}:[n]\to[n]$ from each $P_{i}$ ($1\leq i\leq m$), sample $\vec{a}_{i}^{\prime},\vec{b}_{i}\leftarrow(\{0,1\}^{l})^{n}$ for $11\leq i\leq m-$ 1, and $\vec{a}_{i},\vec{\Delta}_{m}\leftarrow(\{0,1\}^{l})^{n}$ for $2\leq i\leq m$, which satisfy
[…]
    Give $\vec{a}_{1}^{\prime},\vec{b}_{1}$ to $P_{1}$, $\vec{a}_{i}^{\prime},\vec{a}_{i},\vec{b}_{i}$ to $P_{i}$ ($1<i<m$), and $\vec{a}_{m},\vec{\Delta}_{m}$ to $P_{m}$.
Figure 15: Share Correlation Functionality $\mathcal{F}_{\mathsf{sc}}$
  • Parameters: $m$ parties $P_{1},\cdots,P_{m}$. Ideal functionality $\mathcal{F}_{\mathsf{sc}}$ in Figure 15. The dimension of vector $n$. The item length $l$.
  • Inputs: Each party $P_{i}$ has input $\vec{x}_{i}={(x_{i}^{1},\cdots,x_{i}^{n})}$.
  • Protocol:
    1. 1.
      Each party $P_{i}$ samples a random permutation $\pi_{i}:[n]\to[n]$ and invokes $\mathcal{F}_{\mathsf{sc}}$ with input $\pi_{i}$. $P_{1}$ receives $\vec{a}_{1}^{\prime},\vec{b}_{1}\in(\{0,1\}^{l})^{n}$; For $2\leq j<m$, $P_{j}$ receives $\vec{a}_{j}^{\prime},\vec{a}_{j},\vec{b}_{j}\in(\{0,1\}^{l})^{n}$; $P_{m}$ receives $\vec{a}_{m},\vec{\Delta}_{m}\in(\{0,1\}^{l})^{n}$.
    2. 2.
      For $2\leq j\leq m$, $P_{j}$ computes $\vec{c}_{j}=\vec{x}_{j}\oplus\vec{a}_{j}$ and sends $\vec{c}_{j}$ to $P_{1}$.
    3. 3.
      $P_{1}$ computes $\vec{c^{\prime}}_{1}=\pi_{1}(\bigoplus_{j=2}^{m}\vec{c}_{j}\oplus\vec{x}_{1})\oplus\vec{a}_{1}^{\prime}$ and send it to $P_{2}$. $P_{1}$ outputs $\vec{b}_{1}$.
    4. 4.
      For $2\leq j<m$, $P_{j}$ computes $\vec{c^{\prime}}_{j}=\pi_{j}(\vec{c^{\prime}}_{j-1})\oplus\vec{a}_{j}^{\prime}$ and send it to $P_{j+1}$. $P_{j}$ outputs $\vec{b}_{j}$.
    5. 5.
      $P_{m}$ outputs $\pi_{m}(\vec{c^{\prime}}_{m-1})\oplus\vec{\Delta}_{m}$.
Figure 16: Multi-Party Secret-Shared Shuffle Protocol $\Pi_{\mathsf{ms}}$
Appendix 0.D Construction of MKR-PKE from ElGamal
The MKR-PKE is instantiated using ElGamal encryption as follows:
  • The key-generation algorithm $\mathsf{Gen}$ takes as input the security parameter $1^{\lambda}$ and generates $(\mathbb{G},g,p)$, where $\mathbb{G}$ is a cyclic group, $g$ is the generator and $q$ is the order. Outputs $sk$ and $pk=g^{sk}$.
  • The randomized encryption algorithm $\mathsf{Enc}$ takes as input a public key $pk$ and a plaintext message $x\in\mathbb{G}$, samples $r\leftarrow\mathbb{Z}_{q}$, and outputs a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})=(g^{r},x\cdot pk^{r})$.
  • The partial decryption algorithm $\mathsf{ParDec}$ takes as input a secret key share $sk$ and a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})$, and outputs $\mathsf{ct}^{\prime}=(\mathsf{ct}_{1},\mathsf{ct}_{2}\cdot\mathsf{ct}_{1}^{-sk})$.
  • The decryption algorithm $\mathsf{Dec}$ takes as input a secret key $sk$ and a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})$, and outputs $x=\mathsf{ct}_{2}\cdot\mathsf{ct}_{1}^{-sk}$.
  • The rerandomization algorithm $\mathsf{ReRand}$ takes as input $pk$ and a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})$, samples $r\leftarrow\mathbb{Z}_{q}$, and outputs $\mathsf{ct}^{\prime}=(\mathsf{ct}_{1}\cdot g^{r^{\prime}},\mathsf{ct}_{2}\cdot pk^{r^{\prime}})$.
Appendix 0.E Missing Security Proofs
0.E.1 The Proof of Theorem 5.1
  • Parameters: $m$ parties $P_{1},\cdots,P_{m}$. The message length $l$.
  • Inputs: $P_{1}$ has input $b_{1}\in\{0,1\}$ and $P_{m}$ has input $b_{m}\in\{0,1\}$.
  • Protocol:
    1. 1.
      $P_{1}$ and $P_{m}$ invoke $\mathcal{F}_{\mathsf{rot}}$ twice: First, $P_{1}$ acts as receiver with input $b_{1}$ and $P_{m}$ acts as sender. $P_{1}$ receives $r_{m,1}^{b_{1}}\in\{0,1\}^{l}$. $P_{m}$ receives $r_{m,1}^{0},r_{m,1}^{1}\in\{0,1\}^{l}$; Second, $P_{m}$ acts as receiver with input $b_{m}$ and $P_{1}$ acts as sender. $P_{m}$ receives $r_{1,m}^{b_{m}}\in\{0,1\}^{l}$. $P_{1}$ receives $r_{1,m}^{0},r_{1,m}^{1}\in\{0,1\}^{l}$.
    2. 2.
      For $1<j<m$: $P_{1}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{1}$ acts as receiver with input $b_{1}$ and $P_{j}$ as sender without input. $P_{1}$ receives $r_{j,1}^{b_{1}}\in\{0,1\}^{l}$. $P_{j}$ receives $r_{j,1}^{0},r_{j,1}^{1}\in\{0,1\}^{l}$.
    3. 3.
      For $1<j<m$: $P_{m}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{m}$ acts as receiver with input $b_{m}$ and $P_{j}$ as sender. $P_{m}$ receives $r_{j,m}^{b_{m}}\in\{0,1\}^{l}$. $P_{j}$ receives $r_{j,m}^{0},r_{j,m}^{1}\in\{0,1\}^{l}$.
    4. 4.
      For $1<j<m$: $P_{j}$ samples $\Delta_{j}\leftarrow\{0,1\}^{l}$ and computes $r_{j}=r_{j,1}^{0}\oplus r_{j,m}^{0}$. $P_{j}$ sends $u_{j,1}=\Delta_{j}\oplus r_{j,1}^{0}\oplus r_{j,1}^{1}$ to $P_{1}$, and $u_{j,m}=\Delta_{j}\oplus r_{j,m}^{0}\oplus r_{j,m}^{1}$ to $P_{m}$, then outputs $(r_{j},\Delta_{j})$.
    5. 5.
      $P_{1}$ computes $\Delta_{1}=r_{1,m}^{0}\oplus r_{1,m}^{1}$ and $r_{1}=\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1}$ ($\cdot$ denotes bitwise-AND between the repetition code of $b_{1}$ and $u_{j,1}$, which are both strings of length $l$. Similarly hereinafter), then outputs $(r_{1},\Delta_{1})$. $P_{m}$ computes $\Delta_{m}=r_{m,1}^{0}\oplus r_{m,1}^{1}$ and $r_{m}=\bigoplus_{j=2}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus r_{1,m}^{b_{m}}\oplus b_{m}\cdot\Delta_{m}$, then outputs $(r_{m},\Delta_{m})$.
Figure 17: Multi-Party Secret-Shared Random OT $\Pi_{\mathsf{mss\text{-}rot}}$
In order to facilitate the proof, we assume that $\text{ch}_{0}=1$ and $\text{ch}_{1}=m$, and $P_{1}$ and $P_{m}$ provide inputs $b_{1}\in\{0,1\}$ and $b_{m}\in\{0,1\}$ respectively. We turn to proving the correctness and security of the protocol in Figure 17. Note that when $\text{ch}_{0}$ and $\text{ch}_{1}$ are assigned different values, the proof is essentially the same.
  • Correctness. From the description of the protocol, we have the following equations:
    $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}=(\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1})$
    $\oplus(r_{j1}^{0}\oplus r_{j,m}^{0})\oplus(\bigoplus_{j=2}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus r_{m,1}^{b_{m}}\oplus b_{m}\cdot\Delta_{m})$ (1)
    $u_{j,1}=\Delta_{j}\oplus r_{j,1}^{0}\oplus r_{j,1}^{1},u_{j,m}=\Delta_{j}\oplus r_{j,m}^{0}\oplus r_{j,m}^{1}$ (2)
    From the definition of Random OT functionality (Figure 6), we have the following equations:
    $r_{j,1}^{b_{1}}=r_{j,1}^{0}\oplus b_{1}\cdot(r_{j,1}^{0}\oplus r_{j,1}^{1}),r_{j,m}^{b_{m}}=r_{j,m}^{0}\oplus b_{m}\cdot(r_{j,m}^{0}\oplus r_{j,m}^{1}),$ (3)
    $r_{m,1}^{b_{1}}=r_{m,1}^{0}\oplus b_{1}\cdot(r_{j,1}^{0}\oplus r_{j,1}^{1})=r_{m,1}^{0}\oplus b_{1}\cdot\Delta_{1},$ (4)
    $r_{1,m}^{b_{m}}=r_{1,m}^{0}\oplus b_{m}\cdot(r_{1,m}^{0}\oplus r_{1,m}^{1})=r_{1,m}^{0}\oplus b_{m}\cdot\Delta_{m}$ (5)
    Substitute Equation 2345 into Equation 1 and cancel out the same terms, we obtain:
    $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}$
    $=(b_{1}\oplus b_{m})\cdot\Delta_{1}\oplus(\bigoplus_{j=2}^{m-1}(b_{1}\oplus b_{m})\cdot\Delta_{j})\oplus(b_{1}\oplus b_{m})\cdot\Delta_{m}$
    $=\bigoplus_{i=1}^{m}(b_{1}\oplus b_{m})\cdot\Delta_{i}$
    Then we can summarize that if $b_{1}\oplus b_{m}=0$, $r_{1}=\bigoplus_{j=2}^{m}r_{j}$, else $r_{1}=\Delta_{1}\oplus(\bigoplus_{j=2}^{m}(r_{j}\oplus\Delta_{j}))$. This is exactly the functionality $\mathcal{F}_{\mathsf{mss\text{-}rot}}$.
  • Security. We now prove the security of the protocol.
    Proof
    Let $\mathcal{C}orr$ denote the set of all corrupted parties and $\mathcal{H}$ denote the set of all honest parties. $\lvert\mathcal{C}orr\rvert=t$.
    Intuitively, the protocol is secure because all things the parties do are invoking $\mathcal{F}_{\mathsf{rot}}$ and receiving random messages. The simulator can easily simulate these outputs from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages by generating random values, which are independent of honest parties’ inputs.
    To elaborate, in the case that $P_{1}\notin\mathcal{C}orr$ and $P_{m}\notin\mathcal{C}orr$, simulator receives all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$ and needs to emulate each $P_{c}$’s view, including its private $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$. The simulator for corrupted $P_{c}$ runs the protocol honestly except that it sets $P_{c}$’s random tape to be the output $\Delta_{c}$, and simulates uniform outputs from $\mathcal{F}_{\mathsf{rot}}$ under the constraint that $r_{c,1}^{0}\oplus r_{c,m}^{0}=r_{c}$. Clearly, the joint distribution of all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
    In the case that $P_{1}\in\mathcal{C}orr$ or $P_{m}\in\mathcal{C}orr$, since the protocol is symmetric with respect to the roles of $P_{1}$ and $P_{m}$, we focus on the case of corrupted $P_{1}$. The simulator receives $b_{1}$ in addition to all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$. For $P_{1}$, its view consists of its private $\Delta_{1}$, outputs $r_{m,1}^{b_{1}}$, $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j<m}$ from $P_{j}$. For each $P_{c}(c\neq 1)$, its view consists of its private $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$.
    The simulator sets each corrupted party $P_{c}$’s random tape to be its output $\Delta_{c}$. Then for $P_{1}$’s view, it runs the protocol honestly except that it simulates uniform outputs $r_{m,1}^{b_{1}}$, $(r_{1,m}^{0},r_{1,m}^{1})$, $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$. For other corrupted parties’ view, it runs the protocol honestly except that it sets $r_{c,m}^{0}=r_{c,1}^{0}\oplus r_{c}$ and simulates uniform output $r_{c,m}^{1}$ from $\mathcal{F}_{\mathsf{rot}}$.
    In the real execution, $P_{1}$ receives $u_{i,1}=\Delta_{i}\oplus r_{i,1}^{0}\oplus r_{i,1}^{1}$. From the definition of ROT functionality, $r_{i,1}^{0}$ (when $b_{1}=0$) or $r_{i,1}^{1}$(when $b_{1}=0$) is uniform and independent of $P_{1}$’s view. Therefore, $u_{i,1}$ is uniformly at random from the perspective of $P_{1}$. Clearly, the joint distribution of $b_{1}$ and all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
    In the case that $P_{1}\in\mathcal{C}orr$ and $P_{m}\in\mathcal{C}orr$, the simulator receives $b_{1},b_{m}$ and all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$. For $P_{1}$, its view consists of its private $\Delta_{1}$, outputs $r_{m,1}^{b_{1}}$, $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j<m}$ from $P_{j}$. For each $P_{c}(c\neq 1,c\neq m)$, its view consists of its private $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$. For $P_{m}$, its view consists of its private $\Delta_{m}$, outputs $r_{1,m}^{b_{m}}$, $(r_{m,1}^{0},r_{m,1}^{1})$, $\{r_{j,1}^{b_{m}}\}_{1<j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,m}\}_{1<j<m}$ from $P_{j}$.
    The simulator sets each corrupted party $P_{c}$’s random tape to be its output $\Delta_{c}$. Then for $P_{1}$’s view, it runs the protocol honestly except that it simulates uniform outputs $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$. For the view of $P_{c}(c\neq 1,c\neq m)$, it runs the protocol honestly except that it sets $r_{c,m}^{0}=r_{c,1}^{0}\oplus r_{c}$ and simulates uniform output $r_{c,m}^{1}$ from $\mathcal{F}_{\mathsf{rot}}$. For $P_{m}$’s view, it runs the protocol honestly with the following changes:
    • It simulates uniform outputs $r_{i,m}^{b_{m}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniform messages $u_{i,m}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus r_{1,m}^{b_{m}}\oplus b_{m}\cdot\Delta_{m}=r_{m}$, where $P_{i}\in\mathcal{H}$.
    • It sets the output $r_{c,m}^{b_{m}}$ from $\mathcal{F}_{\mathsf{rot}}$ to be consistent with the partial view $(r_{c,m}^{0},r_{c,m}^{1})$ of each corrupted $P_{c}$ in preceding simulation, where $c\neq 1$ and $c\neq m$.
    Clearly, the joint distribution of $b_{1},b_{m}$ and all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
0.E.2 The Proof of Theorem 5.2
Proof
This proof is supposed to be divided into two cases in terms of whether $P_{1}\in\mathcal{C}orr$, since this determines whether the adversary has knowledge of the output. Nevertheless, the simulation of these two cases merely differ in the output reconstruction stage, thus we combine them together for the sake of simplicity. Specifically, the simulator receives the input $X_{c}$ of $P_{c}\in\mathcal{C}orr$ and the output $\bigcup_{i=1}^{m}X_{i}$ if $P_{1}\in\mathcal{C}orr$.
For each $P_{c}$, its view consists of its input $X_{c}$, outputs from $\mathcal{F}_{\mathsf{bssPMT}}$, $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, output $\vec{sh}_{c}^{\prime}$ from $\mathcal{F}_{\mathsf{ms}}$ as its share, and $m-1$ sets of shares $\{\vec{sh}_{i}^{\prime}\}_{1<i\leq m}$($P_{i}$’s output from $\mathcal{F}_{\mathsf{ms}}$) from $P_{i}$ if $c=1$. The simulator emulates each $P_{c}$’s view by running the protocol honestly with the following changes:
  • In step 2, it simulates uniform outputs $\{e_{c,j}^{b}\}_{c<j\leq m}$ and $\{e_{c,i}^{b}\}_{1\leq i<c}$ from $\mathcal{F}_{\mathsf{bssPMT}}$, on condition that $P_{i},P_{j}\in\mathcal{H}$.
  • In step 3, it simulates uniform outputs $\{r_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j},\{\Delta_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, on condition that $\exists 1\leq d\leq j,P_{d}\in\mathcal{H}$.
  • In step 4, it simulates uniformly output $\vec{sh}_{c}^{\prime}$ from $\mathcal{F}_{\mathsf{ms}}$.
Now we discuss the case when $P_{1}\in\mathcal{C}orr$. In step 4 and 5, it computes $Y=\bigcup_{i=1}^{m}X_{i}\setminus X_{1}$ and constructs $\vec{v}\in(\{0,1\}^{l+\kappa})^{(m-1)B}$ as follows:
  • For $\forall x_{i}\in Y$, $v_{i}=x_{i}\|\mathsf{H}(x)$, $1\leq i\leq\lvert Y\rvert$.
  • For $\lvert Y\rvert<i\leq(m-1)B$, samples $v_{i}\leftarrow\{0,1\}^{l+\kappa}$.
Then it samples a random permutation $\pi:[(m-1)B]\to[(m-1)B]$ and computes $\vec{v^{\prime}}=\pi(\vec{v})$. For $1\leq i\leq m$, it samples share $\vec{sh}_{i}^{\prime}\leftarrow(\{0,1\}^{l+\kappa})^{(m-1)B}$, which satisfies $\bigoplus_{i=1}^{m}\vec{sh}_{i}^{\prime}=\vec{v^{\prime}}$ and is consistent with the previous sampled $\vec{sh}_{c}^{\prime}$ for each corrupted $P_{c}$. Add all $\vec{sh}_{i}^{\prime}$ to $P_{1}$’s view and $\vec{sh}_{c^{\prime}}^{\prime}$ to each corrupted $P_{c^{\prime}}$’s view($c^{\prime}\neq 1$) as its output from $\mathcal{F}_{\mathsf{ms}}$, respectively.
The changes of outputs from $\mathcal{F}_{\mathsf{bssPMT}}$ and $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ have no impact on $P_{c}$’s view, for the following reasons. By the definition of $\mathcal{F}_{\mathsf{bssPMT}}$, each output $e_{c,j}^{b}$ and $e_{c,i}^{b}$ from $\mathcal{F}_{\mathsf{bssPMT}}$ is uniformly distributed as a secret-share between $P_{c}$ and $P_{j}$, or $P_{i}$ and $P_{c}$, where $P_{i},P_{j}\in\mathcal{H}$. By the definition of $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, each output $r_{c,ji}^{b}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ is a secret-share of 00 among $P_{1},\cdots,P_{j}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=0$, or a secret-share of $\bigoplus_{d=1}^{j}\Delta_{d,ji}^{b}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=1$. Therefore, even if $P_{c}$ colludes with others, $r_{c,ji}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{d}\in\mathcal{H}(1\leq d\leq j)$ holding one share. Besides, the outputs $\{\Delta_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ are uniformly distributed.
It remains to demonstrate that the output $\vec{sh}_{c}^{\prime}$ from $\mathcal{F}_{\mathsf{ms}}$($P_{1}\notin\mathcal{C}orr$) or all outputs $\{\vec{sh}_{i}^{\prime}\}_{1\leq i\leq m}$ from $\mathcal{F}_{\mathsf{ms}}$($P_{1}\in\mathcal{C}orr$) does not leak any other information except for the union. The former case is easier to tackle with. The output $\vec{sh}_{c}^{\prime}$ is distributed as a secret-share among all parties, so it is uniformly distributed from the perspective of adversary.
We now proceed to explain the latter case. For all $1<j\leq m$, consider an element $x\in X_{j}$ and $x$ is placed in the $b$th bin by $P_{j}$. In the real protocol, if there is no $X_{i}(1\leq i<j)$ s.t. $x\in X_{i}$, then for all $1\leq i<j$, $e_{i,j}^{y}\oplus e_{j,i}^{y}=0$. By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 10, each $r_{d,ji}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=1}^{j}r_{d,ji}^{b}=0$. From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=1}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)$, namely, they are additive shares of $x\|\mathsf{H}(x)$ among all parties. This is exactly identical to the simulation.
If there exists some $X_{i}(1\leq i<j)$ s.t. $x\in X_{i}$, then $e_{i,j}^{b}\oplus e_{j,i}^{b}=1$. By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 10, each $r_{d,ji}^{b}$ is uniform conditioned on $\bigoplus_{d=1}^{j}r_{d,ji}^{b}=\bigoplus_{d=1}^{j}\Delta_{d,ji}^{b}$, where each $\Delta_{d,ji}^{b}$ is uniformly held by $P_{d}$. From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform conditioned on $\bigoplus_{d=1}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)\oplus\bigoplus_{d=1}^{j}\Delta_{d,ji}^{b}\oplus r$, where $r$ is the sum of remaining terms. Then, even if $P_{1}$ colludes with others, $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{d}\in\mathcal{H}(1\leq d\leq j)$ holding one uniform $\Delta_{d,ji}^{b}$ and independent of all honest parties’ inputs. For all empty bins, $u_{d,j}^{b}$ is chosen uniformly random, so the corresponding $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ is also uniformly at random, which is identical to the simulation. By the definition of $\mathcal{F}_{\mathsf{ms}}$, all parties additively share $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ in a random permutation that maintains privacy against a coalition of arbitrary corrupted parties, and receive back $\{\vec{sh}_{i}^{\prime}\}$, respectively. We conclude that all outputs $\{\vec{sh}_{i}^{\prime}\}_{1\leq i\leq m}$ from $\mathcal{F}_{\mathsf{ms}}$ distribute identically between the real and ideal executions.
0.E.3 The Proof of Theorem 6.1
Proof
The simulator receives the input $X_{c}$ of $P_{c}\in\mathcal{C}orr$ and the output $\bigcup_{i=1}^{m}X_{i}$ if $P_{1}\in\mathcal{C}orr$.
For each $P_{c}$, its view consists of its input $X_{c}$, outputs from $\mathcal{F}_{\mathsf{bssPMT}}$ and $\mathcal{F}_{\mathsf{rot}}$, protocol messages $\{u_{j,c,0}^{b}\}_{c<j\leq m},\{u_{j,c,1}^{b}\}_{c<j\leq m}$ from $P_{j}$, rerandomization messages $\{{v}_{c,i}^{\prime b}\}_{1<i<c}$ from $P_{i}$, $\pi_{c}$, permuted partial decryption messages $\vec{\mathsf{ct}}_{c-1}^{\prime\prime}$ from $P_{c-1}$ if $c>1$, or $\vec{\mathsf{ct}}_{m}^{\prime\prime}$ from $P_{m}$ if $c=1$. The simulator emulates each $P_{c}$’s view by running the protocol honestly with the following changes:
  • In step 2, it simulates uniform outputs $\{e_{c,j}^{b}\}_{c<j\leq m}$ and $\{e_{c,i}^{b}\}_{1\leq i<c}$ from $\mathcal{F}_{\mathsf{bssPMT}}$, on condition that $P_{i},P_{j}\in\mathcal{H}$.
  • In step 3, it simulates uniform outputs $\{r_{c,i,0}^{b}\}_{1\leq i<c}$, $\{r_{c,i,1}^{b}\}_{1\leq i<c}$ from $\mathcal{F}_{\mathsf{rot}}$, and $\{r_{j,c,e_{c,j}^{b}}^{b}\}_{c<j\leq m}$ from $\mathcal{F}_{\mathsf{rot}}$, on condition that $P_{i},P_{j}\in\mathcal{H}$. For $c<j\leq m$, it computes $u_{j,c,e_{c,j}^{b}}^{b}=r_{j,c,e_{c,j}^{b}}^{b}\oplus\mathsf{Enc}(pk,\perp)$ and simulates $u_{j,c,e_{c,j}^{b}\oplus 1}^{b}$ uniformly at random, on condition that $P_{j}\in\mathcal{H}$. For $1<i<c$, it simulates ${v}_{c,i}^{\prime b}=\mathsf{Enc}(pk,\perp)$, on condition that $P_{i}\in\mathcal{H}$.
  • If $P_{1}\notin\mathcal{C}orr$, in step 4, for $1\leq i\leq(m-1)B$, it computes ${\mathsf{ct}}_{c-1}^{\prime i}=\mathsf{Enc}(pk,\perp)$, and then simulates the vector $\vec{\mathsf{ct}}_{c-1}^{\prime\prime}={\pi}(\vec{\mathsf{ct}}_{c-1}^{\prime})$ from $P_{c-1}$, where $\pi$ is sampled uniformly random and $P_{c-1}\in\mathcal{H}$.
Now we discuss the case when $P_{1}\in\mathcal{C}orr$. In step 4, assume $d$ is the largest number that $P_{d}\in\mathcal{H}$, namely, $P_{d+1},\cdots,P_{m}\in\mathcal{C}orr$. The simulator emulates the partial decryption messages $\vec{\mathsf{ct}}_{d}^{\prime\prime}$ from $P_{d}$ in the view of $P_{d+1}$ as follows:
  • For $\forall x_{i}\in Y=\bigcup_{j=1}^{m}X_{j}$, ${\mathsf{ct}}_{d}^{\prime i}=\mathsf{Enc}(pk_{A},x_{i})$, $1\leq i\leq\lvert Y\rvert$.
  • For $\lvert Y\rvert<i\leq(m-1)B$, sets ${\mathsf{ct}}_{d}^{\prime i}=\mathsf{Enc}(pk_{A},\perp)$.
where $pk_{A_{d}}=pk_{1}\cdot\prod_{j=d+1}^{m}pk_{j}$. Then it samples a random permutation $\pi:[(m-1)B]\to[(m-1)B]$ and computes $\vec{\mathsf{ct}}_{d}^{\prime\prime}=\pi(\vec{\mathsf{ct}}_{d}^{\prime})$.
For other corrupted $P_{d^{\prime}+1}\in\{P_{2},\cdots,P_{d-1}\}$, if $P_{d^{\prime}}\in\mathcal{H}$, it simulates each partial decryption message ${\mathsf{ct}}_{d^{\prime}}^{\prime i}=\mathsf{Enc}(pk,\perp)$ for $1\leq i\leq(m-1)B$, and then computes the vector $\vec{\mathsf{ct}}_{d^{\prime}}^{\prime\prime}={\pi}(\vec{\mathsf{ct}}_{d^{\prime}}^{\prime})$ from $P_{d^{\prime}}$, where $\pi_{d^{\prime}}$ is sampled uniformly random. Append $\vec{\mathsf{ct}}_{d^{\prime}}^{\prime\prime}$ to the view of $P_{d^{\prime}+1}$.
The changes of outputs from $\mathcal{F}_{\mathsf{bssPMT}}$ and $\mathcal{F}_{\mathsf{rot}}$ have no impact on $P_{c}$’s view, for similar reasons in Theorem 5.2.
Indeed, $u_{j,c,e_{c,j}^{b}\oplus 1}^{b}$ is uniform in the real process, as $r_{j,c,e_{c,j}^{b}\oplus 1}^{b}$ (which is one of $P_{j}$’s output from $\mathcal{F}_{\mathsf{rot}}$ hidden from $P_{c}$, and is used to mask the encrypted message in $u_{j,c,e_{c,j}^{b}\oplus 1}^{b}$) is uniform and independent of $r_{j,c,e_{c,j}^{b}}^{b}$ from $P_{c}$’s perspective.
It’s evident from the descriptions of the protocol and the simulation that the simulated $u_{j,c,e_{c,j}^{b}}^{b}$ is identically distributed to that in the real process, conditioned on the event $e_{c,j}^{b}\oplus e_{j,c}^{b}=1$. The analysis in the case of $e_{c,j}^{b}\oplus e_{j,c}^{b}=0$ can be further divided into two subcases, $c\neq 1$ and $c=1$. We first argue that when $c\neq 1$, $u_{j,c,e_{c,j}^{b}}^{b}$ emulated by simulator is indistinguishable from that in the real process.
In the real process, for $1<c<j\leq m,1\leq b\leq B$, if $\mathsf{Elem}(\mathcal{C}_{j}^{b})\in X_{j}\setminus(X_{2}\cup\cdots\cup X_{c-1})$, $c_{j}^{b}=\mathsf{Enc}(pk,\mathsf{Elem}(\mathcal{C}_{j}^{b}))$, $u_{j,c,e_{c,j}^{b}}^{b}=r_{j,c,e_{c,j}^{b}}^{b}\oplus\mathsf{Enc}(pk,\mathsf{Elem}(\mathcal{C}_{j}^{b}))$; else $c_{j}^{b}=\mathsf{Enc}(pk,\perp)$, $u_{j,c,e_{c,j}^{b}}^{b}=r_{j,c,e_{c,j}^{b}}^{b}\oplus\mathsf{Enc}(pk,\perp)$. In the real process, for $1<c<j\leq m,1\leq b\leq B$, $u_{j,c,e_{c,j}^{b}}^{b}=r_{j,c,e_{c,j}^{b}}^{b}\oplus\mathsf{Enc}(pk,\perp)$. If there exists an algorithm that distinguishes these two process, it implies the existence of an algorithm that can distinguish two lists of encrypted messages, with no knowledge of $sk$ (since $sk$ is secret-shared among $m$ parties, it is uniformly distributed for any coalition of $m-1$ parties). Consequently, this implies the existence of a adversary to break the indistinguishable multiple encryptions of $\mathcal{E}$ in Definition 2 (where $q=(m-1)B$).
When $c=1$, $u_{j,e_{1,j}^{b}}^{b}$ emulated by simulator is indistinguishable from that in the real process for the similar reason as the above analysis when $c>1$.
Next, we start demonstrating that all ${v}_{c,i}^{\prime b}=\mathsf{Enc}(pk,\perp)$ emulated by simulator are indistinguishable from the real ones via the sequences of hybrids:
  • $\mathsf{Hyb}_{0}$ The real interaction. For $1<i<c,1\leq b\leq B$: If $\mathsf{Elem}(\mathcal{C}_{c}^{b})\in X_{c}\setminus(X_{1}\cup\cdots\cup X_{i})$, ${v}_{c,i}^{b}=\mathsf{Enc}(pk,\mathsf{Elem}(\mathcal{C}_{c}^{b}))$; else ${v}_{c,i}^{b}=\mathsf{Enc}(pk,\perp)$. ${v}_{c,i}^{\prime b}=\mathsf{ReRand}(pk,{v}_{c,i}^{b})$.
  • $\mathsf{Hyb}_{1}$ For $1<i<c,1\leq b\leq B$: If $\mathsf{Elem}(\mathcal{C}_{c}^{b})\in X_{c}\setminus(X_{1}\cup\cdots\cup X_{i})$, ${v}_{c,i}^{\prime b}=\mathsf{Enc}(pk,\mathsf{Elem}(\mathcal{C}_{c}^{b}))$; else ${v}_{c,i}^{\prime b}=\mathsf{Enc}(pk,\perp)$. This change is indistinguishable by the rerandomizable property of $\mathcal{E}$.
  • $\mathsf{Hyb}_{2}$ For $1<i<c,1\leq b\leq B$: ${v}_{c,i}^{\prime b}=\mathsf{Enc}(pk,\perp)$. This change is indistinguishable by the indistinguishable multiple encryptions of $\mathcal{E}$.
When $P_{1}\in\mathcal{C}orr$, we prove that $\vec{\mathsf{ct}}_{d}^{\prime\prime}$ emulated by simulator is indistinguishable from that in the real process via the sequences of hybrids:
  • $\mathsf{Hyb}_{0}$ The real interaction. $\vec{\mathsf{ct}}_{1}^{\prime\prime}={\pi_{1}}(\vec{\mathsf{ct}}_{1}^{\prime})$. For $2\leq j\leq d,1\leq i\leq(m-1)B$: ${\mathsf{ct}}_{j}^{i}=\mathsf{ParDec}(sk_{j},{\mathsf{ct}}_{j-1}^{\prime\prime i}),{\mathsf{ct}}_{j}^{\prime i}=\mathsf{ReRand}(pk_{A_{j}},{\mathsf{ct}}_{j}^{i})$, $\vec{\mathsf{ct}}_{j}^{\prime\prime}={\pi_{j}}(\vec{\mathsf{ct}}_{j}^{\prime})$.
  • $\mathsf{Hyb}_{1}$ For $2\leq j\leq d,1\leq i\leq(m-1)B$: ${\mathsf{ct}}_{j}^{i}=\mathsf{ParDec}(sk_{j},{\mathsf{ct}}_{j-1}^{i})$. $\vec{\mathsf{ct}}_{d}^{\prime}=\mathsf{ReRand}(pk_{A_{d}},\vec{\mathsf{ct}}_{d})$, $\vec{\mathsf{ct}}_{d}^{\prime\prime}={\pi}(\vec{\mathsf{ct}}_{d}^{\prime})$, where $\pi=\pi_{1}\circ\cdots\circ\pi_{d}$. It’s easy to see that $\mathsf{Hyb}_{1}$ is identical to $\mathsf{Hyb}_{0}$.
  • $\mathsf{Hyb}_{2}$ $\vec{\mathsf{ct}}_{1}$ is replaced by the following:
    • For $\forall x_{i}\in Y=\bigcup_{j=1}^{m}X_{j}$, $\mathsf{ct}_{1}^{i}=\mathsf{Enc}(pk,x_{i})$, $1\leq i\leq\lvert Y\rvert$.
    • For $\lvert Y\rvert<i\leq(m-1)B$, sets $\mathsf{ct}_{1}^{i}=\mathsf{Enc}(pk,\perp)$.
    $\mathsf{Hyb}_{2}$ rearranges $\vec{\mathsf{ct}}_{1}$ and it is identical to $\mathsf{Hyb}_{1}$ as the adversary is unaware of $\pi_{d}$ s.t. the order of elements in $\vec{\mathsf{ct}}_{1}$ has no effect on the result of $\vec{\mathsf{ct}}_{d}^{\prime\prime}$.
  • $\mathsf{Hyb}_{3}$ $\vec{\mathsf{ct}}_{d}$ is replaced by the following:
    • For $\forall x_{i}\in Y=\bigcup_{j=1}^{m}X_{j}$, $\mathsf{ct}_{d}^{i}=\mathsf{Enc}(pk_{A_{d}},x_{i})$, $1\leq i\leq\lvert Y\rvert$.
    • For $\lvert Y\rvert<i\leq(m-1)B$, sets $\mathsf{ct}_{d}^{i}=\mathsf{Enc}(pk_{A_{d}},\perp)$.
    The indistinguishability between $\mathsf{Hyb}_{3}$ and $\mathsf{Hyb}_{2}$ is implied by the partially decryptable property of $\mathcal{E}$.
  • $\mathsf{Hyb}_{4}$ $\vec{\mathsf{ct}}_{d}^{\prime}$ is replaced by the following:
    • For $\forall x_{i}\in Y=\bigcup_{j=1}^{m}X_{j}$, ${\mathsf{ct}}_{d}^{\prime i}=\mathsf{Enc}(pk_{A_{d}},x_{i})$, $1\leq i\leq\lvert Y\rvert$.
    • For $\lvert Y\rvert<i\leq(m-1)B$, sets ${\mathsf{ct}}_{d}^{\prime i}=\mathsf{Enc}(pk_{A_{d}},\perp)$.
    $\mathsf{Hyb}_{4}$ is indistinguishable to $\mathsf{Hyb}_{3}$ because of the rerandomizable property of $\mathcal{E}$.
  • $\mathsf{Hyb}_{5}$ The only change in $\mathsf{Hyb}_{5}$ is that $\pi$ are sampled uniformly by the simulator. $\mathsf{Hyb}_{5}$ generates the same $\vec{\mathsf{ct}}_{d}^{\prime\prime}$ as in the simulation. Given that $\pi_{d}$ is uniform in the adversary’s perspective, the same holds for $\pi$, so $\mathsf{Hyb}_{5}$ is identical to $\mathsf{Hyb}_{4}$.
When $P_{1}\notin\mathcal{C}orr$, the simulator is unaware of the final union, so it has to emulate the partial decryption messages $\vec{\mathsf{ct}}_{c-1}^{\prime\prime}$ as permuted $\mathsf{Enc}(pk,\perp)$ in the view of $P_{c}$. Compared to the above hybrid argument, we only need to add one additional hybrid after $\mathsf{Hyb}_{4}$ to replace all rerandomized partial decryption messages $\vec{\mathsf{ct}}_{c-1}^{\prime}$ with $\mathsf{Enc}(pk,\perp)$. This change is indistinguishable by the indistinguishable multiple encryptions of $\mathcal{E}$, as the adversary cannot distinguish two lists of messages with no knowledge of the partial secret key $sk_{A_{c-1}}=sk_{1}+sk_{c}+\cdots+sk_{m}$ (it is unaware of $sk_{1}$).
The same applies for the simulation of the partial decryption messages $\vec{\mathsf{ct}}_{d^{\prime}}^{\prime\prime}$ in the view of corrupted $P_{d^{\prime}+1}$ ($d^{\prime}\neq d$) when $P_{1}\in\mathcal{C}orr$. $\vec{\mathsf{ct}}_{d^{\prime}}^{\prime\prime}$ is also emulated by permuted $\mathsf{Enc}(pk,\perp)$. To avoid repetition, we omit the analysis here.
Appendix 0.F Complete Complexity Analyses of Protocols
0.F.1 Theoretical Analysis of Protocol $\Pi_{\mathsf{bssPMT}}$
The costs of each stage in $\Pi_{\mathsf{bssPMT}}$ (Figure 9) are calculated as follows.
  • Batch OPPRF. The cost of batch OPPRF mainly consists of two parts:
    • Batch OPRF: There are several options for instantiating batch OPRF functionality [KKRT16, BC23]. We opt for subfield vector oblivious linear evaluation (subfield-VOLE) [BCG+19a, BCG+19b, RRT23] to instantiate batch OPRF using the approach in [BC23].
      For $B=O(n)$ instances of OPRF, we require performing a subfield-VOLE of size $B$ in the offline phase, and sending $B$ derandomization messages of length $l$ in the online phase. We resort to Dual LPN with Fixed Weight and Regular Noise [BCG+19a, BCG+19b] to improve the efficiency of subfield-VOLE, and instantiate the code family with Expand-Convolute codes presented by [RRT23], which enables near-linear time syndrome computation. The computation complexity is $O(n\log n)$. The computation complexity is $O(t\lambda\log n/t)$, where $t$ is the fixed noise weight [^20] For instance, $t\approx 128$.. The round complexity is $O(1)$.
    • OKVS: The size of key-value pairs encoded into OKVS is $\lvert K_{1}\rvert+\cdots+\lvert K_{B}\rvert$. Taking into account the invocation of batch ssPMT in our two MPSU protocols in advance, this size is $3n$. We use the OKVS construction of [RR22], and the computation complexities of $\mathsf{Encode_{H}}$ and $\mathsf{Decode_{H}}$ algorithms are both $O(n)$. As we employ their $w=3$ scheme with a cluster size of $2^{14}$, the size of OKVS is $1.28\gamma\cdot(3n)$ bits, where $\gamma$ is the output length of OPPRF and is set as $\sigma+\log_{2}B$ to avoid wrong matching.
  • Offline phase. The computation complexity of each party is $O(n\log n)$. The communication complexity of each party is $O(t\lambda\log(n/t))$. The round complexity is $O(1)$.
  • Online phase. The computation complexity of each party is $O(n)$. The communication complexity of each party is $O(\gamma n)$. The round complexity is $O(1)$.
  • Secret-shared private equality tests. Like [PSTY19], we instantiate ssPEQT (Figure 5) using the generic MPC techniques. The circuit of $\mathcal{F}_{\mathsf{ssPEQT}}$ is composed of $\gamma-1$ AND gates in GMW [GMW87], where the inputs are already in the form of secret-shaing. Executing $\gamma-1$ AND gates in sequence would incur $\gamma-1$ rounds. To reduce the round complexity, we leverage a divide-and-conquer strategy and recursively organize the AND gates within a binary tree structure, where each layer requires one round. This ensures that the number of rounds is directly related to the depth of the tree (i.e., $O(\log{\gamma})$). To sum up, each party invoke $B$ instances of ssPEQT, which amounts to $(\gamma-1)B$ AND gates and takes $O(\log{\gamma})$ rounds.
    We use silent OT extension [BCG+19a] to generate Beaver triples in offline phase, then each AND gate only requires 4 bits communication and $O(1)$ computation in the online phase. As a result, an invocation of ssPEQT requires $4\gamma$ bits communication and $O(1)$ computation in the online phase [^21] When calculating computational complexity, one evaluation on PRG or one hash operation is usually considered as $O(1)$ operation. However, here, the computational unit is bitwise XOR operation, and it’s evident that performing $O(\gamma)$ bitwise XOR operations is much faster than executing $O(1)$ PRG evaluation or hash operation. Therefore, we count the online computation complexity of ssPEQT as $O(1)$.
    • Offline phase. The computation complexity of each party is $O(\gamma n\log n)$. The communication complexity of each party is $O(t\lambda\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(n)$. The communication complexity of each party is $O(\gamma n)$. The round complexity is $O(\log\gamma)$.
  • 0.F.2 Theoretical Analysis of Protocol $\Pi_{\mathsf{SK\text{-}MPSU}}$
    The costs of each stage in $\Pi_{\mathsf{SK\text{-}MPSU}}$ (Figure 12) are calculated as follows.
  • Batch ssPMT. To achieve linear communication of this stage, we use stash-less Cuckoo hashing [PSTY19]. To render the failure probability (failure is defined as the event where an item cannot be stored in the table and must be stored in the stash) less than $2^{-40}$, we set $B=1.27n=O(n)$ for 3-hash Cuckoo hashing. The cost of batch ssPMT follows the complexity analysis in Section 4. For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke batch ssPMT. Overall, each party $P_{j}$ engages in $m-1$ invocations of batch ssPMT, acting as $\mathcal{R}$ in the first $j-1$ invocations, and acting as $\mathcal{S}$ in the last $m-j$ invocations.
    • Offline phase. The computation complexity of each party is $O(\gamma mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(mn)$. The communication complexity of each party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
  • Multi-party secret-shared random oblivious transfers. Each invocation of mss-ROT involves pairwise executions of two-party OT. If $P_{i}$ and $P_{j}$ hold the two choice bits, then $P_{i}$ and $P_{j}$ invoke two instances of OT and they separately invoke one instance of OT with the remaining parties.
    Each OT execution consists of two parts: In the offline phase, the parties engages in a random-choice-bit ROT, then $\mathcal{S}$ sends one messages of length $l+\kappa$ ($\kappa=\sigma+\log(m-1)+\log n$) to $\mathcal{R}$ [^22] The OT execution between two choice-bit-holder does not include the sending step. This does not affect the total complexity, so we ignore this.; In the online phase, $\mathcal{R}$ sends a 1-bit derandomization message to transform ROT into a chosen-input version.
    For $1\leq i<j\leq m$, $P_{1},\cdots,P_{j}$ engage in $B$ instances of mss-ROT, where $P_{i}$ and $P_{j}$ hold the two choice bits. Considering one instance of mss-ROT and fixing $j$, there are overall $j-1$ invocations of mss-ROT. For $1\leq d<j$, each $P_{d}$ has one chance to act as the choice-bit-holder (who invokes two instances of OT with another and one with $j-2$ remaining parties separately in one invocation of mss-ROT) and acts as a nomal party (who invokes one instances of OT with the two choice-bit-holders separately in one invocation of mss-ROT) in the rest $j-2$ invocations. $P_{j}$ act as the choice-bit-holder in $j-1$ invocations (who invokes $2+(j-2)=j$ instances of OT).
    Overall, each party $P_{j^{\prime}}$ invokes $[{j^{\prime}}({j^{\prime}}-1)+\sum_{j={j^{\prime}}+1}^{m}(j+2(j-2))]B=(\frac{3m^{2}-{j^{\prime}}^{2}-5m+3{j^{\prime}}}{2})B$ instances of two-party OT.
    • Offline phase. The computation complexity of each party is $O(m^{2}n\log n)$. The communication complexity of each party is $O(t\lambda m^{2}\log(n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(m^{2}n)$. The communication complexity of each party is $O(m^{2}n)$. The round complexity is $O(1)$.
  • Multi-party secret-shared shuffle. We use the construction in [EB22]. In the offline phase, each pair of parties runs a Share Translation protocol [CGP20] of size $(m-1)B$ and length $l+\kappa$.
    • Offline phase. The computation complexity of each party is $O(m^{2}n\log(mn))$. The communication complexity of each party is $O(\lambda m^{2}n\log(mn))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of $P_{1}$ is $O(m^{2}n)$. The communication complexity of $P_{1}$ is $O((l+\kappa)m^{2}n)$. The computation complexity of $P_{j}$ is $O(mn)$. The communication complexity of $P_{j}$ is $O((l+\kappa)mn)$. The round complexity is $O(m)$.
  • Output reconstruction. For $1<j\leq m$, $P_{j}$ sends $\vec{sh}_{j}^{\prime}\in(\{0,1\}^{l+\kappa})^{(m-1)B}$ to $P_{1}$. $P_{1}$ reconstructs $(m-1)B$ secrets, each having $m$ shares.
    • Online phase. The computation complexity of $P_{1}$ is $O(m^{2}n)$. The communication complexity of $P_{1}$ is $O((l+\kappa)m^{2}n)$. The communication complexity of $P_{j}$ is $O((l+\kappa)mn)$. The round complexity is $O(1)$.
  • Total costs.
    • Offline phase. The offline computation complexity per party is $O(\gamma mn\log n+m^{2}n(\log m+\log n))$. The offline communication complexity per party is $O(t\lambda m\log(\gamma n/t)+t\lambda m^{2}\log(n/t)+\lambda m^{2}n(\log m+\log n))$. The offline round complexity is $O(1)$.
    • Online phase. The online computation complexity per party is $O(\gamma mn+m^{2}n)$. The online communication complexity of $P_{1}$ is $O(\gamma mn+(l+\kappa)m^{2}n)$. The online communication complexity of $P_{j}$ is $O(\gamma mn+m^{2}n+(l+\kappa)mn)$. The online round complexity is $O(\log\gamma+m)$.
  • 0.F.3 Theoretical Analysis of Protocol $\Pi_{\mathsf{PK\text{-}MPSU}}$
    The costs of each stage in $\Pi_{\mathsf{PK\text{-}MPSU}}$ (Figure 13) are calculated as follows.
  • Batch ssPMT The cost of this stage is the same as that in $\Pi_{\mathsf{SK\text{-}MPSU}}$ (cf. Appendix 0.F.3).
    • Offline phase. The computation complexity of each party is $O(\gamma mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(mn)$. The communication complexity of each party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
  • Random oblivious transfers and messages rerandomization. We use EC ElGamal encryption to instantiate the MKR-PKE scheme. So each of encryption, partial decryption and rerandomization takes one point scalar operation. The length of ciphertext is $4\lambda$.
    For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke $B$ instances of silent ROT correlations with random inputs during the offline phase. In the online phase, $P_{i}$ sends 1-bit derandomization message to transform each ROT into a chosen-input version. For each OT correlation, each $P_{j}$ executes two encryptions and sends two $4\lambda$-bit messages to $P_{i}$. $P_{i}$ executes one rerandomization. If $i\neq 1$, $P_{i}$ sends one ciphertext to $P_{j}$ and $P_{j}$ executes one rerandomization as well.
    • Offline phase. The computation complexity of each party is $O(mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(mn)$ public-key operations. The communication complexity of each party is $O(\lambda mn)$. The round complexity is $O(1)$.
  • Messages decryptions and shufflings. Each party shuffles $(m-1)B$ ciphertexts and executes $(m-1)B$ partial decryptions before sending them to the next party.
    • Online phase. The computation complexity of each party is $O(mn)$ public-key operations. The communication complexity of each party is $O(\lambda mn)$. The round complexity is $O(m)$.
  • Total costs.
    • Offline phase. The computation complexity of each party is $O(\gamma mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(\gamma n/t)$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(mn)$ symmetric-key operations and $O(mn)$ public-key operations. The communication complexity of each party is $O((\gamma+\lambda)mn)$. The round complexity is $O(\log\gamma+m)$.
  • 0.F.4 The Complete Analysis of LG Protocol
    The costs of each stage in LG protocol are calculated as follows.
    • mq-ssPMT. The cost of this stage mainly consists of three parts:
      • SKE encryption: The computation complexity of encrypting $10,1,\cdots,n-$ 1 is $O(n)$.
      • OKVS: The computation complexities of Encode and Decode algorithms are both $O(n)$. The size of OKVS is $1.28\kappa n$ bits, where $\kappa$ is the size of a SKE ciphertext with the same range as before.
      • ssVODM: The ssVODM protocol requires a total of $(T+l-\log n-1)n$ AND gates, where $T$ is the number of AND gates in the SKE decryption circuit.
      For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke mq-ssPMT of size $B$. Overall, each party $P_{j}$ engages in $(m-1)$ instances of mq-ssPMT.
      • Offline phase. The computation complexity of each party is $O((T+l-\log n)mn\log((T+l-\log n)n))$. The communication complexity of each party is $O(t\lambda m\log(((T+l-\log n)n)/t))$. The round complexity is $O(1)$.
      • Online phase. The computation complexity of each party is $O((T+l-\log n)mn)$. The communication complexity of each party is $O((T+l+\kappa-\log n)mn)$. The round complexity is $O(\log(l-\log n))$.
    • Random oblivious transfers. For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke ROT extension of size $B$. Overall, each party $P_{j}$ engages in $(m-1)$ independent instances of ROT extension of size $B$.
      • Offline phase. The computation complexity of each party is $O(mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
      • Online phase. The computation complexity of each party is $O(mn)$. The communication complexity of each party is $O(mn)$. The round complexity is $O(1)$.
      The costs of the remaining two steps, multi-party secret-shared shuffle and output reconstruction, are exactly the same as those in our SK-MPSU protocol.(cf. Appendix  0.F.2).
    • Total costs.
      • Offline phase. The computation complexity of each party is $O((T+l-\log n)mn\log((T+l-\log n)n)+m^{2}n\log mn)$. The communication complexity of each party is $O(t\lambda m\log(((T+l-\log n)n)/t)+\lambda m^{2}n(\log m+\log n))$. The round complexity is $O(1)$.
      • Online phase. The computation complexity of $P_{1}$ is $O((T+l-\log n)mn+m^{2}n)$. The communication complexity of $P_{1}$ is $O((T+l+\kappa-\log n)mn+(l+\kappa)m^{2}n)$. The computation complexity of $P_{j}$ is $O((T+l-\log n)mn)$. The communication complexity of $P_{j}$ is $O((T+l+\kappa-\log n)mn)$. The round complexity is $O(\log(l-\log n)+m)$.
    Appendix 0.G Multi-Party Private-ID
    We generalize the two-party private-ID [BKM+20] to multi-party setting. Suppose there are $m$ parties, each possessing a set of $n$ elements. The multi-party private-ID functionality assigns a unique random identifier to each element across all input sets, ensuring that identical elements in different sets obtain the same identifier. Each party receives identifiers associated with its own input set, as well as identifiers associated with the union of all parties’ input sets. With multi-party private-ID, the parties can sort their private sets based on a global set of identifiers and perform desired private computations item by item, ensuring alignment of identical elements across their sets. The formal definition of the multi-party private-ID is depicted in Figure 18. We build a concrete multi-party private-ID protocol based on the DDH assumption (described in Figure 19) by extending the “distributed OPRF+PSU” paradigm [CZZ+24b] to multi-party setting.
    • Parameters. $m$ parties $P_{1},\cdots P_{m}$. Size $n$ of input sets. The bit length $l$ of set elements. The range $D$ of identifiers.
    • Functionality. On input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$ from $P_{i}$,
      • For every element $x\in\bigcup_{i=1}^{m}X_{i}$, choose a random identifier $R(x)\leftarrow D$.
      • Define $R^{*}=\{R(x)|x\in\bigcup_{i=1}^{m}X_{i}\}$ and $R_{i}=\{R(x)|x\in X_{i}\}$ for $i\in[m]$.
      • Give output $(R^{*},R_{i})$ to $P_{i}$.
    Figure 18: Multi-Party Private ID Functionality $\mathcal{F}_{\mathsf{MPID}}$
    • Parameters. $m$ parties $P_{1},\cdots P_{m}$. Size $n$ of input sets. The bit length $l$ of set elements. A cyclic group $\mathbb{G}$, where $g$ is the generator and $q$ is the order. The identifie range $D=\mathbb{G}$. Hash function $\mathsf{H}(x):\{0,1\}^{l}\to\mathbb{G}$.
    • Inputs. Each party $P_{i}$ has input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$.
    • Protocol.
      1. 1.
        For $1\leq i\leq m$, $P_{i}$ samples $a_{i}\leftarrow\mathbb{Z}_{q}$ and $k_{i}\leftarrow\mathbb{Z}_{q}$, then sends $\{\mathsf{H}(x_{i}^{1})^{a_{i}},\cdots,\mathsf{H}(x_{i}^{n})^{a_{i}}\}$ to $P_{(i+1)\mod m}$. For $1\leq j<m$, $P_{(i+j)\mod m}$ receives $\{y_{i}^{1},\cdots,y_{i}^{n}\}$ from $P_{(i+j-1)\mod m}$. $P_{(i+j)\mod m}$ computes $\{(y_{i}^{1})^{k_{(i+j)\mod m}},\cdots,(y_{i}^{n})^{k_{(i+j)\mod m}}\}$ and sends to $P_{(i+j+1)\mod m}$.
      2. 2.
        For $1\leq i\leq m$, $P_{i}$ receives $\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}}$ for $j\in[n]$, then it computes $(\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}})^{-a_{i}k_{i}}=\mathsf{H}(x_{i}^{j})^{k_{1}\cdots k_{m}}$. We denote the set $\{\mathsf{H}(x_{i}^{1})^{k_{1}\cdots k_{m}},\mathsf{H}(x_{i}^{n})^{k_{1}\cdots k_{m}}\}$ as $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$, where each $r_{i}^{j}\in\mathbb{G}$.
      3. 3.
        The parties invoke $\Pi_{\mathsf{PK\text{-}MPSU}}$ where $P_{i}$ inputs $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$. $P_{1}$ receives the union $R^{*}=\bigcup_{i=1}^{m}Y_{i}$, and sends it to other parties.
      4. 4.
        Each party $P_{i}$ outputs $(R^{*},R_{i})$.
    Figure 19: DH-based Multi-Party Private ID $\Pi_{\mathsf{MPID}}$
    • Correctness. The first two steps essentially realize a multi-party distributed OPRF protocol, where each party $P_{i}$ inputs a set $\{x_{i}^{1},\cdots,x_{i}^{n}\}$ and receives its own PRF key $k_{i}$ and the PRF values computed on its input set using all parties’ keys $k_{1},\cdots,k_{m}$, denoted as $\{\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{1}),\cdots,\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{n})\}$. In this case, $\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x)=\mathsf{H}(x)^{k_{1}\cdots k_{m}}$. Note that even if $m-1$ parties collude, there exist one exponent private to the adversary, ensuring that the PRF values remain pseudorandom (The proof is analogous to that of the DH-based OPRF, so we omit it here).
    • Security. Assuming $\mathsf{H}$ is a random oracle, the security of protocol follows immediately from the DDH assumption and the security of our PK-MPSU protocol.

### v4

@@
 Multi-party private set union (MPSU) protocol enables $m$ $(m>2)$ parties, each holding a set, to collectively compute the union of their sets without revealing any additional information to other parties.
-There are two main categories of MPSU protocols: The first builds on public-key techniques.
-All existing works in this category involve a super-linear number of public-key operations, resulting in poor practical efficiency.
-The second builds on oblivious transfer and symmetric-key techniques.
-The only existing work in this category is proposed by Liu and Gao (ASIACRYPT 2023), which features the best concrete performance among all existing protocols, despite its super-linear computation and communication.
-Unfortunately, it does not achieve the standard semi-honest security, as it inherently relies on a non-collusion assumption, which is unlikely to hold in practice.
-Therefore, the problem of constructing a practical MPSU protocol based on oblivious transfer and symmetric-key techniques in standard semi-honest model remains open.
-Furthermore, there is no MPSU protocol achieving both linear computation and linear communication complexity, which leaves another unresolved problem.
-In this work, we resolve these two open problems.
+There are two main categories of multi-party private set union (MPSU) protocols: The first category builds on public-key techniques, where existing works require a super-linear number of public-key operations, resulting in their poor practical efficiency.
+The second category builds on oblivious transfer and symmetric-key techniques.
+The only work in this category, proposed by Liu and Gao (ASIACRYPT 2023), features the best concrete performance among all existing protocols, but still has super-linear computation and communication.
+Moreover, it does not achieve the standard semi-honest security, as it inherently relies on a non-collusion assumption, which is unlikely to hold in practice.
+There remain two significant open problems so far: no MPSU protocol achieves semi-honest security based on oblivious transfer and symmetric-key techniques, and no MPSU protocol achieves both linear computation and linear communication complexity.
+In this work, we resolve both of them.
   • We propose the first MPSU protocol based on oblivious transfer and symmetric-key techniques in the standard semi-honest model.
-This protocol is 4.99.3×4.9-9.3$\times$4.9 - 9.3 × faster than Liu and Gao in the LAN setting.
-Concretely, our protocol requires only $3.6$ seconds in online phase for 3 parties with sets of $2^{20}$ items each.
+This protocol is 3.910.0×3.9-10.0$\times$3.9 - 10.0 × faster than Liu and Gao in the LAN setting.
+Concretely, our protocol requires only $4.4$ seconds in online phase for 3 parties with sets of $2^{20}$ items each.
   • We propose the first MPSU protocol achieving both linear computation and linear communication complexity, based on public-key operations.
@@
 We implement our protocols and conduct an extensive experiment to compare the performance of our protocols and the state-of-the-art.
-To the best of our knowledge, our implementation is the first correct and secure implementation of MPSU that reports on large-size experiments.
+To the best of our knowledge, our code is the first correct and secure implementation of MPSU that reports on large-size experiments.
 1 Introduction
@@
 Meanwhile, multi-party PSI [KMP+17, NTY21, CDG+21, BNOP22] is also well-studied.
-In contrast, the advancement of PSU has been sluggish until recently, several works proposed efficient two-party PSU protocols [KRTW19, GMR+21, JSZ+22, ZCL+23, CZZ+24a].
+In contrast, the advancement of PSU has been sluggish until recently, several works proposed efficient two-party PSU protocols [KRTW19, GMR+21, JSZ+22, ZCL+23, CZZ+24a, JSZG24].
 However, multi-party PSU has still not been extensively studied.
@@
 Multi-party private set union (MPSU) enables $m$ $(m>2)$ mutually untrusted parties, each holding a private set of elements, to compute the union of their sets without revealing any additional information.
-MPSU and its variants have numerous applications, such as information security risk assessment [LV04], IP blacklist and vulnerability data aggregation [HLS+16], joint graph computation [BS05], distributed network monitoring [KS05], building block for private DB supporting full join [KRTW19], private ID [GMR+21] etc.
-According to the underlying techniques, existing MPSU protocols can be mainly divided into two categories: The first category, denoted PK-MPSU, is primarily based on public-key techniques, and has been explored in a series of works [KS05, Fri07, VCE22, GNT23].
-The drawbacks of these works are that each party has to perform a substantial number of public-key operations, leading to super-linear computation complexity and poor practical efficiency.
-The second category, denoted SK-MPSU, is primarily based on symmetric-key techniques, and has only one existing work [LG23] to date.
-This work exhibits much better performance than all prior works.
-However, it fails to achieve the standard semi-honest security due to its inherent reliance on a non-collusion assumption, assuming the party who obtains the union (we call it leader hereafter) not to collude with other parties.
-Furthermore, it has super-linear complexity as well.
-Motivated by the above, we raise the following two questions:
-Can we construct a MPSU protocol based on oblivious transfer and symmetric-key operations, without any non-collusion assumptions?
-Can we construct a MPSU protocol with both linear computation and linear communication complexity [^1] In the context of MPSU, linear complexity means that the complexity for each party scales linearly with the total size of all parties’ sets.
-In this paper, we consider the balanced setting where each party holds sets of equal size, thus linear complexity means that the complexity for each party scales linearly with both the number of parties $m$ and the set size $n$.
-Meanwhile, following current conventions in the area of PSO, linear complexity only considers the online phase.?
+MPSU and its variants have numerous applications, such as information security risk assessment [LV04], IP blacklist and vulnerability data aggregation [HLS+16], joint graph computation [BS05], distributed network monitoring [KS05], building block for private DB supporting full join [KRTW19], private ID [GMR+21], etc.
+According to the underlying techniques, existing MPSU protocols [^1] We only consider the special-purpose solutions for MPSU, excluding those based on circuit-based generic techniques of secure computation, due to their unacceptable performance.
+can be divided into two categories:
+  • PK-MPSU: This category is primarily based on public-key techniques, and has been explored in a series of works [KS05, Fri07, VCE22, GNT23].
+A common drawback of these works is that each party has to perform a substantial number of public-key operations, leading to unsatisfactory practical efficiency.
+  • SK-MPSU: This category is primarily based on symmetric-key techniques, and includes only one existing work [LG23] to date.
+This work exhibits much better performance than all prior works but fails to achieve standard semi-honest security due to its inherent reliance on a non-collusion assumption, assuming the party who obtains the union (we call it leader hereafter) not to collude with other parties.
+Both categories share one common limitation: neither of them includes a protocol achieving linear computation and communication complexity [^2] In the context of MPSU, linear complexity means that the complexity per party scales linearly with the total size of all parties’ sets.
+In this paper, we consider the balanced setting where each party holds sets of equal size, thus linear complexity denotes the complexity per party to scale linearly with both the number of parties $m$ and the set size $n$.
+Meanwhile, following current conventions, linear complexity only considers the online phase..
+Motivated by the context, we raise the following two questions:
+Can we construct an MPSU protocol based on oblivious transfer and symmetric-key operations, without any non-collusion assumptions?
+Can we construct an MPSU protocol with both linear computation and linear communication complexity?
 1.1 Our Contribution
-In this work, we resolve the above two open problems by first presenting a new primitive for MPSU, called batch secret-shared private membership test (batch ssPMT), then employing the batch ssPMT to build a SK-MPSU in the standard semi-honest model, and a PK-MPSU with linear computation and communication complexity.
-Our contributions are summarized as follows:
-1.1.1 Efficient Batch ssPMT.
-At the technical core of the state-of-the-art MPSU protocol [LG23] (hereafter referred to as LG protocol) is the multi-query secret-shared private membership test (mq-ssPMT), which dominates both computation and communication costs of LG protocol.
-In analogy of the relation between batched oblivious pseudorandom function (batch OPRF) [KKRT16] and multi-point oblivious pseudorandom function (multi-point OPRF) [PRTY19, CM20], we abstract a new functionality called batch secret-shared private membership test (batch ssPMT), which allows a fairly efficient construction and can be used to build an alternative to mq-ssPMT in the context of MPSU.
-Looking ahead, our batch ssPMT serves as a core building block in our two MPSU protocols, and significantly contributes to our speedup compared to LG protocol.
-1.1.2 SK-MPSU in Standard Semi-Honest Model.
-We generalize random OT (ROT) into multi-party setting, which we call multi-party secret-shared random oblivious transfer (mss-ROT).
-Based on batch ssPMT and mss-ROT, we propose the first SK-MPSU protocol in the standard semi-honest model.
-Compared to LG protocol, our SK-MPSU has superior online performance with a 4.99.3×4.9-9.3$\times$4.9 - 9.3 × improvement in the LAN setting.
-1.1.3 PK-MPSU with Linear Complexity.
+In this work, we answer the above two questions affirmatively. Our contributions are summarized as follows:
+  • Efficient batch ssPMT.
+We present a new primitive called batch secret-shared private membership test (batch ssPMT).
+Compared to multi-query secret-shared private membership test (mq-ssPMT), which is the technical core of the state-of-the-art MPSU protocol [LG23] (hereafter referred to as LG), batch ssPMT admits a much more efficient construction.
+Combined with hashing to bins, batch ssPMT can replace mq-ssPMT in the context of MPSU.
+Looking ahead, our batch ssPMT serves as a core building block in our two MPSU protocols, and significantly contributes to our speedup to LG.
+  • SK-MPSU in standard semi-honest model.
+We generalize random oblivious transfer (ROT) into multi-party setting and present a new primitive called multi-party secret-shared random oblivious transfer (mss-ROT).
+Based on batch ssPMT and mss-ROT, we propose the first SK-MPSU in the standard semi-honest model.
+In addition to enhanced security, our SK-MPSU has superior online / total performance with a 3.910.0×3.9-10.0$\times$3.9 - 10.0 × / 1.27.8×1.2-7.8$\times$1.2 - 7.8 × improvement in the LAN setting.
+  • PK-MPSU with linear complexity.
 Based on batch ssPMT and multi-key rerandomizable public-key encryption (MKR-PKE) [GNT23], we propose the first MPSU protocol with both linear computation and communication.
-Our PK-MPSU has the lowest overall communication costs with a factor of 3.036.5×3.0-36.5$\times$3.0 - 36.5 × improvement compared to LG protocol.
-It also achieves a 1.85.4×1.8-5.4$\times$1.8 - 5.4 × speedup in terms of overall running time in the WAN setting.
-Along the way, we find that the PK-MPSU protocol of Gao et al.
-[GNT23] is insecure against arbitrary collusion and give a practical attack to demonstrate that it necessitates the same non-collusion assumption as LG protocol as well.
+Our PK-MPSU has the lowest overall communication costs with a factor of 3.036.5×3.0-36.5$\times$3.0 - 36.5 × improvement compared to LG.
+Along the way, we find that the MPSU protocol of Gao et al.
+[GNT23] is insecure against arbitrary collusion and give a practical attack to demonstrate that it necessitates non-collusion assumption as well.
+[^3] After we pointed out the security flaw of the GNT protocol [GNT23] with a concrete attack, the authors of [GNT23] contacted us and confirmed our attack.
+Subsequently, they updated their paper and revised their original protocol (c.f.
+Appendix B in the new version of [GNT23]) to a new one, which is similar to our PK-MPSU.
+The main difference lies in that they implement our batch ssPMT by invoking multiple instances of ssPMT separately, which renders their new protocol still has superlinear complexities.
+See also the summary of the relationship in their updated paper.
     Figure 1 depicts the technical overview of our new MPSU framework. We will elaborate the details in Section 2.
@@
     Figure 1: Technical overview of our MPSU framework. The newly introduced primitives are marked with solid boxes. The existing primitives are marked with dashed boxes.
-1.2 Related Works
-We review the existing semi-honest MPSU protocols in the literature.
-1.2.1 PK-MPSU.
+  • 1.2 Related Works
+    We review the existing semi-honest MPSU protocols below.
+    • PK-MPSU.
 Kisser and Song [KS05] introduced the first MPSU protocol, based on polynomial representations and additively homomorphic encryption (AHE).
@@
       Vos et al.
-[VCE22] proposed a MPSU protocol based on the bit-vector representations.
-The parties collectively compute the union by performing the private OR operations on the bit-vectors, realized by ElGamal encryption.
-The original version of this protocol is merely applicable for small universes.
-Even after leveraging divide-and-conquer approach, the protocol requires quadratic computation and communication complexity in the number of parties for the leader and shows poor concrete efficiency reported by [LG23].
+[VCE22] proposed an MPSU protocol based on the bit-vector representations, where the parties collectively compute the union by performing the private OR operations on the bit-vectors, using ElGamal encryption.
+It shows poor concrete efficiency reported by [LG23] and the leader requires quadratic computation and communication in the number of parties.
       Recently, Gao et al.
-[GNT23] proposed a MPSU protocol based on three newly introduced cryptographic tools, including membership Oblivious Transfer (mOT), conditional Oblivious Pseudorandom Function (cOPRF) and MKR-PKE.
-This protocol achieves near-linear complexity in the set size and linear complexity in the number of parties, and is the most advanced MPSU in terms of theoretical complexity.
+[GNT23] proposed an MPSU protocol, representing the most advanced MPSU in terms of asymptotic complexity.
 Unfortunately, their protocol turns out to be insecure against arbitrary collusion.
-We propose a practical attack to show that it requires the same non-collusion assumption as LG protocol (see Appendix 0.A for details).
-1.2.2 SK-MPSU.
+We propose a practical attack to show that it requires the same non-collusion assumption as LG (see Appendix 0.A for details).
+    • SK-MPSU.
 Recently, Liu and Gao [LG23] proposed a practical MPSU protocol based on oblivious transfer and symmetric-key operations.
@@
 However, their protocol is not secure in the standard semi-honest model.
-1.2.3 Other Related Works.
+    • Other Related Works.
 Blanton et al.
-[BA12] proposed a MPSU protocol based on oblivious sorting and generic MPC.
-Their work focuses on devising the circuit for MPSU.
+[BA12] proposed an MPSU protocol based on oblivious sorting and generic multi-party computation (MPC).
 The heavy dependency on general MPC leads to inefficiency.
@@
 pub: public-key operations; sym: symmetric cryptographic operations.
-We ignore the offline phase cost in all SK-MPSU protocols and the symmetric-key operations in all PK-MPSU protocols.
-[KS05] and [Fri07] use Paillier while [VCE22], [GNT23] and our PK-MPSU use ElGamal.
 $n$ is the set size.
-$m$ is the number of participants.
-$\lambda$ and $\sigma$ are computational and statistical security parameter respectively.
+$m$ is the number of parties.
+$\lambda$ and $\sigma$ are computational and statistical security parameter.
 $T$ is the number of AND gate in a SKE decryption circuit in [LG23].
@@
     2 Technical Overview
-2.1 LG Protocol Revisit
-We start by recalling the high-level idea of LG protocol.
-For the sake of simplicity, we focus here on the case of three parties $P_{1},P_{2},P_{3}$, whose inputs are $X_{1},X_{2},X_{3}$ respectively.
+    2.1 LG Revisit
+    We start by abstracting the high-level idea of LG.
+For the sake of simplicity, we focus here on the case of three parties $P_{1},P_{2},P_{3}$, with respective inputs $X_{1},X_{2},X_{3}$.
 We designate $P_{1}$ as the leader.
 Since $P_{1}$ already holds $X_{1}$, it needs to obtain the set difference $Y_{1}=(X_{1}\cup X_{2}\cup X_{3})\setminus X_{1}$ from $P_{2}$ and $P_{3}$.
-Roughly speaking, their protocol enables $P_{2}$ to somehow secret-share $Y_{2}=X_{2}\setminus X_{1}$ and $P_{3}$ to somehow secret-share $Y_{3}=X_{3}\setminus(X_{1}\cup X_{2})$ among all parties.
-Since $\{Y_{2},Y_{3}\}$ is a partition of $Y_{1}$, each party holds a share of $Y_{1}$ in the order of $Y_{2},Y_{3}$ eventually, and any two parties cannot collude to obtain information of the last one’s inputs.
-A naive approach to reconstructing $Y_{1}$ to $P_{1}$ is to let $P_{2}$ and $P_{3}$ send their shares to $P_{1}$ straightforwardly.
-However, in this way, $P_{1}$ can determine the party that each obtained element $x\in Y_{1}$ belongs.
-Their solution is to let the parties invoke multi-party secret-shared shuffle to randomly permute and re-share all shares.
-Multi-party secret-shared shuffle guarantees that any two parties have no knowledge of the permutation and all shares are refreshed after the invocation, hence the adversary is unaware of the correspondence between shares and the individual difference sets $Y_{2},Y_{3}$.
-Then $P_{2}$ and $P_{3}$ can send their shuffled shares to $P_{1}$, who reconstructs $Y_{1}$ and obtains the union by appending elements of $X_{1}$.
-LG protocol utilizes two main ingredients: (1) The secret-shared private membership test (ssPMT) [CO18, ZMS+21], where the sender $\mathcal{S}$ inputs a set $X$, and the receiver $\mathcal{R}$ inputs an element $y$.
-If $y\in X$, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of 1, otherwise they receive secret shares of 0.
+    LG can be summarized as a secret-sharing based MPSU framework with two phases: The first phase involves two secret-sharing processes, where $P_{2}$ somehow secret-shares $Y_{2}=X_{2}\setminus X_{1}$ and $P_{3}$ somehow secret-shares $Y_{3}=X_{3}\setminus(X_{1}\cup X_{2})$ among all parties.
+Since $\{Y_{2},Y_{3}\}$ is a partition of $Y_{1}$, as a result, the parties secret-share $Y_{1}$ in the order of $Y_{2},Y_{3}$.
+The second phase is to reconstruct $Y_{1}$ to $P_{1}$.
+In this case, a straightforward approach is insufficient because $P_{1}$ can identify which party each element $x\in Y_{1}$ originates from.
+The solution is to let the parties invoke multi-party secret-shared shuffle to randomly permute and re-share all shares.
+This shuffle ensures that any two parties have no knowledge of the permutation, thereby concealing the correspondence between the secret shares and individual difference sets $Y_{2},Y_{3}$.
+Afterwards, $P_{2}$ and $P_{3}$ can send their shuffled shares to $P_{1}$, who then reconstructs $Y_{1}$.
+    LG utilizes two ingredients to realize the secret-sharing processes of the framework: (1) The secret-shared private membership test (ssPMT) [CO18, ZMS+21], where the sender $\mathcal{S}$ inputs a set $X$, and the receiver $\mathcal{R}$ inputs an element $y$.
+If $y\in X$, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of 1, otherwise secret shares of 0.
 Liu and Gao proposed a multi-query ssPMT (mq-ssPMT), which supports the receiver querying multiple elements’ memberships of the sender’s set simultaneously.
@@
 $\mathcal{S}$ receives two random messages $r_{0},r_{1}$.
-If $e_{0}\oplus e_{1}=0$, $\mathcal{R}$ receives $r_{0}$, otherwise $r_{1}$ [^2] This special ROT is identical to the standard 1-out-of-2 ROT, where $e_{0}$ is determined by $\mathcal{S}$ to indicate whether it would swap the order of $r_{0}$ and $r_{1}$.
-We use the standard ROT to formally describe our protocol specifications in the later sections.
-Here, we boil down a number of steps to the two-choice-bit ROT for facilitating illustration..
-The following is to elaborate that how to utilize these ingredients to realize the secret-sharing processes.
-The process for $P_{2}$ to secret-share $Y_{2}$ is as follows: $P_{2}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$.
+If $e_{0}\oplus e_{1}=0$, $\mathcal{R}$ receives $r_{0}$, otherwise $r_{1}$ [^4] This special ROT is identical to the standard 1-out-of-2 ROT, where $e_{0}$ is determined by $\mathcal{S}$ to indicate whether to swap the order of $r_{0}$ and $r_{1}$..
+The following is to elaborate the concrete constructions of the two processes.
+    The process for $P_{2}$ to secret-share $Y_{2}$: $P_{2}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$.
 For each item $x\in X_{2}$, $P_{2}$ and $P_{1}$ receive shares $e_{2,1}$ and $e_{1,2}$.
@@
 $P_{1}$ sets the output as its share $s_{2,1}$.
-$P_{2}$ sets its share $s_{2,2}$ to be $r_{2,1}^{0}\oplus x\|\mathsf{H}(x)$ [^3] Suppose a lack of specific structure for set elements, then $P_{1}$ cannot distinguish a set element $x$ and a random value $r$ when it reconstructs the difference set.
-To address this, all parties append the hash value when secret sharing an element, i.e., sharing $x\|\mathsf{H}(x)$.
-It is provable that if the output length of $\mathsf{H}$ is sufficiently long (according to [LG23], at least $\sigma+\log(m-1)+\log n$ bits), the probability of existing $x$ that satisfies $x\|\mathsf{H}(x)=r$ is negligible.
-Therefore, $P_{1}$ can distinguish set elements from random values with overwhelming probability..
+$P_{2}$ sets its share $s_{2,2}$ to be $r_{2,1}^{0}\oplus x\|\mathsf{H}(x)$ [^5] Suppose a lack of specific structure for set elements, then $P_{1}$ cannot distinguish a set element $x$ and a random value $r$ in the reconstruction.
+To address this, the parties append the hash value when secret sharing an element.
+.
 $P_{3}$ sets $s_{2,3}$ to 0.
-If $x\notin X_{1}$, $e_{2,1}\oplus e_{1,2}=0$, $s_{2,1}\oplus s_{2,2}\oplus s_{2,3}=r_{2,1}^{0}\oplus(r_{2,1}^{0}\oplus x\|\mathsf{H}(x))=x\|\mathsf{H}(x)$.
+If $x\notin X_{1}$, $e_{2,1}\oplus e_{1,2}=0$, $s_{2,1}\oplus s_{2,2}\oplus s_{2,3}=x\|\mathsf{H}(x)$.
 Otherwise, $e_{2,1}\oplus e_{1,2}=1$, $s_{2,1}\oplus s_{2,2}\oplus s_{2,3}=r_{2,1}^{1}\oplus r_{2,1}^{0}\oplus x\|\mathsf{H}(x)$ is uniformly random.
-That is, $Y_{2}$ is secret-shared among all parties [^4] Since $X_{2}\setminus X_{1}$ merely contains the information of $X_{1}$ and $X_{2}$, it is sufficient to secret-share $X_{2}\setminus X_{1}$ among $P_{1}$ and $P_{2}$.
-Other difference sets are shared similarly., and the other elements in $X_{2}$ are masked by random values before being secret-share.
-We proceed to the process for $P_{3}$ to secret-share $Y_{3}$: $P_{3}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$ and $P_{2}$ separately.
+That is, $Y_{2}$ is secret-shared among all parties [^6] After the re-sharing of multi-party secret-shared shuffle., and the elements in $X_{1}\cap X_{2}$ are masked by random values.
+    The process for $P_{3}$ to secret-share $Y_{3}$: $P_{3}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$ and $P_{2}$ separately.
 For each $x\in X_{3}$, $P_{3}$ and $P_{1}$ receive shares $e_{3,1}$ and $e_{1,3}$, while $P_{3}$ and $P_{2}$ receive shares $e_{3,2}$ and $e_{2,3}$.
@@
 $P_{3}$ sets its share $s_{3,3}$ to be $r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x)$.
-If $x\notin X_{1}$ and $x\notin X_{2}$, $e_{3,1}\oplus e_{1,3}=0$ and $e_{3,2}\oplus e_{2,3}=0$, $s_{3,1}\oplus s_{3,2}\oplus s_{3,3}=r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus(r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x))=x\|\mathsf{H}(x)$.
-Otherwise, there is at least one random value $r$ ($r_{3,1}^{1}$ or $r_{3,2}^{1}$) cannot be canceled out from the summation.
-The LG protocol described above has two main drawbacks: First, as the heart of LG protocol, mq-ssPMT is given a heavy instantiation which is fed much computation task through expensive general MPC machinery and renders it the bottleneck of the entire protocol.
-Second, LG protocol fails to achieve security against arbitrary collusion.
-In the following two sections, we are devoted to improving the efficiency of mq-ssPMT and enhancing LG protocol into standard semi-honest security, respectively.
+If $x\notin X_{1}$ and $x\notin X_{2}$, $e_{3,1}\oplus e_{1,3}=0$ and $e_{3,2}\oplus e_{2,3}=0$, $s_{3,1}\oplus s_{3,2}\oplus s_{3,3}=x\|\mathsf{H}(x)$.
+Otherwise, there is at least one random value ($r_{3,1}^{1}$ or $r_{3,2}^{1}$) cannot be canceled out from the summation that makes it random.
+    The above outlines how LG works.
+The protocol has two main drawbacks: First, the core component, mq-ssPMT heavily relies on expensive general MPC machinery, which becomes the bottleneck of the entire protocol.
+Second, LG fails to achieve security against arbitrary collusion.
+The next two sections are devoted to addressing these two drawbacks.
     2.2 Efficient Batch ssPMT
-To improve the efficiency of mq-ssPMT, we abstract a new functionality called batch ssPMT, which is essentially the batched version of single-query ssPMT.
-In the batch ssPMT functionality, the sender $\mathcal{S}$ inputs $n$ disjoint sets $X_{1},\cdots,X_{n}$, and the receiver $\mathcal{R}$ inputs $n$ elements $y_{1},\cdots,y_{n}$.
-$\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $n$, whose $i$th bit is 1 if $y_{i}\in X_{i}$, otherwise 0.
-The relationship between batch ssPMT and mq-ssPMT is two-fold: In terms of functionality, batch ssPMT is to mq-ssPMT what batch OPRF is to multi-point OPRF, testing a batch of elements’ memberships across distinct sets rather than a common set; In terms of efficiency, batch ssPMT admits a more efficient construction, providing a superior alternative to mq-ssPMT in the context of MPSU.
-We adopt the following two-step approach to build the batch ssPMT protocol: (1) $\mathcal{S}$ and $\mathcal{R}$ invoke batched oblivious programmable pseudorandom function (batch OPPRF) [KMP+17, PSTY19].
-To elaborate, for the $i$th ssPMT instance ($i=1,\cdots,n$), $\mathcal{S}$ choose a random $s_{i}$, and creates a set $E_{i}$ comprising key-value pairs with elements from $X_{i}$ as keys and $s_{i}$ as all elements’ corresponding values.
-Then $\mathcal{S}$ and $\mathcal{R}$ invoke batch OPPRF of size $n$, where in the $i$th instance of OPPRF, $\mathcal{S}$ inputs $E_{i}$, while $\mathcal{R}$ inputs $y_{i}$ and receives $t_{i}$.
-According to the definition of OPPRF (cf.
-Section 3.5), if $y_{i}\in X_{i}$, $t_{i}$ equals $s_{i}$, otherwise $t_{i}$ is pseudorandom to $s_{i}$.
-If $y_{i}\notin X_{i}$, the probability of $t_{i}=s_{i}$ is $2^{-\gamma}$, where $\gamma$ is the output length of OPPRF, i.e., the bit length of $s_{i}$ and $t_{i}$.
-To ensure that the probability of any $t_{i}\neq s_{i}$ occurring is less than $2^{-\sigma}$ ($\sigma$ is the statistical security parameter), we set $\gamma\geq\sigma+\log n$ so that $n\cdot 2^{-\gamma}\leq 2^{-\sigma}$.
-(2) $\mathcal{S}$ and $\mathcal{R}$ invoke $n$ instances of secret-shared private equality test (ssPEQT) [PSTY19, CGS22], where in the $i$th instance of ssPEQT, $\mathcal{S}$ inputs $s_{i}$ and $\mathcal{R}$ inputs $t_{i}$.
-If $s_{i}=t_{i}$, they receive secret shares of 1, otherwise 0.
-It’s easy to verify that if $y_{i}\in X_{i}$, $\mathcal{S}$ and $\mathcal{R}$ receive shares of 1, otherwise they receive shares of 0 with overwhelming probability.
-Now we start to build an alternative to mq-ssPMT using our batch ssPMT.
-Recall that in mq-ssPMT, the sender $\mathcal{S}$’s input is a single set $X$.
-First, $\mathcal{S}$ and $\mathcal{R}$ preprocess their inputs through hashing to bin technique: $\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign its input elements $y_{1},\cdots,y_{n}$ to $B$ bins via Cuckoo hashing [PR04], which ensures that each bin accommodates at most one element.
+    To improve the efficiency of mq-ssPMT, we introduce a new functionality called batch ssPMT: The sender $\mathcal{S}$ inputs $n$ disjoint sets $X_{1},\cdots,X_{n}$, and the receiver $\mathcal{R}$ inputs $n$ elements $y_{1},\cdots,y_{n}$.
+$\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $n$, where the $i$th bit is 1 if $y_{i}\in X_{i}$, and 0 otherwise.
+    Compared to mq-ssPMT, batch ssPMT enables the testing of element membership across multiple distinct sets rather than within a single common set.
+Their relationship is analogous to that between batched oblivious pseudorandom function (batch OPRF) [KKRT16] and multi-point oblivious pseudorandom function (multi-point OPRF) [PRTY19, CM20].
+Thanks to its batching nature, batch ssPMT admits a much more efficient construction, lending it a superior alternative to mq-ssPMT in the context of MPSU, by coupling with hashing to bins.
+    The alternative to mq-ssPMT works as follows: First, $\mathcal{S}$ and $\mathcal{R}$ preprocess inputs through hashing to bins technique.
+$\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign $y_{1},\cdots,y_{n}$ to $B$ bins via Cuckoo hashing [PR04], ensuring that each bin accommodates at most one element.
 At the same time, $\mathcal{S}$ assigns each $x\in X$ to all bins determined by $h_{1}(x),h_{2}(x),h_{3}(x)$.
-Then the parties invoke $B$ instances of ssPMT, where in the $j$th instance, $\mathcal{S}$ inputs the subset $X_{j}\in X$ containing all elements mapped into its bin $j$ and $\mathcal{R}$ inputs the sole element mapped into its bin $j$.
-If some $y_{i}$ is mapped to bin $j$, and $y_{i}\in X$, then $\mathcal{S}$ certainly maps $y_{i}$ to bin $j$ (and other bins) as well, i.e., $y_{i}\in X_{j}$.
-Therefore, for each bin $j$ of $\mathcal{R}$, if the inside element $y_{i}$ is in $X$, we have $y_{i}\in X_{j}$, then $\mathcal{S}$ and $\mathcal{R}$ receive shares of 1 from the $i$th instance of batch ssPMT.
-Otherwise, we have $y_{i}\notin X_{j}$, then $\mathcal{S}$ and $\mathcal{R}$ receive shares of 0.
-The above process in fact realizes a different functionality with mq-ssPMT.
-In the above construction, the query sequence of $\mathcal{R}$ is determined by the Cuckoo hash positions of its input elements.
-Since the Cuckoo hash positions depends on the whole input set of $\mathcal{R}$, and all shares is arranged in the order of the parties’ Cuckoo hash positions, the reconstruction stage may leak some information about the parties’ input sets to $P_{1}$ and hence cannot be simulated.
-Therefore, it is crucial to eliminate the dependence of shares’ order on the parties’ Cuckoo hash positions before the reconstruction, while fortunately, it has been realized by the execution of multi-party secret-shared shuffle as we discussed before.
-For the difference in efficiency, refer to Section 7.1.
+Then they invoke $B$ instances of ssPMT, where in the $j$th instance, $\mathcal{S}$ inputs the subset $X_{j}\in X$ containing all elements mapped into its bin $j$, while $\mathcal{R}$ inputs the sole element mapped into its bin $j$.
+If some $y_{i}$ is mapped to bin $j$ and $y_{i}\in X$, then $\mathcal{S}$ certainly maps $y_{i}$ to bin $j$ (and other bins) as well, ensuring $y_{i}\in X_{j}$.
+Therefore, for each bin $j$ of $\mathcal{R}$, if the inside element $y_{i}$ is in $X$, $y_{i}\in X_{j}$, $\mathcal{S}$ and $\mathcal{R}$ receive shares of 1 from the $i$th instance of batch ssPMT.
+Otherwise, $y_{i}\notin X_{j}$, $\mathcal{S}$ and $\mathcal{R}$ receive shares of 0.
+    The functionality realized by the above construction differs slightly from mq-ssPMT, as the query sequence of $\mathcal{R}$ is determined by the Cuckoo hash positions of its input elements.
+Since the Cuckoo hash positions depends on the entire input set of $\mathcal{R}$, and all shares is arranged according to the parties’ Cuckoo hash positions, a straightforward reconstruction may leak information about the parties’ input sets to $P_{1}$.
+Therefore, we have to eliminate the dependence of shares’ order on the parties’ Cuckoo hash positions during the reconstruction phase.
+Our crucial insight is that the multi-party secret-shared shuffle used in the reconstruction phase eliminates not only the correspondence between secret shares and difference sets, but also this dependence on the hash positions.
+As a result, our construction can plug in the secret-sharing based framework seamlessly, without introducing any information leakage or additional overhead.
+    Unlike the mq-ssPMT, which heavily relies on general 2PC, our batch ssPMT is built on two highly efficient primitives, batched oblivious programmable pseudorandom function (batch OPPRF) [KMP+17, PSTY19] and secret-shared private equality test (ssPEQT) [PSTY19, CGS22], where the former has an extremely fast specialized instantiation through vector oblivious linear evaluation (VOLE), while the later is composed of a small circuit, allowing for efficient implementation with general 2PC.
+Therefore, our batch ssPMT greatly reduces dependency on general 2PC, leading to a substantial performance improvement.
+For more details, refer to Appendix 0.E.2.2.
     2.3 SK-MPSU from Batch ssPMT and mss-ROT
-By replacing mq-ssPMT with the construction using batch ssPMT plus hash to bin technique as previous, we can improve the performance of LG protocol significantly.
-However, the improved protocol still relies on the non-collusion assumption.
-So it remains to show that how to eliminate the non-collusion assumption to obtain a SK-MPSU protocol in the standard semi-honest model.
-First, let us figure out that why LG protocol needs the non-collusion assumption.
-We still use the aforementioned three-party PSU protocol as illustration: After the secret sharing process of $P_{3}$, all parties hold the secret shares of
-$\begin{cases}x\|\mathsf{H}(x)&x\in Y_{3}\\ r_{3,1}^{0}\oplus r_{3,1}^{1}\oplus x\|\mathsf{H}(x)&x\in Y_{1}\\ r_{3,2}^{0}\oplus r_{3,2}^{1}\oplus x\|\mathsf{H}(x)&x\in Y_{2}\\ r_{3,1}^{1}\oplus r_{3,2}^{1}\oplus r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x)&x\in X_{2}\cap X_{1}\end{cases}$
-In the final stage, if $x\in Y_{3}$, $P_{1}$ reconstructs the secret $x\|\mathsf{H}(x)$.
-Otherwise, the secret is uniformly at random in the $P_{1}$’s view so that $P_{1}$ gains no information about $x$.
-Nevertheless, this is only true when $P_{1}$ does not collude with $P_{3}$.
-Since there is no randomness in $P_{3}$’s view (In the two invocations of ROT, $P_{3}$ receives $r_{3,1}^{0},r_{3,1}^{1}$ and $r_{3,2}^{0},r_{3,2}^{1}$), for each $x\in X_{3}$, the coalition of $P_{1}$ and $P_{3}$ can easily check the reconstructed values, and distinguish the cases that $x\in Y_{1}$, $x\in Y_{2}$, and $x\in X_{2}\cap X_{1}$, which reveals the information of $P_{2}$’s inputs.
-To resolve the above security problem, we generalize ROT into multi-party setting, which we call multi-party secret-shared random oblivious transfer (mss-ROT).
-Suppose there are $d<m$ involved parties and two of them, denoted $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$, possessing choice bits $b_{0}$ and $b_{1}$ respectively.
-The mss-ROT functionality gives each involved party $P_{i}$ ($i\in\{1,\cdots,d\}$) two outputs $r_{i}$ and $\Delta_{i}$, such that if $b_{0}\oplus b_{1}=0$, $r_{1}\oplus\cdots\oplus r_{d}=0$.
-Otherwise, $r_{1}\oplus\cdots\oplus r_{d}=\Delta_{1}\oplus\cdots\oplus\Delta_{d}$.
-Note that if $b_{0}\oplus b_{1}=1$, the parties share a value which is uniformly random in the view of a coalition of any $d-1$ parties.
-Equipped with this new primitive, the above protocol can be amended to achieve security under arbitrary collusion: After $P_{3}$ and $P_{1}$ receiving shares $e_{3,1}$ and $e_{1,3}$, while $P_{3}$ and $P_{2}$ receiving shares $e_{3,2}$ and $e_{2,3}$, $P_{1},P_{2},P_{3}$ invoke mss-ROT twice.
-In the first invocation, $P_{3}$ and $P_{1}$ act as $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ holding as choice bits $e_{3,1}$ and $e_{1,3}$.
-$P_{1},P_{2},P_{3}$ receives $r_{1,31},r_{2,31},r_{3,31}$ and $\Delta_{1,31},\Delta_{2,31},\Delta_{3,31}$ separately.
-If $e_{3,1}\oplus e_{1,3}=0$, $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=0$, otherwise $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=\Delta_{1,31}\oplus\Delta_{2,31}\oplus\Delta_{3,31}$.
-Likewise, in the second invocation, $P_{3}$ and $P_{2}$ hold as choice bits $e_{3,2}$ and $e_{2,3}$.
-$P_{1},P_{2},P_{3}$ receives $r_{1,32},r_{2,32},r_{3,32}$ and $\Delta_{1,32},\Delta_{2,32},\Delta_{3,32}$ separately.
-If $e_{3,2}\oplus e_{2,3}=0$, $r_{1,32}\oplus r_{2,32}\oplus r_{3,32}=0$, otherwise $r_{1,32}\oplus r_{2,32}\oplus r_{3,32}=\Delta_{1,32}\oplus\Delta_{2,32}\oplus\Delta_{3,32}$.
-Finally, $P_{1}$ sets its share $sh_{1}$ to be $r_{1,31}^{0}\oplus r_{1,32}^{0}$.
-$P_{2}$ sets its share $sh_{2}$ to be $r_{2,31}^{0}\oplus r_{2,32}^{0}$.
-$P_{3}$ sets its share $sh_{3}$ to be $r_{3,31}^{0}\oplus r_{3,32}^{0}\oplus x\|\mathsf{H}(x)$.
+    By replacing mq-ssPMT with the construction in the previous section, we improve LG’s performance significantly.
+However, the resulting protocol still relies on the non-collusion assumption.
+We proceed to show how to eliminate it.
+    First, we explain why LG requires the non-collusion assumption using the three-party example: After the secret-sharing process of $P_{3}$, all parties hold the secret shares of
+    $\begin{cases}x\|\mathsf{H}(x)&x\in X_{3}\setminus(X_{1}\cup X_{2})\\ r_{3,1}^{0}\oplus r_{3,1}^{1}\oplus x\|\mathsf{H}(x)&x\in(X_{2}\cup X_{3})\setminus X_{1}\\ r_{3,2}^{0}\oplus r_{3,2}^{1}\oplus x\|\mathsf{H}(x)&x\in(X_{1}\cup X_{3})\setminus X_{2}\\ r_{3,1}^{1}\oplus r_{3,2}^{1}\oplus r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x)&x\in X_{1}\cap X_{2}\cap X_{3}\end{cases}$
+    If $x\in Y_{3}$, $P_{1}$ reconstructs the secret $x\|\mathsf{H}(x)$.
+If $x\notin Y_{3}$, the secret is random so that $P_{1}$ gains no information about $x$.
+Nevertheless, this is only guaranteed when $P_{1}$ does not collude with $P_{3}$.
+Recall that $P_{3}$ receives $r_{3,1}^{0},r_{3,1}^{1}$ and $r_{3,2}^{0},r_{3,2}^{1}$ from the ROT execution with $P_{1}$ and $P_{2}$ respectively, so the coalition of $P_{1}$ and $P_{3}$ can easily check the secret, and distinguish the distinct cases, which reveals information of $P_{2}$’s inputs.
+    To fix this security issue, we generalize ROT into multi-party setting, which we call multi-party secret-shared random oblivious transfer (mss-ROT).
+Consider $d$ involved parties, with two parties, denoted as $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$, possessing choice bits $b_{0}$ and $b_{1}$ respectively.
+There is a subset $J$ of $\{1,\cdots,d\}$ so that each party $P_{j}$ ($j\in J$) holds $\Delta_{j}$ as input.
+The mss-ROT functionality outputs a random $r_{i}$ to each party $P_{i}$, s.t.
+if $b_{0}\oplus b_{1}=0$, $\bigoplus_{i=1}^{d}r_{i}=0$.
+Otherwise, $\bigoplus_{i=1}^{d}r_{i}=\bigoplus_{j\in J}\Delta_{j}$.
+    Equipped with this new primitive, the secret-sharing process of $P_{3}$ is revised as follows: After $P_{3}$ and $P_{1}$ receiving shares $e_{3,1}$ and $e_{1,3}$, $\{P_{1},P_{2},P_{3}\}$ invoke mss-ROT, where $P_{3}$ and $P_{1}$ act as $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ holding $e_{3,1}$ and $e_{1,3}$, while $P_{2}$ and $P_{3}$ samples $\Delta_{2,31},\Delta_{3,31}$ uniformly as inputs.
+$P_{1},P_{2},P_{3}$ receives $r_{1,31},r_{2,31},r_{3,31}$ separately.
+If $e_{3,1}\oplus e_{1,3}=0$, $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=0$, otherwise $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=\Delta_{2,31}\oplus\Delta_{3,31}$.
+After $P_{3}$ and $P_{2}$ receiving shares $e_{3,2}$ and $e_{2,3}$, $\{P_{2},P_{3}\}$ invoke mss-ROT, where $P_{3}$ and $P_{2}$ hold choice bits $e_{3,2}$ and $e_{2,3}$, while $P_{2}$ and $P_{3}$ samples $\Delta_{2,32},\Delta_{3,32}$ uniformly as inputs.
+$P_{2}$ and $P_{3}$ receives $r_{2,32}$ and $r_{3,32}$ separately.
+If $e_{3,2}\oplus e_{2,3}=0$, $r_{2,32}\oplus r_{3,32}=0$, otherwise $r_{2,32}\oplus r_{3,32}=\Delta_{2,32}\oplus\Delta_{3,32}$.
+Finally, $P_{1}$ sets $s_{3,1}=r_{1,31}$.
+$P_{2}$ sets $s_{3,2}=r_{2,31}\oplus r_{2,32}$.
+$P_{3}$ sets $s_{3,3}=r_{3,31}\oplus r_{3,32}\oplus x\|\mathsf{H}(x)$.
     We conclude that all parties hold the secret shares of
-$\begin{cases}x\|\mathsf{H}(x)&x\in Y_{3}\\ \Delta_{1,31}\oplus\Delta_{2,31}\oplus\Delta_{3,31}\oplus x\|\mathsf{H}(x)&x\in Y_{1}\\ \Delta_{1,32}\oplus\Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in Y_{2}\\ \Delta_{1,31}\oplus\Delta_{2,31}\oplus\Delta_{3,31}\oplus\Delta_{1,32}\oplus\Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in X_{2}\cap X_{1}\end{cases}$
-It is immediate that if $x\notin Y_{3}$, the secret-shared value is uniform and independent of $x$ in the view of a coalition of any two parties.
-Hence, the reconstruction reveals no additional information to $P_{1}$ even if it colludes with $P_{2}$ or $P_{3}$.
+    $\begin{cases}x\|\mathsf{H}(x)&x\in X_{3}\setminus(X_{1}\cup X_{2})\\ \Delta_{2,31}\oplus\Delta_{3,31}\oplus x\|\mathsf{H}(x)&x\in(X_{2}\cup X_{3})\setminus X_{1}\\ \Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in(X_{1}\cup X_{3})\setminus X_{2}\\ \Delta_{2,31}\oplus\Delta_{3,31}\oplus\Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in X_{1}\cap X_{2}\cap X_{3}\end{cases}$
+    It is immediate that if $x\notin Y_{3}$, the secret is random in the view of the coalition of $\{P_{1},P_{2}\}$ or $\{P_{1},P_{3}\}$.
+Since the coalition of $\{P_{2},P_{3}\}$ cannot know the secret, the protocol achieves security under arbitrary collusion.
     2.4 PK-MPSU from Batch ssPMT and MKR-PKE
-As previously mentioned, there are no existing MPSU works that achieve both linear computation and communication complexity.
-Thereinto, LG protocol and our SK-MPSU fail to achieve linear complexity because in the secret-sharing based framework, each party holds shares of $(m-1)n$ elements and in the final reconstruction stage, all parties except $P_{1}$ send their shares to $P_{1}$, then $P_{1}$ reconstruct $(m-1)n$ secrets to pick up all elements comprising the union.
-It is clear that even in this single stage, the computation and communication complexity of $P_{1}$ deviate from the linear criteria.
-So we have to turn to another approach.
-We notice that the most advanced work in terms of asymptotic complexity is [GNT23], achieving near-linear complexity.
-We start by analyzing [GNT23]: The first phase enables $P_{1}$ to somehow acquire encrypted $X_{i}\setminus(X_{1}\cup\cdots\cup X_{i-1})$ for $2\leq i\leq m$, interspersed with encrypted dummy messages filling positions of the duplicate elements.
-Note that if $P_{1}$ decrypts these ciphertexts by itself, then it can establish a mapping associating each element $x\in X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$ with the party to whom $x$ belongs.
-To address this problem, they introduce a PKE variant, MKR-PKE, with several tailored properties, enabling all parties to conduct a collaborative decryption and shuffle procedure.
-The second phase, which is the aforemetioned collaborative decryption and shuffle procedure, works as follows: $P_{1}$ sends the ciphertexts to $P_{2}$ after permuting them using a random permutation $\pi_{1}$.
-$P_{2}$ performs partial decryption on the received ciphertexts before rerandomization, and permutes them using a random permutation $\pi_{2}$.
-Then it forwards the permuted partially decrypted ciphertexts to $P_{3}$.
-This iterative process continues until the last party, $P_{m}$, who returns its permuted partially decrypted ciphertexts to $P_{1}$.
-$P_{1}$ decrypts the ultimate ciphertexts, filters out the dummy elements, retains the desired set $X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$, and appends the elements in $X_{1}$, to obtain the union.
-We identify that their non-linear complexity stems from their construction of the first phase.
-Despite that they also use hashing to bin technique, their construction does not support amortizing the batching cost.
-In the case $B=O(n)$ bins are considered, each party has to execute mOT and cOPRF of size $O(\log\log n)$ per bin, resulting in overall $O(n\log\log n)$ computation and communication complexity.
-Besides, their insecurity against arbitrary collusion roots in the construction of the first phase (cf.
-Appendix 0.A).
-In contrast, the second phase already has linear computation and communication complexity, as well as security against arbitrary collusion.
-[^5] As we have said, the second phase addresses the same problem solved by the multi-party secret-shared shuffle, meanwhile there currently exists no construction for multi-party secret-shared shuffle that achieves total linear complexity.
-This is the root why our PK-MPSU protocol can achieve total linear complexity while our SK-MPSU protocol cannot.
-The key distinction lies in our PK-MPSU protocol’s departure from the secret-shared paradigm, opting instead for a PKE-based approach to replace the non-linear component.
-Therefore, the task of devising linear-complexity MPSU reduces to devising a linear-complexity and secure construction against arbitrary collusion for the first phase.
-Our first attempt is as follows: For $2\leq i\leq m$, each $P_{i}$ and $P_{1}$ execute batch ssPMT (after preprocessing their inputs using hashing to bin, we ignore this for the sake of simplicity in the later explanation) and receive secret shares as the choice bits for the subsequent execution of two-choice-bit OT [^6] The difference of functionality between two-choice-bit OT and two-choice-bit ROT is that $\mathcal{S}$ inputs two messages $m_{0},m_{1}$ instead of obtaining two uniform messages., where for each $x\in X_{i}$, $P_{i}$ acts as $\mathcal{S}$ and inputs the encrypted $x$ and an encrypted dummy message.
-As a result, if $x$ belongs to the difference set $X_{i}\setminus X_{1}$, $P_{1}$ receives the encrypted $x$; otherwise, it receives the encrypted dummy message.
-Nevertheless, the goal of the first phase is to obtain the encrypted difference set $X_{i}\setminus(X_{1}\cup\cdots\cup X_{i-1})$, which is the subset of $X_{i}\setminus X_{1}$.
-In order to attain the above goal, our next strategy is to let each $P_{i}$ engage in a procedure with each $P_{j}$ ($1<j<i$) before executing the batch ssPMT and two-choice-bit OT with $P_{1}$, so that $P_{j}$ can “obliviously” replace the encrypted items in the intersection $X_{j}\cap X_{i}$ with encrypted dummy messages.
-The procedure unfolds as follows.
-$P_{i}$ plays the role of $\mathcal{R}$ and executes the batch ssPMT with $P_{j}$.
+    There are no existing MPSU protocols with linear computation and communication complexity.
+Notably, it is impossible to achieve this goal within the secret-sharing based framework, given that $P_{1}$ must reconstruct $(m-1)n$ secrets with each secret comprising $m$ shares.
+Therefore, we have to seek for another technical route.
+    Our start point is the existing work [GNT23] with the best asymptotic complexity $mn(\log n/\log\log n)$.
+We distill their protocol into a framework with two phases: The first phase allows $P_{1}$ to somehow obtain encrypted $X_{i}\setminus(X_{1}\cup\cdots\cup X_{i-1})$ for $2\leq i\leq m$, interspersed with encrypted dummy messages filling the positions of duplicate elements.
+However, if $P_{1}$ decrypts ciphertexts by itself, then it could associate each element in $X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$ with the specific party it originates from.
+This problem is addressed by the second phase, which is the collaborative decryption and shuffle procedure, making use of a PKE variant MKR-PKE: $P_{1}$ sends the ciphertexts to $P_{2}$ after permuting them using a random permutation $\pi_{1}$.
+$P_{2}$ performs a partial decryption on the received ciphertexts, rerandomizes them, permutes them with a random permutation $\pi_{2}$, and forwards the ciphertexts to $P_{3}$.
+This iterative process continues until the last party, $P_{m}$, returns its permuted partially decrypted ciphertexts to $P_{1}$.
+Finally, $P_{1}$ fully decrypts the ciphertexts, filters out the dummy elements, retains the set $X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$, and appends $X_{1}$ to compute the union.
+    We identify that both the non-linear complexities and insecurity against arbitrary collusion of [GNT23] arise from their construction in the first phase.
+Therefore, the task of building a linear-complexity MPSU reduces to devising a linear-complexity and secure construction for this phase.
+    A rough idea is as follows: Let $P_{i}$ hold a set of encrypted elements in $X_{i}$.
+For $2\leq i\leq m$, each $P_{i}$ engage in a procedure with each $P_{j}$ ($1<j<i$) such that $P_{j}$ can “obliviously” replace the encrypted elements in $X_{i}\cap X_{j}$ with encrypted dummy messages.
+After the procedure, $P_{i}$ holds a set of encrypted elements in $X_{i}\setminus(X_{2}\cup\cdots\cup X_{i-1})$.
+The replacement of encrypted elements in $X_{1}$ is handled at the end.
+    The replacing procedure between each $P_{i}$ and $P_{j}$: (1) $P_{i}$ acts as $\mathcal{R}$ and executes batch ssPMT (after preprocessing their inputs using hashing to bins, which we omit here for simplicity) with $P_{j}$.
 For each $x\in X_{i}$, $P_{i}$ and $P_{j}$ receive shares $e_{i,j}$ and $e_{j,i}$.
-$P_{i}$ initializes a message $m_{0}$ to the encrypted $x$ and $m_{1}$ to an encrypted dummy message, then executes two-choice-bit OT with $P_{j}$.
-$P_{i}$ and $P_{j}$ input $e_{i,j}$ and $e_{j,i}$ as their choice bits.
-$P_{i}$ acts as $\mathcal{S}$ and inputs $m_{0}$ and $m_{1}$.
-$P_{j}$ acts as $\mathcal{R}$ and receives $\mathsf{ct}=m_{e_{i,j}\oplus e_{j,i}}$.
+(2) If $j=2$, $P_{i}$ initializes $m_{0}$ as the encrypted $x$ and $m_{1}$ as an encrypted dummy message.
+(3) $P_{i}$ acts as $\mathcal{S}$ and executes two-choice-bit OT [^7] The difference from two-choice-bit ROT is that $\mathcal{S}$ inputs two specific messages $m_{0},m_{1}$ instead of obtaining two uniform messages.
+with $P_{j}$, where $P_{i}$ and $P_{j}$ input $e_{i,j}$ and $e_{j,i}$ as choice bits.
+$P_{i}$ inputs $m_{0},m_{1}$ and $P_{j}$ receives $\mathsf{ct}=m_{e_{i,j}\oplus e_{j,i}}$.
 If $x\notin X_{j}$, $\mathsf{ct}$ is the ciphertext of $x$; otherwise, the ciphertext of dummy message.
-$P_{j}$ rerandomizes $\mathsf{ct}$ to $\mathsf{ct}^{\prime}$ and returns $\mathsf{ct}^{\prime}$ to $P_{i}$.
-$P_{i}$ rerandomizes $\mathsf{ct}^{\prime}$, then updates $m_{0}$ to $\mathsf{ct}^{\prime}$ and $m_{1}$ to the rerandomized $m_{1}$ before repeating the above procedure with the next party $P_{j+1}$.
-The final values of $m_{0}$ and $m_{1}$ shall be used during the subsequent invocation of two-choice-bit OT with $P_{1}$.
+(4) $P_{j}$ rerandomizes $\mathsf{ct}$ to $\mathsf{ct}^{\prime}$ and returns $\mathsf{ct}^{\prime}$ to $P_{i}$.
+(5) $P_{i}$ rerandomizes $\mathsf{ct}^{\prime}$, updates $m_{0}$ to $\mathsf{ct}^{\prime}$, and rerandomizes $m_{1}$ before repeating the procedure with the next party $P_{j+1}$.
+Note that only if all conditions of $x\notin X_{j}$ are satisfied, the final $m_{0}$ is an encrypted message of $x$.
+    For $2\leq i\leq m$, each $P_{i}$ and $P_{1}$ execute batch ssPMT and receive secret shares as choice bits for the subsequent execution of two-choice-bit OT.
+For each $x\in X_{i}$, $P_{i}$ acts as $\mathcal{S}$ and inputs the final $m_{0}$ of the above procedure and an encrypted dummy message.
+As a result, if $x\notin X_{1}$, $P_{1}$ receives the final $m_{0}$; otherwise, it receives the encrypted dummy message.
+This excludes all elements in $X_{1}$ from the encrypted set $X_{i}\setminus(X_{2}\cup\cdots\cup X_{i-1})$ and lets $P_{1}$ obtain the results.
+    The linear complexities of our construction are attributed to two key facts: First, this framework ensures that each interaction between $P_{i}$ and $P_{j}$ only involves the input sets of themselves.
+Second, our batch ssPMT protocol has linear computation and communication complexity (cf.
+Section 4).
     3 Preliminaries
@@
     The ideal functionality of MPSU is formalized in Figure 2.
-  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$, where $P_{\textnormal{ld}}$ is the leader, $\textnormal{ld}\in[m]$. Size $n$ of input sets. The bit length $l$ of set elements.
-  • Functionality. On input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$ from $P_{i}$, give $\bigcup_{i=1}^{m}X_{i}$ output to $P_{\textnormal{ld}}$.
+    • Parameters. $m$ parties $P_{1},\cdots,P_{m}$, where $P_{1}$ is the leader. Size $n$ of input sets. The bit length $l$ of set elements.
+    • Functionality. On input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$ from $P_{i}$ ($i\in[m]$), output $\bigcup_{i=1}^{m}X_{i}$ to $P_{1}$.
     Figure 2: Multi-party Private Set Union Functionality $\mathcal{F}_{\mathsf{mpsu}}$
-3.3 Batch Oblivious Pseudorandom Function
+    3.3 Batch Oblivious Programmable Pseudorandom Function
     Oblivious pseudorandom function (OPRF) [FIPR05] is a central primitive in the area of PSO.
 Kolesnikov et al.
-[KKRT16] introduced batched OPRF, which provides a batch of OPRF instances in the following way.
-In the $i$th instance, the receiver $\mathcal{R}$ has an input $x_{i}$; the sender $\mathcal{S}$ learns a PRF key $k_{i}$ and $\mathcal{R}$ learns $\mathsf{PRF}(k_{i},x_{i})$.
-Note that $\mathcal{R}$ evaluates the PRF on only one point per key.
-The batch OPRF functionality is described formally in Figure 3.
-  • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of set elements. Some PRF family $\mathsf{PRF}:\{0,1\}^{*}\times\{0,1\}^{l}\to\{0,1\}^{\gamma}$.
-  • Functionality. On input $\vec{x}\subseteq(\{0,1\}^{l})^{B}$ from $\mathcal{R}$,
-    • For each $i\in[B]$, choose uniform PRF key $k_{i}$.
-    • For each $i\in[B]$, define $f_{i}=\mathsf{PRF}(k_{i},x_{i})$.
-    • Give vector $\vec{k}=(k_{1},\cdots,k_{B})$ to $\mathcal{S}$ and vector $\vec{f}=(f_{1},\cdots,f_{B})$ to $\mathcal{R}$.
-Figure 3: Batch OPRF Functionality $\mathcal{F}_{\mathsf{bOPRF}}$
-3.4 Oblivious Key-Value Stores
-A key-value store [PRTY20, GPR+21, RR22, BPSY23] is a data structure that compactly represents a desired mapping from a set of keys to corresponding values.
-Definition 1
-A key-value store is parameterized by a set $\mathcal{K}$ of keys, a set $\mathcal{V}$ of values, and a set of functions $\mathsf{H}$, and consists of two algorithms:
-  • $\mathsf{Encode_{H}}$ takes as input a set of key-value pairs $\{(k_{i},v_{i})|i\in[n]\}$ and outputs an object $D$ (or, with statistically small probability, an error indicator $\perp$).
-  • $\mathsf{Decode_{H}}$ takes as input an object $D$, a key $k$, and outputs a value $v$.
-Correctness. For all $A\subseteq\mathcal{K}\times\mathcal{V}$ with distinct keys:
-$(k,v)\in A\ \ and\ \perp\neq D\leftarrow\mathsf{Encode_{H}}(A)\Longrightarrow\mathsf{Decode_{H}}(D,k)=v$
-Obliviousness.
-For all distinct $\{k_{1}^{0},\cdots,k_{n}^{0}\}$ and all distinct $\{k_{1}^{1},\cdots,k_{n}^{1}\}$, if $\mathsf{Encode_{H}}$ does not output $\perp$ for $\{k_{1}^{0},\cdots,k_{n}^{0}\}$ or $\{k_{1}^{1},\cdots,k_{n}^{1}\}$, then the distribution of $D_{0}\leftarrow\mathsf{Encode_{H}}(\{(k_{1}^{0},v_{1}),\cdots,(k_{n}^{0},v_{n})\})$ is computationally indistinguishable to $D_{1}\leftarrow\mathsf{Encode_{H}}(\{(k_{1}^{1},v_{1}),\cdots,(k_{n}^{1},v_{n})\})$, where $v_{i}\leftarrow\mathcal{V}$ for $i\in[n]$.
-A KVS is an Oblivious KVS (OKVS) if it satisfies the obliviousness property. we also require an additional randomness property [ZCL+23] from the OKVS.
-Randomness.
-For any $A=\{(k_{1},v_{1}),\cdots,(k_{n},v_{n})\}$ and $k^{*}\notin\{k_{1},\cdots,k_{n}\}$, the output of $\mathsf{Decode_{H}}(D,k^{*})$ is statistically indistinguishable to that of uniform distribution over $\mathcal{V}$, where $D\leftarrow\mathsf{Encode_{H}}(A)$.
-3.5 Batch Oblivious Programmable Pseudorandom Function
-Oblivious programmable pseudorandom function (OPPRF) [KMP+17, PSTY19, CGS22, RS21, RR22] is an extension of OPRF, which lets $\mathcal{S}$ program a PRF $F$ so that it has specific random outputs for some specific inputs and pseudorandom outputs for all other inputs.
-This kind of PRF with the additional property that on a certain programmed set of inputs the function outputs programmed values is called programmable PRF (PPRF) [PSTY19].
+[KKRT16] introduced batched OPRF, which provides a batch of OPRF instances.
+In the $i$th instance, the sender $\mathcal{S}$ learns a PRF key $k_{i}$, while the receiver $\mathcal{R}$ inputs $x_{i}$ and learns $\mathsf{PRF}(k_{i},x_{i})$.
+    Oblivious programmable pseudorandom function (OPPRF) [KMP+17, PSTY19, CGS22, RS21, RR22] is an extension of OPRF, which lets $\mathcal{S}$ program a PRF $F$ so that it has specific uniform outputs for some specific inputs and pseudorandom outputs for all other inputs.
+This kind of PRF with the additional property that the function outputs programmed values on a certain programmed set of inputs is called programmable PRF (PPRF) [PSTY19].
 $\mathcal{R}$ evaluates the OPPRF with no knowledge of whether it learns a programmed output of $F$ or just a pseudorandom value.
 The batch OPPRF functionality is given in Figure 3.
-  • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of keys. The bit length $\gamma$ of values. An OKVS scheme $(\mathsf{Encode_{H}},\mathsf{Decode_{H}})$.
+    • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of keys. The bit length $\gamma$ of values.
     • Sender’s inputs. $\mathcal{S}$ inputs $B$ sets of key-value pairs including:
       • Disjoint key sets $K_{1},\cdots,K_{B}$.
-    • The value sets $V_{1},\cdots,V_{B}$, where $\lvert K_{i}\rvert=\lvert V_{i}\rvert$ for every $i\in[B]$, and $V_{i}(j)\in\{0,1\}^{\gamma}$ for every $i\in[B]$ and $1\leq j\leq\lvert K_{i}\rvert$.
+      • The value sets $V_{1},\cdots,V_{B}$, where $\lvert K_{i}\rvert=\lvert V_{i}\rvert$, $i\in[B]$.
     • Receiver’s inputs. $\mathcal{R}$ inputs $B$ queries $\vec{x}\subseteq(\{0,1\}^{l})^{B}$.
     • Functionality: On input $(K_{1},\cdots,K_{B})$ and $(V_{1},\cdots,V_{B})$ from $\mathcal{S}$ and $\vec{x}\subseteq(\{0,1\}^{l})^{B}$ from $\mathcal{R}$,
-    • Choose uniformly random and independent PPRF key $k_{i}$, for each $i\in[B]$;
+      • Choose a uniform PPRF key $k_{i}$, for $i\in[B]$;
       • Sample a PPRF $F:\{0,1\}^{*}\times\{0,1\}^{l}\to\{0,1\}^{\gamma}$ such that $F(k_{i},K_{i}(j))=V_{i}(j)$ for $i\in[B],1\leq j\leq\lvert K_{i}\rvert$;
-    • Define $f_{i}=F(k_{i},x_{i})$, for each $i\in[B]$;
+      • Define $f_{i}=F(k_{i},x_{i})$, for $i\in[B]$;
       • Give vector $\vec{f}=(f_{1},\cdots,f_{B})$ to $\mathcal{R}$.
     Figure 3: Batch OPPRF Functionality $\mathcal{F}_{\mathsf{bOPPRF}}$
-3.6 Hashing to Bin
-The hashing to bin technique was introduced by Pinkas et al.
-[PSZ14, PSSZ15], which is originally applied to construct two-party PSI protocol.
-At the high level, the receiver $\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign its items to $B$ bins via Cuckoo hashing [PR04], so that each bin has at most one item.
-The hashing process uses eviction and the choice of which of the bins is used depends on the entire set.
-Using the same hash functions and simple hashing, sender $\mathcal{S}$ assigns each of its items $x$ to all of the bins $h_{1}(x),h_{2}(x),h_{3}(x)$, so that if the item $x$ is also in $\mathcal{R}$’s set and is mapped into the bin $b\in\{h_{1}(x),h_{2}(x),h_{3}(x)\}$ by Cuckoo hashing, then the bin $b$ of $\mathcal{S}$’s simple hash table certainly contains $x$.
-We write simple hashing with the following notation:
+    3.4 Hashing to Bins
+    The hashing to bins technique was introduced by Pinkas et al.
+[PSZ14, PSSZ15] to construct two-party PSI.
+At a high level, the receiver $\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign its items to $B$ bins via Cuckoo hashing [PR04], so that each bin has at most one item.
+The Cuckoo hashing process uses eviction and the choice of bins for each item depends on the entire set.
+On the other hand, the sender $\mathcal{S}$ assigns each of its items $x$ to all bins $h_{1}(x),h_{2}(x),h_{3}(x)$ via simple hashing.
+This guarantees that for each item $x$ of $\mathcal{R}$, if $x$ is mapped into the $b$th bin of Cuckoo hash table ($b\in\{h_{1}(x),h_{2}(x),h_{3}(x)\}$), and $x$ is in $\mathcal{S}$’s set, then the $b$th of simple hash table certainly contains $x$.
+    We denote simple hashing with the following notation:
     $\mathcal{T}^{1},\cdots,\mathcal{T}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X)$
-This expression means to hash the items of $X$ into $B$ bins using simple hashing with hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$.
-The output is the simple hash table denoted by $\mathcal{T}^{1},\cdots,\mathcal{T}^{B}$, where for each $x\in X$ we have $\mathcal{T}^{h_{i}(x)}=\{x\|i|i=1,2,3\}$.
+    This expression represents hashing the items of $X$ into $B$ bins using simple hashing with hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$.
+The output is a simple hash table denoted by $\mathcal{T}^{1},\cdots,\mathcal{T}^{B}$, where for each $x\in X$ we have $\mathcal{T}^{h_{i}(x)}\supseteq\{x\|i|i=1,2,3\}$.
 [^8] Appending the index of the hash function is helpful for dealing with edge cases like $h_{1}(x)=h_{2}(x)=i$, which happen with non-negligible probability.
-Without appending $i$, the OKVS would contain the identical value $F(k_{i},x)$, which leaks the fact that a collision $h_{1}(x)=h_{2}(x)$ occurred.
-Such an event is input-dependent so cannot be simulated.
-We write Cuckoo hashing with the following notation:
+    We denote Cuckoo hashing with the following notation:
     $\mathcal{C}^{1},\cdots,\mathcal{C}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X)$
-This expression means to hash the items of $X$ into $B$ bins using Cuckoo hashing on hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$.
-The output is the Cuckoo hash table denoted by $\mathcal{C}^{1},\cdots,\mathcal{C}^{B}$, where for each $x\in X$ there is some $i\in\{1,2,3\}$ such that $\mathcal{C}^{h_{i}(x)}=x\|i$.
-Some Cuckoo hash positions do not matter, corresponding to empty bins.
+    This expression represents hashing the items of $X$ into $B$ bins using Cuckoo hashing with hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$.
+The output is a Cuckoo hash table denoted by $\mathcal{C}^{1},\cdots,\mathcal{C}^{B}$, where for each $x\in X$ there is some $i\in\{1,2,3\}$ such that $\mathcal{C}^{h_{i}(x)}=\{x\|i\}$.
+Some Cuckoo hash positions are irrelevant, corresponding to empty bins.
 We use these symbols throughout the subsequent sections.
     3.5 Secret-Shared Private Equality Test
-Secret-shared private equality test protocol (ssPEQT) [PSTY19, CGS22] can be viewed as an extreme case of ssPMT when the sender $\mathcal{S}$’s input set has only one item.
-In Figure 5 we formally define this functionality.
-  • Parameters. Two parties $P_{1},P_{2}$. The bit length $\gamma$ of inputs.
-  • Functionality.
-On input $x$ from $P_{1}$ and input $y$ from $P_{2}$, sample two random bits $a,b$ under the constraint that $a\oplus b=\bigwedge_{i=1}^{\gamma}(1\oplus x[i]\oplus y[i])$.
-Namely, if $x=y$, $a\oplus b=1$.
-Otherwise $a\oplus b=0$.
-Give $a$ to $P_{1}$ and $b$ to $P_{2}$.
+    Secret-shared private equality test protocol (ssPEQT) [PSTY19, CGS22] can be viewed as an extreme case of ssPMT when the sender $\mathcal{S}$’s input set size is 1.
+Figure 4 defines its functionality.
+    • Parameters. Two parties $P_{1},P_{2}$. The input bit length $\gamma$.
+    • Functionality. On input $x$ from $P_{1}$ and input $y$ from $P_{2}$, sample two random bits $a,b$ s.t. if $x=y$, $a\oplus b=1$. Otherwise $a\oplus b=0$. Give $a$ to $P_{1}$ and $b$ to $P_{2}$.
     Figure 4: Secret-Shared Private Equality Test Functionality $\mathcal{F}_{\mathsf{ssPEQT}}$
@@
     3.7 Multi-Party Secret-Shared Shuffle
-Multi-party secret-shared shuffle has the capability to permute the share vectors of all parties in a random permutation unknown to any coalition of $m-1$ parties.
-Then it refreshs all shares.
-The functionality is given in Figure 7.
+    Multi-party secret-shared shuffle functionality works by randomly permuting the share vectors of all parties and then refreshing all shares, ensuring that the permutation remains unknown to any coalition of $m-1$ parties.
+The formal functionality is given in Figure 6.
     Eskandarian et al.
-[EB22] propose an online-efficient protocol by extending [CGP20] to the multi-party setting.
-In the offline phase, each party generates a random permutation and a set of correlated vectors called share correlation.
-In the online phase, each party permutes and refreshes the share vectors efficiently using share correlation.
-We give the functionality of share correlation and details of the online protocol in Appendix 0.C.
+[EB22] proposed an online-efficient protocol by extending [CGP20] to the multi-party setting.
+In the offline phase, each party generates a random permutation and a set of correlated vectors called share correlations.
+In the online phase, each party efficiently permutes and refreshes the share vectors using these share correlations.
     • Parameters. $m$ parties $P_{1},\cdots P_{m}$. The dimension of vector $n$. The item length $l$.
@@
     Figure 6: Multi-Party Secret-Shared Shuffle Functionality $\mathcal{F}_{\mathsf{ms}}$
-3.10 Multi-Key Rerandomizable Public-Key Encryption
     Gao et al.
-[GNT23] introduce the concept of multi-key rerandomizable public-key encryption (MKR-PKE), which is a variant of PKE with several additional properties.
-Let $\mathcal{SK}$ denote the space of secret keys which forms an abelian group under the operation $+$, $\mathcal{PK}$ the space of public keys which forms an abelian group under the operation $\cdot$, $\mathcal{M}$ the space of plaintexts, and $\mathcal{C}$ the space of ciphertexts.
+[GNT23] introduced the concept of multi-key rerandomizable public-key encryption (MKR-PKE), a variant of PKE with several additional properties.
+Let $\mathcal{SK}$ denote the space of secret keys, which forms an abelian group under the operation $+$, and $\mathcal{PK}$ denote the space of public keys, which forms an abelian group under the operation $\cdot$.
+$\mathcal{M}$ denotes the space of plaintexts, and $\mathcal{C}$ denotes the space of ciphertexts.
 MKR-PKE is a tuple of PPT algorithms $(\mathsf{Gen},\mathsf{Enc},\mathsf{ParDec},\mathsf{Dec},\mathsf{ReRand})$ such that:
@@
     • The rerandomization algorithm $\mathsf{ReRand}$ takes as input a public key $pk\in\mathcal{PK}$ and a ciphertext $\mathsf{ct}\in\mathcal{C}$, outputs another ciphertext $\mathsf{ct}^{\prime}\in\mathcal{C}$.
-MKR-PKE requires the following additional properties besides those of PKE with indistinguishable multiple encryptions. (cf. Appendix 0.B):
-3.10.1 Partially Decryptable
-For any two pairs of keys $(sk_{1},pk_{1})\leftarrow\mathsf{Gen}(1^{\lambda})$ and $(sk_{2},pk_{2})\leftarrow\mathsf{Gen}(1^{\lambda})$ and any $x\in\mathcal{M}$, it holds that
+    MKR-PKE is an IND-CPA secure PKE scheme that requires the following additional properties:
+    • Partially Decryptable. For any two pairs of keys $(sk_{1},pk_{1})\leftarrow\mathsf{Gen}(1^{\lambda}),(sk_{2},pk_{2})\leftarrow\mathsf{Gen}(1^{\lambda})$ and any $x\in\mathcal{M}$,
       $\mathsf{ParDec}(sk_{1},\mathsf{Enc}(pk_{1}\cdot pk_{2},x))=\mathsf{Enc}(pk_{2},x)$
-3.10.2 Rerandomizable
-For any $pk\in\mathcal{PK}$ and any $x\in\mathcal{M}$, it holds that
+    • Rerandomizable. For any $pk\in\mathcal{PK}$ and any $x\in\mathcal{M}$,
 […]
-[GNT23] uses elliptic curve based ElGamal encryption [ElG85] to instantiate MKR-PKE. The concrete EC MKR-PKE is described in Appendix 0.D.
+      Gao et al. [GNT23] make use of elliptic curve (EC) based ElGamal encryption [ElG85] to instantiate MKR-PKE. The concrete EC MKR-PKE is described in Appendix 0.C.
       Remark 1
-We note that only a few of elliptic curves support efficient encoding from bit-strings to EC points.
+      To the best of our knowledge, there is few elliptic curves that support efficient encoding and decoding between bit-strings and EC points.
 Therefore, the plaintext space of EC MKR-PKE is generally restricted to EC points to guarantee rerandomizable property.
     4 Batch Secret-Shared Private Membership Test
-The batch secret-shared private membership test (batch ssPMT) is a central building block in both our SK-MPSU and PK-MPSU protocols.
-In this section, we formally introduce this functionality and provide a construction built by the aforementioned primitives.
-The batch ssPMT is a two-party protocol implementing a batch of ssPMT instances between a sender and a receiver.
-Assuming a batch size of $B$, the sender $\mathcal{S}$ inputs $B$ sets $X_{1},\cdots,X_{B}$, while the receiver $\mathcal{R}$ inputs $B$ elements $x_{1},\cdots,x_{B}$.
-Consequently, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $B$, whose $i$th bit is 1 if $x_{i}\in X_{i}$, otherwise 0.
-The batch ssPMT functionality is presented in Figure 8 and the construction is given in Figure 9.
+    The batch secret-shared private membership test (batch ssPMT) is a central building block in our SK-MPSU and PK-MPSU protocols.
+It is a two-party protocol that implements multiple instances of ssPMT between a sender $\mathcal{S}$ and a receiver $\mathcal{R}$.
+Given a batch size of $B$, $\mathcal{S}$ inputs $B$ sets $X_{1},\cdots,X_{B}$, while $\mathcal{R}$ inputs $B$ elements $x_{1},\cdots,x_{B}$.
+As a result, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $B$, where the $i$th bit is 1 if $x_{i}\in X_{i}$, 0 otherwise.
+The batch ssPMT functionality is presented in Figure 7 and the construction is given in Figure 8, built with batch OPPRF and ssPEQT.
     • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of set elements. The output length $\gamma$ of OPPRF.
@@
     • Functionality.
-On inputs $X_{1},\cdots,X_{B}$ from $\mathcal{S}$ and input $\vec{x}$ from $\mathcal{R}$, for each $i\in[B]$, sample two random bits $e_{S}^{i},e_{R}^{i}$ under the constraint that if $x_{i}\in X_{i},e_{S}^{i}\oplus e_{R}^{i}=1$, otherwise $e_{S}^{i}\oplus e_{R}^{i}=0$.
+On inputs $X_{1},\cdots,X_{B}$ from $\mathcal{S}$ and input $\vec{x}$ from $\mathcal{R}$, for each $i\in[B]$, sample two random bits $e_{S}^{i},e_{R}^{i}$ s.t.
+if $x_{i}\in X_{i},e_{S}^{i}\oplus e_{R}^{i}=1$, otherwise $e_{S}^{i}\oplus e_{R}^{i}=0$.
 Give $\vec{e}_{S}=(e_{S}^{1},\cdots,e_{S}^{B})$ to $\mathcal{S}$ and $\vec{e}_{R}=(e_{R}^{1},\cdots,e_{R}^{B})$ to $\mathcal{R}$.
@@
         The parties invoke $B$ instances of $\mathcal{F}_{\mathsf{ssPEQT}}$, where in the $i$th instance $\mathcal{S}$ inputs $s_{i}$ and $\mathcal{R}$ inputs $t_{i}$.
-In the end, $\mathcal{S}$ receives $e_{S}^{i}\in\{0,1\}$ and $\mathcal{R}$ receives $e_{R}^{i}\in\{0,1\}$.
+In the end, $\mathcal{S}$ and $\mathcal{R}$ receives $e_{S}^{i},e_{R}^{i}\in\{0,1\}$ respectively.
     Figure 8: Batch ssPMT $\Pi_{\mathsf{bssPMT}}$
+    Theorem 4.1
+    Protocol $\Pi_{\mathsf{bssPMT}}$ securely realizes $\mathcal{F}_{\mathsf{bssPMT}}$ in the $(\mathcal{F}_{\mathsf{bOPPRF}},\mathcal{F}_{\mathsf{ssPEQT}})$-hybrid model.
     • Correctness.
-According to the functionality of batch OPPRF, if $x_{i}\in X_{i}$, then $\mathcal{R}$ receives $t_{i}=s_{i}$.
-Subsequently, in the $i$th instance of ssPEQT, they input $s_{i}$ and $t_{i}$ which are equal, hence their outputs satisfy $e_{S}^{i}\oplus e_{R}^{i}=1$.
-Conversely, if $x_{i}\notin X_{i}$, then $\mathcal{R}$ receives a pseudorandom value $t_{i}$.
-The probability of $t_{i}=s_{i}$ equals $2^{-\gamma}$ and any $t_{i}=s_{i},i\in[B]$ occurs with a probability of $B\cdot 2^{-\gamma}$.
-As we set $\gamma\geq\sigma+\log B$, so $B\cdot 2^{-\gamma}\leq 2^{-\sigma}$, which means the probability of any $t_{i}=s_{i}$ occurring is negligible.
-After the invocation of ssPEQT, we conclude that if $x_{i}\notin X_{i}$, $e_{S}^{i}$ and $e_{R}^{i}$ satisfy $e_{S}^{i}\oplus e_{R}^{i}=0$ with overwhelming probability.
-Theorem 4.1
-Protocol $\Pi_{\mathsf{bssPMT}}$ securely realizes $\mathcal{F}_{\mathsf{bssPMT}}$ in the $(\mathcal{F}_{\mathsf{bOPRF}},\mathcal{F}_{\mathsf{ssPEQT}})$-hybrid model.
-The security of the protocol follows immediately from the security of the batch OPPRF and the ssPEQT functionalities.
-  • Complexity Analysis.
-We set $B=O(n)$ to be consistent with our MPSU protocols.
-One thing to clarify is that the asymptotic complexity, to a great extent, depends on the choices of the instantiation of batch OPRF and the generation method of Beaver Triples [Bea91] in ssPEQT.
-In consideration of concrete efficiency, we employ subfield vector oblivious linear evaluation (subfield-VOLE) [BCG+19a, BCG+19b, RRT23] for both options.
-A comprehensive complexity analysis is in Appendix 0.F.1.
-We summarize the total costs as below:
-    • Offline phase. The computation complexity of each party is $O(\gamma n\log n)$. The communication complexity of each party is $O(t\lambda\log(\gamma n/t))$. The round complexity is $O(1)$.
-    • Online phase. The computation complexity of each party is $O(n)$. The communication complexity of each party is $O(\gamma n)$. The round complexity is $O(\log\gamma)$.
+According to the batch OPPRF functionality, if $x_{i}\in X_{i}$, $\mathcal{R}$ receives $t_{i}=s_{i}$.
+Then in the $i$th instance of ssPEQT, since $s_{i}=t_{i}$, $e_{S}^{i}\oplus e_{R}^{i}=1$.
+Conversely, if $x_{i}\notin X_{i}$, $\mathcal{R}$ receives a pseudorandom value $t_{i}$.
+The probability that any $t_{i}=s_{i}$ for $i\in[B]$ is $B\cdot 2^{-\gamma}$.
+By setting $\gamma\geq\sigma+\log B$, we have $B\cdot 2^{-\gamma}\leq 2^{-\sigma}$, meaning the probability of any such collision is negligible [^9] In MPSU, the batch OPPRF is invoked multiple times, and the lower bound of $\gamma$ is relevant to the total number of batch OPPRF invocations.
+Refer to  0.E.2 for more details..
+After the invocation of ssPEQT, we have that if $x_{i}\notin X_{i}$, $e_{S}^{i}\oplus e_{R}^{i}=0$ with overwhelming probability.
+    • Security. The security of the protocol follows immediately from the security of batch OPPRF and ssPEQT functionalities.
+    In our MPSU protocols, the batch ssPMT is always combined with the hashing to bins technique.
+The combined construction has good efficiency, together with linear computation and communication in term of $n$, which mainly benefits from the following technical advances: First, we follow the paradigm in [PSTY19] to construct batch OPPRF from batch OPRF and oblivious key-value store (OKVS) [PRTY20, GPR+21, RR22, BPSY23].
+By leveraging the technique to amortize communication, the total communication of computing $B=O(n)$ instances of OPPRF is equal to the total number of items $3n$.
+Second, we utilize subfield vector oblivious linear evaluation (subfield-VOLE) [BCG+19a, BCG+19b, RRT23] to instantiate batch OPRF and the construction in [RR22] to instantiate OKVS.
+This ensures the computation complexity of batch OPPRF of size $O(n)$ to scale linearly with $n$.
+A comprehensive analysis is in Appendix 0.E.2.
     5 MPSU from Symmetric-Key Techniques
-In this section, we introduce a new primitive called multi-party secret-shared random oblivious transfer (mss-ROT), then we utilize its two-choice-bit version to construct a SK-MPSU based on oblivious transfer and symmetric-key operations in the standard semi-honest model.
+    In this section, we introduce a new primitive called multi-party secret-shared random oblivious transfer (mss-ROT), then we utilize it to build SK-MPSU based on oblivious transfer and symmetric-key operations in the standard semi-honest model.
     5.1 Multi-Party Secret-Shared Random Oblivious Transfer
-The abstraction of the general version of mss-ROT is inspired by the generation of Beaver Triples in multi-party setting.
-In the two-party GMW protocol, a Beaver Triple is produced by two role-switching executions of 1-bit ROT [ALSZ13].
-As a result, $P_{1}$ holds $a_{1},b_{1},c_{1}$ and $P_{2}$ holds $a_{2},b_{2},c_{2}$, such that $(a_{1}\oplus a_{2})\cdot(b_{1}\oplus b_{2})=c_{1}\oplus c_{2}$.
-In the multi-party setting, the functionality of Beaver Triple is extended to let each party $P_{i}$ hold $a_{i},b_{i},c_{i}$ with the correlation $(\bigoplus_{i=1}^{m}a_{i})\cdot(\bigoplus_{i=1}^{m}b_{i})=\bigoplus_{i=1}^{m}c_{i}$.
-Given that $(\bigoplus_{i=1}^{m}a_{i})\cdot(\bigoplus_{i=1}^{m}b_{i})$ can be written as $(2-m)(\bigoplus_{i=1}^{m}a_{i}\cdot b_{i})+\bigoplus_{1\leq i\leq j\leq m}(a_{i}+a_{j})\cdot(b_{i}+b_{j})$, multi-party Beaver Triple reduces to pairwise two-party Beaver Triples.
-In consideration of the simple fact that the functionality of ROT can be interpreted as $r_{0}\oplus r_{b}=b\cdot\Delta$, where $\Delta=r_{0}\oplus r_{1}$, we follow the approach to producing multi-party Beaver Triples to extend ROT into multi-party setting, which is to let each party $P_{i}$ to individually input a choice bit $b_{i}$ and receive the output shares $r_{i}$ and $\Delta_{i}$ such that $r_{1}\oplus\cdots\oplus r_{m}=(b_{1}\oplus\cdots\oplus b_{m})\cdot(\Delta_{1}\oplus\cdots\oplus\Delta_{m})$.
-In the context of MPSU, we only need two-choice-bit version of mss-ROT, which allows a more efficient construction with less pairwise invocations of two-party ROT compared to the general version.
-To elaborate, the two-choice-bit version of mss-ROT (for simplicity, we call it mss-ROT for short hereafter) allows two parties $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ to hold their choice bits $b_{0}$ and $b_{1}$ s.t.
-$r_{1}\oplus\cdots\oplus r_{m}=(b_{0}\oplus b_{1})\cdot(\Delta_{1}\oplus\cdots\oplus\Delta_{m})$.
-We give the formal functionality in Figure 10 and the detailed construction in Figure 11.
+    In Figure 5, the ROT functionality can be interpreted as $r_{0}\oplus r_{b}=b\cdot\Delta$, where $\Delta=r_{0}\oplus r_{1}$.
+As the aforementioned two-choice-bit ROT, it can be consider to secret-share the choice bit $b$ between two parties and interpreted as $r_{0}\oplus r_{b}=(b_{0}\oplus b_{1})\cdot\Delta$.
+We further extend this secret-sharing idea to multiple parties, allowing not only any non-empty subset of the involved parties to secret-share $b$, but also any non-empty subset of the involved parties to secret-share $\Delta$.
+    To build our SK-MPSU, we only need two-choice-bit version of mss-ROT.
+For simplicity, we will refer to this as mss-ROT hereafter.
+The formal functionality [^10] We choose to let the parties input the shares of $\Delta$ instead of outputting them, as the former functionality simply implies the latter.
+is given in Figure 9 and the detailed construction is given in Figure 10.
     Theorem 5.1
@@
 And the correctness and security of the mss-ROT protocol stems from the correctness and security of ROT.
-For the complete proof, refer to Appendix 0.E.1.
+For the complete proof, refer to Appendix 0.D.1.
     5.2 Construction of Our SK-MPSU
-We now turn our attention to construct a SK-MPSU. The construction of our SK-MPSU follows the high-level ideas we introduced in the technical overview and is formally presented in Figure 12.
+    We now turn our attention to construct a SK-MPSU. The construction follows the high-level ideas we introduced in the technical overview and is formally presented in Figure 11.
     Theorem 5.2
-Protocol $\Pi_{\mathsf{SK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{mss\text{-}rot}},\mathcal{F}_{\mathsf{ms}})$-hybrid model, where $P_{1}=P_{\textnormal{ld}}$.
-The security proof of Theorem 5.2 is deferred to Appendix 0.E.2.
-  • Complexity Analysis. We provide a comprehensive complexity analysis for our SK-MPSU protocol in Appendix 0.F.2. The total costs are summarized below:
-    • Offline phase.
-The offline computation complexity per party is $O(\gamma mn\log n+m^{2}n(\log m+\log n))$.
-The offline communication complexity per party is $O(t\lambda m\log(\gamma n/t)+t\lambda m^{2}\log(n/t)+\lambda m^{2}n(\log m+\log n))$.
-The offline round complexity is $O(1)$.
-    • Online phase.
-The online computation complexity per party is $O(m^{2}n)$.
-The online communication complexity of $P_{1}$ is $O(\gamma mn+(l+\kappa)m^{2}n)$.
-The online communication complexity of $P_{j}$ is $O(\gamma mn+m^{2}n+(l+\kappa)mn)$.
-The online round complexity is $O(\log\gamma+m)$.
+    Protocol $\Pi_{\mathsf{SK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{mss\text{-}rot}},\mathcal{F}_{\mathsf{ms}})$-hybrid model.
+    The proof of Theorem 5.2 is in Appendix 0.D.2. We provide a comprehensive complexity analysis for our SK-MPSU / LG and an exhaustive comparison between them in Appendix 0.E.
     • Parameters.
 $m$ parties $P_{1},\cdots,P_{m}$, where $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ provide inputs as shares of the choice bit, $\text{ch}_{0},\text{ch}_{1}\in[m]$.
+We use $J=\{j_{1},\cdots,j_{d}\}$ to denote the set of indices for the parties who provide inputs as shares of $\Delta$, where $d\leq m$.
 The message length $l$.
-  • Functionality. On input $b_{0}\in\{0,1\}$ from $P_{\text{ch}_{0}}$ and $b_{1}\in\{0,1\}$ from $P_{\text{ch}_{1}}$,
-    • Sample $r_{2},\cdots r_{m},\Delta_{1},\Delta_{2},\cdots,\Delta_{m}\leftarrow\{0,1\}^{l}$ and give $(r_{j},\Delta_{j})$ to $P_{j}$ for $2\leq j\leq m$.
-    • If $b_{0}\oplus b_{1}=0$, compute $r_{1}=\bigoplus_{j=2}^{m}r_{j}$, else compute $r_{1}=\Delta_{1}\oplus(\bigoplus_{j=2}^{m}(r_{j}\oplus\Delta_{j}))$. Give $r_{1}$ to $P_{1}$.
+    • Functionality. On input $b_{0}\in\{0,1\}$ from $P_{\text{ch}_{0}}$, $b_{1}\in\{0,1\}$ from $P_{\text{ch}_{1}}$, and $\Delta_{j}$ from each $P_{j}$ where $j\in J$,
+      • Sample $r_{i}\leftarrow\{0,1\}^{l}$ and give $r_{i}$ to $P_{i}$ for $2\leq i\leq m$.
+      • If $b_{0}\oplus b_{1}=0$, compute $r_{1}=\bigoplus_{i=2}^{m}r_{i}$, else compute $r_{1}=\bigoplus_{i=2}^{m}r_{i}\oplus(\bigoplus_{j\in J}\Delta_{j})$. Give $r_{1}$ to $P_{1}$.
     Figure 9: Multi-Party Secret-Shared Random OT Functionality $\mathcal{F}_{\mathsf{mss\text{-}rot}}$
@@
 where $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ provide inputs as shares of the choice bit, $\text{ch}_{0},\text{ch}_{1}\in[m]$.
-We use $J$ to denote the set of indices for the remaining parties who do not provide a choice-bit input, i.e., $J=[m]\setminus\{\text{ch}_{0},\text{ch}_{1}\}$.
+We use $J=\{j_{1},\cdots,j_{d}\}$ to denote the set of indices for the parties who provide inputs as shares of $\Delta$, where $d\leq m$.
 The message length $l$.
-  • Inputs. $P_{\text{ch}_{0}}$ has input $b_{0}\in\{0,1\}$ and $P_{\text{ch}_{1}}$ has input $b_{1}\in\{0,1\}$.
+    • Inputs. $P_{\text{ch}_{0}}$ has input $b_{0}\in\{0,1\}$ and $P_{\text{ch}_{1}}$ has input $b_{1}\in\{0,1\}$. Each $P_{j}$ for $j\in J$ has input $\Delta_{j}$,
     • Protocol.
       1. 1.
-      $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ invoke $\mathcal{F}_{\mathsf{rot}}$ twice: First, $P_{\text{ch}_{0}}$ acts as receiver with input $b_{0}$ and $P_{\text{ch}_{1}}$ acts as sender.
-$P_{\text{ch}_{0}}$ receives $r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{b_{0}}\in\{0,1\}^{l}$.
-$P_{\text{ch}_{1}}$ receives $r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{0},r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{1}\in\{0,1\}^{l}$; Second, $P_{\text{ch}_{1}}$ acts as receiver with input $b_{1}$ and $P_{\text{ch}_{0}}$ acts as sender.
-$P_{\text{ch}_{1}}$ receives $r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{b_{1}}\in\{0,1\}^{l}$.
-$P_{\text{ch}_{0}}$ receives $r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{0},r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{1}\in\{0,1\}^{l}$.
+        For $1\leq i\leq m$, $P_{i}$ initializes $r_{i}=0$. If $ch_{j=0,1}\in J$, $P_{ch_{j}}$ sets $r_{ch_{j}}=b_{j}\cdot\Delta_{ch_{j}}$ ($\cdot$ denotes bitwise-AND).
       2. 2.
-      For $j\in J$: $P_{\text{ch}_{0}}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{\text{ch}_{0}}$ acts as receiver with input $b_{0}$ and $P_{j}$ as sender without input.
-$P_{\text{ch}_{0}}$ receives $r_{j,{\text{ch}_{0}}}^{b_{0}}\in\{0,1\}^{l}$.
-$P_{j}$ receives $r_{j,{\text{ch}_{0}}}^{0},r_{j,{\text{ch}_{0}}}^{1}\in\{0,1\}^{l}$.
+        For $j\in J$: $P_{\text{ch}_{0}}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{\text{ch}_{0}}$ acts as receiver with input $b_{0}$ and $P_{j}$ as sender.
+$P_{\text{ch}_{0}}$ receives $u_{j,{\text{ch}_{0}}}^{b_{0}}\in\{0,1\}^{l}$.
+$P_{j}$ receives $u_{j,{\text{ch}_{0}}}^{0},u_{j,{\text{ch}_{0}}}^{1}\in\{0,1\}^{l}$.
+$P_{j}$ updates $r_{j}=r_{j}\oplus u_{j,{\text{ch}_{0}}}^{0}$ and sends $\Delta^{\prime}_{j,{\text{ch}_{0}}}=u_{j,{\text{ch}_{0}}}^{0}\oplus u_{j,{\text{ch}_{0}}}^{1}\oplus\Delta_{j}$ to $P_{\text{ch}_{0}}$.
+$P_{\text{ch}_{0}}$ updates $r_{\text{ch}_{0}}=r_{\text{ch}_{0}}\oplus u_{j,{\text{ch}_{0}}}^{b_{0}}\oplus b_{0}\cdot\Delta^{\prime}_{j,{\text{ch}_{0}}}$.
       3. 3.
         For $j\in J$: $P_{\text{ch}_{1}}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{\text{ch}_{1}}$ acts as receiver with input $b_{1}$ and $P_{j}$ as sender.
-$P_{\text{ch}_{1}}$ receives $r_{j,{\text{ch}_{1}}}^{b_{1}}\in\{0,1\}^{l}$.
-$P_{j}$ receives $r_{j,{\text{ch}_{1}}}^{0},r_{j,{\text{ch}_{1}}}^{1}\in\{0,1\}^{l}$.
+$P_{\text{ch}_{1}}$ receives $u_{j,{\text{ch}_{1}}}^{b_{1}}\in\{0,1\}^{l}$.
+$P_{j}$ receives $u_{j,{\text{ch}_{1}}}^{0},u_{j,{\text{ch}_{1}}}^{1}\in\{0,1\}^{l}$.
+$P_{j}$ updates $r_{j}=r_{j}\oplus u_{j,{\text{ch}_{1}}}^{0}$ and sends $\Delta^{\prime}_{j,{\text{ch}_{1}}}=u_{j,{\text{ch}_{1}}}^{0}\oplus u_{j,{\text{ch}_{1}}}^{1}\oplus\Delta_{j}$ to $P_{\text{ch}_{1}}$.
+$P_{\text{ch}_{1}}$ updates $r_{\text{ch}_{1}}=r_{\text{ch}_{1}}\oplus u_{j,{\text{ch}_{1}}}^{b_{1}}\oplus b_{1}\cdot\Delta^{\prime}_{j,{\text{ch}_{1}}}$.
       4. 4.
-      For $j\in J$: $P_{j}$ samples $\Delta_{j}\leftarrow\{0,1\}^{l}$ and computes $r_{j}=r_{j,{\text{ch}_{0}}}^{0}\oplus r_{j,{\text{ch}_{1}}}^{0}$.
-$P_{j}$ sends $u_{j,{\text{ch}_{0}}}=\Delta_{j}\oplus r_{j,{\text{ch}_{0}}}^{0}\oplus r_{j,{\text{ch}_{0}}}^{1}$ to $P_{\text{ch}_{0}}$, and $u_{j,{\text{ch}_{1}}}=\Delta_{j}\oplus r_{j,{\text{ch}_{1}}}^{0}\oplus r_{j,{\text{ch}_{1}}}^{1}$ to $P_{\text{ch}_{1}}$, then outputs $(r_{j},\Delta_{j})$.
+        If $\lvert I=\{ch_{0},ch_{1}\}\cup J\rvert<m$, each $P_{i}$ ($i\in I$) samples $r^{\prime}_{i,i^{\prime}}$ and sends $r^{\prime}_{i,i^{\prime}}$ to each $P_{i^{\prime}}$ ($i^{\prime}\in[m]\setminus I$) and updates $r_{i}=r_{i}\oplus r^{\prime}_{i,i^{\prime}}$.
+$P_{i^{\prime}}$ updates $r_{i^{\prime}}=r_{i^{\prime}}\oplus r^{\prime}_{i,i^{\prime}}$.
       5. 5.
-      $P_{\text{ch}_{0}}$ computes $\Delta_{\text{ch}_{0}}=r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{0}\oplus r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{1}$ and $r_{\text{ch}_{0}}=\bigoplus_{j\in J}(r_{j,{\text{ch}_{0}}}^{b_{0}}\oplus b_{0}\cdot u_{j,{\text{ch}_{0}}})\oplus r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{0}\oplus r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{b_{0}}\oplus b_{0}\cdot\Delta_{\text{ch}_{0}}$ ($\cdot$ denotes bitwise-AND between the repetition code of $b_{0}$ and $u_{j,{\text{ch}_{0}}}$, which are both strings of length $l$.
-Similarly hereinafter), then outputs $(r_{\text{ch}_{0}},\Delta_{\text{ch}_{0}})$.
-$P_{\text{ch}_{1}}$ computes $\Delta_{\text{ch}_{1}}=r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{0}\oplus r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{1}$ and $r_{\text{ch}_{1}}=\bigoplus_{j\in J}(r_{j,{\text{ch}_{1}}}^{b_{1}}\oplus b_{1}\cdot u_{j,{\text{ch}_{1}}})\oplus r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{0}\oplus r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{b_{1}}\oplus b_{1}\cdot\Delta_{\text{ch}_{1}}$, then outputs $(r_{\text{ch}_{1}},\Delta_{\text{ch}_{1}})$.
+        For $1\leq i\leq m$, $P_{i}$ outputs the final value of $r_{i}$.
     Figure 10: Multi-Party Secret-Shared Random OT $\Pi_{\mathsf{mss\text{-}rot}}$
@@
 Cuckoo hashing parameters: hash functions $h_{1},h_{2},h_{3}$ and number of bins $B$.
-An OKVS scheme $(\mathsf{Encode_{H}},\mathsf{Decode_{H}})$.
 A collision-resisitant hash function $\mathsf{H}(x):\{0,1\}^{l}\to\{0,1\}^{\kappa}$.
-  • Inputs. Each party $P_{i}$ has input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$.
+    • Inputs. Each party $P_{i}$ inputs $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$.
     • Protocol.
       1. 1.
-      Hashing to bin.
-For $1\leq j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
+        Hashing to bins.
+$P_{1}$ does $\mathcal{T}_{1}^{1},\cdots,\mathcal{T}_{1}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{1})$.
+For $1<j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
       2. 2.
@@
         Multi-party secret-shared random oblivious transfers.
-For $1\leq i<j\leq m,1\leq b\leq B$: $P_{1},\cdots,P_{j}$ invoke $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ where $P_{i}$ acts as $P_{\text{ch}_{0}}$ with input $e_{i,j}^{b}$ and $P_{j}$ acts as $P_{\text{ch}_{1}}$ with input $e_{j,i}^{b}$.
-For $1\leq d<j$, $P_{d}$ receives $r_{d,ji}^{b},\Delta_{d,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{d,j}^{b}=\bigoplus_{i=1}^{j-1}r_{d,ji}^{b}$.
-$P_{j}$ receives $r_{j,ji}^{b},\Delta_{j,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{j,j}^{b}=\bigoplus_{i=1}^{j-1}r_{j,ji}^{b}\oplus(\mathsf{Elem}(\mathcal{C}_{j}^{b})\|\mathsf{H}(\mathsf{Elem}(\mathcal{C}_{j}^{b})))$ if $\mathcal{C}_{j}^{b}$ is not corresponding to an empty bin, otherwise chooses $u_{j,j}^{b}$ at random.
+For $1\leq i<j\leq m,1\leq b\leq B$: $P_{min(2,i)},\cdots,P_{j}$ invoke $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ where $P_{i}$ acts as $P_{\text{ch}_{0}}$ with input $e_{i,j}^{b}$ and $P_{j}$ acts as $P_{\text{ch}_{1}}$ with input $e_{j,i}^{b}$.
+For $1<j^{\prime}\leq j$, $P_{j^{\prime}}$ samples $\Delta_{j^{\prime},ji}^{b}\leftarrow\{0,1\}^{l+\kappa}$ and inputs $\Delta_{j^{\prime},ji}^{b}$ as shares of $\Delta$.
+For $min(2,i)\leq d<j$, $P_{d}$ receives $r_{d,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{d,j}^{b}=\bigoplus_{i=1}^{j-1}r_{d,ji}^{b}$.
+$P_{j}$ receives $r_{j,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{j,j}^{b}=\bigoplus_{i=1}^{j-1}r_{j,ji}^{b}\oplus(\mathsf{Elem}(\mathcal{C}_{j}^{b})\|\mathsf{H}(\mathsf{Elem}(\mathcal{C}_{j}^{b})))$ if $\mathcal{C}_{j}^{b}$ is not corresponding to an empty bin, otherwise chooses $u_{j,j}^{b}$ at random.
 $\mathsf{Elem}(\mathcal{C}_{j}^{b})$ denotes the element in $\mathcal{C}_{j}^{b}$.
@@
     6 MPSU from Public-Key Techniques
-In this section, we describe how to construct a PK-MPSU achieving both linear computation and linear communication complexity. The construction of our PK-MPSU is formally presented in Figure 13.
+    In this section, we describe how to construct a PK-MPSU achieving linear computation and linear communication complexity. The construction is formally presented in Figure 12.
     Theorem 6.1
-Protocol $\Pi_{\mathsf{PK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{rot}})$-hybrid model, assuming the security of MKR-PKE scheme, where $P_{1}=P_{\textnormal{ld}}$.
-The security of the above protocol is based on security of the underlying building blocks, batch ssPMT and ROT, along with the rerandomizable property and indistinguishable multiple encryptions of MKR-PKE.
-For the complete proof, refer to Appendix 0.E.3.
-  • Complexity Analysis. We provide a comprehensive complexity analysis for our PK-MPSU protocol in Appendix 0.F.3. The total costs are summarized below:
-    • Offline phase.
-The offline computation complexity per party is $O(\gamma mn\log n)$.
-The offline communication complexity per party is $O(t\lambda m\log(\gamma n/t))$.
-The offline round complexity is $O(1)$.
-    • Online phase.
-The online computation complexity per party is $O(mn)$ symmetric-key operations and $O(mn)$ public-key operations.
-The onlne communication complexity per party is $O((\gamma+\lambda)mn)$.
-The online round complexity is $O(\log\gamma+m)$.
+    Protocol $\Pi_{\mathsf{PK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{rot}})$-hybrid model, assuming the rerandomizable property and indistinguishable multiple encryptions (cf.
+Appendix 0.B) of MKR-PKE scheme.
+    For the complete proof, refer to Appendix 0.D.3. A comprehensive complexity analysis and an exhaustive comparison with [GNT23] can be found in Appendix 0.E.
+    Remark 2
+    When instantiating our PK-MPSU framework with EC MKR-PKE, the element space is set as EC points accordingly, which might limit its usage in applications.
+We also identified the same issue in [GNT23].
+However, we argue that this limitation is not inherent in our PK-MPSU (nor to the protocol in[GNT23]), as it can be addressed by finding an elliptic curve with efficient encoding and decoding algorithms or by exploring alternative instantiations of MKR-PKE.
+    Despite the limitation, our EC-element-space PK-MPSU remains useful in many scenarios.
+For instance, we present the first multi-party private ID protocol in Appendix 0.F, where our EC-element-space PK-MPSU MPSU serves as a core subprotocol.
+Our multi-party private-ID takes advantages of our PK-MPSU and achieves linear computation and communication complexity and low communication costs as well.
     • Parameters.
@@
 Cuckoo hashing parameters: hash functions $h_{1},h_{2},h_{3}$ and number of bins $B$.
-An OKVS scheme $(\mathsf{Encode_{H}},\mathsf{Decode_{H}})$.
 A MKR-PKE scheme $\mathcal{E}=(\mathsf{Gen},\mathsf{Enc},\mathsf{ParDec},\mathsf{Dec},\mathsf{ReRand})$.
@@
       1. 1.
-      Hashing to bin.
-For $1\leq j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
+        Hashing to bins.
+$P_{1}$ does $\mathcal{T}_{1}^{1},\cdots,\mathcal{T}_{1}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{1})$.
+For $1<j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
       2. 2.
@@
         1. (a)
-        For $2\leq j\leq m,1\leq b\leq B$: $P_{j}$ defines a vector $\vec{c}_{j}$ and sets $c_{j}^{b}=\mathsf{Enc}(pk,\mathsf{Elem}(\mathcal{C}_{j}^{b}))$.
+          For $2\leq j\leq m,1\leq b\leq B$: $P_{j}$ defines $\vec{c}_{j}$ and sets $c_{j}^{b}=\mathsf{Enc}(pk,\mathsf{Elem}(\mathcal{C}_{j}^{b}))$.
 $\mathsf{Elem}(\mathcal{C}_{j}^{b})$ denotes the element in $\mathcal{C}_{j}^{b}$.
           For $2\leq i<j$: $P_{i}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{i}$ acts as receiver with input $e_{i,j}^{b}$ and $P_{j}$ acts as sender.
-$P_{i}$ receives $r_{j,i,e_{i,j}^{b}}^{b}\in\{0,1\}^{\lambda}$.
+$P_{i}$ receives $r_{i,j}^{b}=r_{j,i,e_{i,j}^{b}}^{b}\in\{0,1\}^{\lambda}$.
 $P_{j}$ receives $r_{j,i,0}^{b},r_{j,i,1}^{b}\in\{0,1\}^{\lambda}$.
-$P_{j}$ computes $u_{j,i,e_{j,i}^{b}}^{b}=r_{j,i,e_{j,i}^{b}}^{b}\oplus c_{j}^{b}$, $u_{j,i,e_{j,i}^{b}\oplus 1}^{b}=r_{j,e_{j,i}^{b}\oplus 1}^{b}\oplus\mathsf{Enc}(pk,\perp)$, then sends $u_{j,i,0}^{b},u_{j,i,1}^{b}$ to $P_{i}$.
+$P_{j}$ computes $u_{j,i,e_{j,i}^{b}}^{b}=r_{j,i,e_{j,i}^{b}}^{b}\oplus c_{j}^{b}$, $u_{j,i,e_{j,i}^{b}\oplus 1}^{b}=r_{j,i,e_{j,i}^{b}\oplus 1}^{b}\oplus\mathsf{Enc}(pk,\perp)$, then sends $u_{j,i,0}^{b},u_{j,i,1}^{b}$ to $P_{i}$.
           $P_{i}$ defines $v_{j,i}^{b}=u_{j,i,e_{i,j}^{b}}^{b}\oplus r_{i,j}^{b}$ and sends ${v}_{j,i}^{\prime b}=\mathsf{ReRand}(pk,v_{j,i}^{b})$ to $P_{j}$.
@@
         Messages decryptions and shufflings.
+        1. (a)
           $P_{1}$ samples $\pi_{1}:[(m-1)B]\to[(m-1)B]$ and computes $\vec{\mathsf{ct}}_{1}^{\prime\prime}=\pi_{1}(\vec{{\mathsf{ct}}_{1}^{\prime}})$.
 $P_{1}$ sends $\vec{\mathsf{ct}}_{1}^{\prime\prime}$ to $P_{2}$.
-      1. (a)
+        2. (b)
           For $2\leq j\leq m,1\leq i\leq(m-1)B$: $P_{j}$ computes ${\mathsf{ct}}_{j}^{i}=\mathsf{ParDec}(sk_{j},{\mathsf{ct}}_{j-1}^{\prime\prime i})$, $pk_{A_{j}}=pk_{1}\cdot\prod_{d=j+1}^{m}pk_{d}$, and ${\mathsf{ct}}_{j}^{\prime i}=\mathsf{ReRand}(pk_{A_{j}},{\mathsf{ct}}_{j}^{i})$.
@@
 If $j\neq m$, $P_{j}$ sends $\vec{\mathsf{ct}}_{j}^{\prime\prime}$ to $P_{j+1}$; else, $P_{m}$ sends $\vec{\mathsf{ct}}_{m}^{\prime\prime}$ to $P_{1}$.
-      2. (b)
+        3. (c)
           For $1\leq i\leq(m-1)B$: $P_{1}$ computes $\mathsf{pt}_{i}=\mathsf{Dec}(sk_{1},{\mathsf{ct}}_{m}^{\prime\prime i})$.
@@
     Figure 12: Our PK-MPSU $\Pi_{\mathsf{PK\text{-}MPSU}}$
-Remark 2
-When instantiating our PK-MPSU framework with EC MKR-PKE, the element space has to be set as EC points accordingly, which may limit its usage in a wide range of applications.
-We argue that the resulting PK-MPSU protocol is still useful in scenarios that the element space is exactly EC points.
-We demonstrate this by building the first multi-party private ID protocol (as described in Appendix 0.G), in which $\Pi_{\mathsf{PK\text{-}MPSU}}$ for EC point elements is used as a core building block.
-7 Theoretical Comparison
-In this section, we compare the construction of mq-ssPMT in [LG23] (mq-ssPMT for short) and our alternative construction built on batch ssPMT (batch ssPMT for short) theoretically.
-Then we compare our two protocols with the recent works [LG23] (which represents the MPSU protocol with the best concrete performance) and [GNT23] (which represents the MPSU protocol with the best asymptotic complexity), respectively.
-We emphasize that both of the two protocols are not in the standard semi-honest model.
-7.1 Comparison Between mq-ssPMT and Batch ssPMT
-Liu and Gao adapt the SKE-based mq-RPMT in [ZCL+23] to construct mq-ssPMT.
-The most expensive part of their construction is the secret-shared oblivious decryption-then-matching (ssVODM) [ZCL+23] protocol, which is to implement a decryption circuit and a comparison circuit by the GMW protocol [GMW87].
-In total, the ssVODM protocol requires $(T+l-\log n-1)n$ AND gates, where $l$ is the bit length of set elements, and $T$ is the number of AND gates in the SKE decryption circuit and is set to be considerably large ($\approx 600$) according to their paper.
-The costs of the batch ssPMT of size $B=1.27n$ [^8] We use stash-less Cuckoo hashing [PSTY19] with 3 hash functions, where $B=1.27n$.
-consist of the costs of batch OPPRF of size $1.27n$ and the costs of $1.27n$ instances of ssPEQT, where the batch OPPRF can be instantiated by extremely fast specialized protocol and the ssPEQT is implemented by the GMW protocol.
-Moreover, the state-of-the-art batch OPPRF construction [CGS22, RS21, RR22] can achieve linear computation and communication complexity with respect to $n$ In the ssPEQT, the parties engage in $1.27(\gamma-1)n$ AND gates, where $\gamma$ is the output length of OPPRF.
-Note that in the typical setting where $n\leq 2^{24},l\leq 128,\gamma\leq 64$, we have $(T+l-\log n-1)n\gg 1.27(\gamma-1)n$, which means that the number of AND gates desired involved in batch ssPMT is far smaller than mq-ssPMT.
-Therefore, the construction built on batch ssPMT greatly reduces the dependency on general 2PC and significantly decreases both computation and communication complexity by a considerable factor, compared to the mq-ssPMT (For a finer-grained asymptotic analysis of mq-ssPMT and batch ssPMT, refer to Appendix 0.F.4 and Appendix 0.F.1).
-7.2 Comparison Between LG Protocol and Our SK-MPSU
-We provide a comprehensive complexity analysis for the state-of-the-art LG protocol in Appendix 0.F.4.
-In Table 2, we present a comparison of the theoretical computation and communication complexity for each party in both offline and online phases between LG protocol and our SK-MPSU.
-Computation Communication
-Offline Online Offline Online
-Leader Client Leader Client
- [LG23] $(T+l+m)mn\log n$ $(T+l+m)mn$ $(T+l)mn$ $\lambda m^{2}n\log n$ $(T+l)mn+lm^{2}n$ $(T+l)mn$
-Ours $(\gamma+m)mn\log n$ $m^{2}n$ $\lambda m^{2}n\log n$ $\gamma mn+lm^{2}n$ $(\gamma+l+m)mn$
-Table 2: Asymptotic communication (bits) and computation costs of LG protocol and our SK-MPSU in the offline and online phases.
-$n$ is the set size.
-$m$ is the number of participants.
-$\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$.
-$T$ is the number of AND gate in an SKE decryption circuit, $T\approx 600$.
-$l$ is the bit length of input elements.
-$\gamma$ is the output length of OPPRF.
-$t$ is the noise weight in dual LPN, $t\approx 128$.
-As depicted in Table 2, our offline communication complexity is comparable to theirs, our offline computation complexity and online computation and communication complexity of the leader are superior to theirs.
-Nevertheless, our online computation and communication complexity of client seem to be surpassed by theirs, as ours are quadratic in the number of parties $m$, while their is linear.
-We argue that it is because that in order to achieve security against arbitrary collusion, our protocol introduces extra overhead (from mss-ROT).
-Even so, for most applications when $m$ is moderate, and our online computation and communication costs still outperform theirs.
-This shall give the credit to our speedup by replacing their mq-ssPMT with batch ssPMT.
-7.3 Comparison Between [GNT23] and Our PK-MPSU
-In Table 3, we present a comparison of the theoretical computation and communication complexity for each party in both offline and online phases between  [GNT23] and our PK-MPSU.
-Computation Communication
-Offline Online Offline Online
- [GNT23] $\gamma mn\log n(\log n/\log\log n)$ $mn(\log n/\log\log n)$ $t\lambda m\log n(\log n/\log\log n)$ $(\gamma+\lambda)mn(\log n/\log\log n)$
-Ours $\gamma mn\log n$ $mn$ $t\lambda m\log n$ $(\gamma+\lambda)mn$
-Table 3: Asymptotic communication (bits) and computation costs of [GNT23] and our PK-MPSU protocol in the offline and online phases.
-In the offline phase, the computation is composed of symmetric-key operations; In the online phase, the computation is composed of public-key operations since we ignore symmetric-key operations.
-$n$ is the set size.
-$m$ is the number of participants.
-$\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$.
-$l$ is the bit length of input elements.
-$\gamma$ is the output length of OPPRF.
-$t$ is the noise weight in dual LPN, $t\approx 128$.
-We conclude that the complexity of our PK-MPSU surpasses theirs in all respects, including the computation and communication complexity of the leader and clients in the offline and online phases, as depicted in the Table 3.
     7 Performance Evaluation
-In this section, we provide implementation details and experimental results for our SK-MPSU and PK-MPSU protocols.
-Most previous works [KS05, Fri07, BA12, GNT23] lack open-source codes.
-The work of [VCE22] shows fairly poor performance in comparison with the state-of-the-art LG protocol.
-Therefore, we only compare our works with LG protocol whose complete implementation is available on https://github.com/lx-1234/MPSU.
-In the implementation of LG protocol, they designate one party to generate share correlations in the offline phase and store them as local files, so that other parties can read these files and consume these share correlations in the online phase [^9] Refer to the function ShareCorrelation::generate() in ShareCorrelationGen.cpp of the repository https://github.com/lx-1234/MPSU.git.
-This implementation is not faithful to the protocol specifications, which is to let the parties pairwise execute share translation protocols [CGP20].
-It has two pitfalls: (1) Their code does not support distributed execution.
-(2) It would lead to serious information leakage.
-Concretely, if it is the leader who generates share correlations, then it can learn the party to whom each element in the union belongs (which is the same security problem addressed by the execution of multi-party secret shuffle as we mentioned in the technical overview).
-If it is a client who generates share correlations, then it can learn the union.
-Moreover, their implementation gives incorrect results in several test cases when the set size $n$ is quite large (cf.
-Table 4).
-As a result, their code cannot be considered as a correct or secure implementation for MPSU.
-To conduct a fair comparison, we replace their implementation of share correlation generation with our correct one.
+    In this section, we provide the implementation details and experimental results for our works.
+For most previous works [KS05, Fri07, BA12, GNT23] do not provide open-source code, and the protocol in [VCE22] shows fairly inferior performance in comparison with the state-of-the-art LG, here we only compare our protocols with LG, whose implementation is available on https://github.com/lx-1234/MPSU.
+    During our experiments, we identify several issues with the LG implementation:
+    1. 1.
+      The code is neither a complete nor a secure implementation for MPSU, as it lacks the offline share correlation generation of the multi-party secret-shared shuffle protocol [EB22].
+Instead of adhering to the protocol specifications, the implementation designates one party to generate and store some “fake” share correlations as local files during the offline phase, while the other parties read these files and consume the share correlations in the online phase [^11] Refer to the function ShareCorrelation::generate() in ShareCorrelationGen.cpp.
+This results in two serious consequences:
+      • The distributed execution is not supported.
+      • The security of multi-party secret-shared shuffle is compromised, leading to information leakage.
+      To conduct a fair and complete comparison, we integrated our correct version of the share correlation generation implementation into their code.
+    2. 2.
+      The code is not a correct implementation, as it produces incorrect results when the set size $n$ increases beyond a certain threshold.
+We figure out that this issue arises from the use of a fixed parameter [^12] The variable batchsize in line 30 of circuit/TripleGen.cpp for generating Beaver Triples in the offline phase, which leads to an insufficient number of Beaver Triples for the online phase when $n$ is large.
+We modified their code to ensure correct execution and mark the cases where the original code fails with *.
     7.1 Experimental Setup
-We run LG and our two protocols on Ubuntu 22.04 with a single Intel i7-13700 2.10 GHz CPU (16 physical cores) and 64 GB RAM.
+    We conduct our experiments on an Alibaba Cloud virtual machine with Intel(R) Xeon(R) 2.70GHz CPU (32 physical cores) and 128 GB RAM.
 We emulate the two network connections using Linux tc command.
@@
 In the WAN setting, the bandwidth is set to be 400 Mbps with 80 ms RTT latency.
-We record the running time as the maximal time from protocol begin to end, including messages transmission time.
-We compute the communication costs of the leader as the sum of the data it sent and received.
-For a fair comparison, we stick to the following setting for all protocols being evaluated:
+We measure the running time as the maximal time from protocol begin to end, including messages transmission time, and the communication costs as the total data that leader sent and received.
+For a fair comparison, we stick to the following settings for all protocols:
     • We set the computational security parameter $\kappa=128$ and the statistical security parameter $\lambda=40$.
     • We test the balanced scenario by setting all $m$ input sets to be of equal size.
-In LG protocol and our SK-MPSU, each party holds $n$ 64-bit strings.
-In our PK-MPSU, we assume that each party holds $n$ elements encoded as elliptic curve points (We provide two versions of the implementation of our PK-MPSU, including encoding elements in uncompressed form and in compressed form respectively).
-  • For each party, we use $m-1$ threads to interact with all other parties simultaneously and 4 threads to perform share correlation generation (in LG protocol and our SK-MPSU), Beaver Triple generation (in all three), parallel SKE encryption (in LG protocol), ciphertext rerandomization and partial decryption (in our PK-MPSU).
+In LG and our SK-MPSU, each party holds $n$ 64-bit strings.
+In our PK-MPSU, each party holds $n$ elements encoded as EC points in compressed form.
+    • Each party uses $m-1$ threads to interact simultaneously with all other parties and 4 threads to perform share correlation generation (in LG and our SK-MPSU), Beaver Triple generation (in all three), parallel SKE encryption (in LG), ciphertext rerandomization and partial decryption (in our PK-MPSU).
     7.2 Implementation Details
@@
     7.3 Experimental Results
-We conduct an extensive experiment for the numbers of parties $\{3,4,5,7,9,10\}$ and a wide range of set sizes $\{2^{6},2^{8},2^{10},2^{12},2^{14},2^{16},2^{18},2^{20}\}$ in the LAN and WAN settings.
-We compare the performance of the protocols from four dimensions: online and total running time, and online and total communication costs.
-The results of online and total running time are depicted in Table 4.
-The results of online and total communication costs are depicted in Table 5.
+    We conduct extensive experiments across various numbers of parties $\{3,4,5,7,9,10\}$ and a wide range of set sizes $\{2^{6},2^{8},2^{10},2^{12},2^{14},2^{16},2^{18},2^{20}\}$ in the LAN and WAN settings.
+The performance of protocols is evaluated from four dimensions: online and total running time, and online and total communication costs.
+The results for running time / communication costs are depicted in Table 2.
+As we can see in the table, our protocols outperform LG in all the case studies, even with enhanced security.
+  • Online computation improvement.
+Our SK-MPSU achieves a 3.910.0×3.9-10.0$\times$3.9 - 10.0 × speedup compared to LG in the LAN setting, and a $\times1$ 2.1 × speedup compared to LG in the WAN setting.
+For example, in the online phase to compute the union of $2^{20}$-size sets among 3 parties, our SK-MPSU runs in 4.4 seconds (LAN) and 33.1 seconds (WAN), which is 5.8×5.8$\times$5.8 × and 1.64×1.64$\times$1.64 × faster than LG that takes 25.2 seconds (LAN) and 54.1 seconds (WAN), respectively.
+  • Total computation improvement.
+Our SK-MPSU achieves a 1.27.8×1.2-7.8$\times$1.2 - 7.8 × speedup compared to LG in the LAN setting.
+Our PK-MPSU achieves a 1.24.0×1.2-4.0$\times$1.2 - 4.0 × speedup compared to LG in the WAN setting.
+For example, to compute the union of $2^{20}$-size sets among 3 parties, our SK-MPSU runs in 96.2 seconds overall (LAN), which is 6.4×6.4$\times$6.4 × faster than LG that takes 610.8 seconds.
+To compute the union of $2^{18}$-size sets among 9 parties, our PK-MPSU runs in 2304 seconds overall (WAN), which is 2.8×2.8$\times$2.8 × faster than LG that takes 6381 seconds.
+  • Online communication improvement.
+The online communication costs of our SK-MPSU is 1.24.9×1.2-4.9$\times$1.2 - 4.9 × smaller than LG.
+For example, in the online phase to compute the union of $2^{20}$-size sets among 3 parties, our SK-MPSU requires 455.6 MB communication, which is a 2.0×2.0$\times$2.0 × improvement of LG that requires 917.0 MB communication.
+  • Total communication improvement.
+The total communication costs of our PK-MPSU is 3.036.5×3.0-36.5$\times$3.0 - 36.5 × smaller than LG.
+For example, to compute the union of $2^{18}$-size sets among 9 parties, our PK-MPSU requires 960.1 MB communication, a 34.8×34.8$\times$34.8 × improvement of LG that requires 33360 MB.
+    In conclusion, our protocols exhibit distinct advantages and are suitable for different scenarios:
+    • Our SK-MPSU: This protocol excels in online performance.
+Additionally, it also outperforms other protocols in nearly all cases in the LAN setting.
+Notably, the online phase takes only $4.4$ seconds for 3 parties with sets of $2^{20}$ items each, and $7$ seconds for 5 parties with sets of $2^{20}$ items each in the LAN setting.
+Therefore, our SK-MPSU is the optimal choice when online performance is the primary concern, or in high-bandwidth networks.
+    • Our PK-MPSU: This protocol has the best total performance in the WAN setting, so it is preferred when the total cost is crucial (e.g.
+in the small-sets setting), at the same time in low-bandwidth networks.
+Note that our PK-MPSU is the only protocol with the capability of running among 10 parties with sets of $2^{18}$ items each in our experiments, which sufficiently indicates the superiority of its linear complexities.
     Sett. $m$ Protocol Set size $n$
-Online time (seconds) Total time (seconds)
+    Online Total = Offline + Online
     $2^{6}$ $2^{8}$ $2^{10}$ $2^{12}$ $2^{14}$ $2^{16}$ $2^{18}$ $2^{20}$ $2^{6}$ $2^{8}$ $2^{10}$ $2^{12}$ $2^{14}$ $2^{16}$ $2^{18}$ $2^{20}$
-LAN 3 LG23 0.030 0.038 0.056 0.106 0.286 1.024 3.899 $\times$ 0.631 0.660 0.822 1.188 4.709 22.62 95.98 $\times$
-Our SK 0.004 0.005 0.006 0.012 0.033 0.138 0.749 3.641 0.232 0.254 0.282 0.344 0.824 3.042 14.32 119.2
-Our PK 0.061 0.134 0.406 1.466 5.452 22.29 87.23 357.3 0.086 0.163 0.440 1.543 5.753 23.55 92.74 389.6
-Our PK 0.083 0.200 0.562 2.039 8.030 31.53 127.4 514.3 0.108 0.229 0.596 2.116 8.331 32.79 133.0 546.6
-4 LG23 0.050 0.055 0.063 0.117 0.336 1.274 5.572 $\times$ 0.948 0.970 1.098 1.744 6.939 32.12 135.4 $\times$
-Our SK 0.006 0.007 0.008 0.017 0.049 0.202 1.043 4.905 0.442 0.458 0.489 0.665 1.700 7.167 36.16 344.71
-Our PK 0.125 0.243 0.734 2.642 10.35 40.83 166.7 -- 0.161 0.280 0.784 2.760 11.00 43.51 178.3 --
-Our PK 0.160 0.346 1.033 3.936 15.50 61.19 246.3 -- 0.196 0.383 1.083 4.054 16.16 63.87 257.9 --
-5 LG23 0.064 0.066 0.077 0.133 0.383 1.509 7.613 -- 1.325 1.355 1.419 2.680 10.82 48.95 213.6 --
-Our SK 0.009 0.010 0.013 0.021 0.061 0.290 1.478 -- 0.677 0.708 0.760 1.114 3.260 13.10 62.35 --
-Our PK 0.170 0.401 1.229 4.191 16.41 64.84 262.5 -- 0.211 0.443 1.273 4.348 17.13 67.89 276.3 --
-Our PK 0.192 0.509 1.611 6.226 24.62 98.04 400.0 -- 0.233 0.551 1.655 6.383 25.33 101.1 413.8 --
-7 LG23 0.099 0.111 0.118 0.233 0.676 3.684 14.45 -- 2.356 2.515 2.752 5.247 23.77 103.6 452.1 --
-Our SK 0.014 0.017 0.021 0.040 0.109 0.545 2.701 -- 1.312 1.357 1.510 2.336 8.368 37.17 146.0 --
-Our PK 0.273 0.649 2.225 8.325 31.62 125.7 521.2 -- 0.338 0.715 2.314 8.695 33.02 131.7 552.7 --
-Our PK 0.339 1.031 3.230 12.25 48.71 195.2 798.1 -- 0.404 1.097 3.319 12.62 50.11 201.3 829.6 --
-9 LG23 0.136 0.142 0.18 0.49 1.584 6.121 -- -- 4.169 4.251 4.716 9.418 44.19 191.6 -- --
-Our SK 0.025 0.029 0.034 0.058 0.186 0.972 -- -- 2.183 2.272 2.557 5.085 18.93 80.01 -- --
-Our PK 0.429 1.064 3.553 13.23 51.74 206.8 -- -- 0.495 1.133 3.660 13.79 54.37 217.9 -- --
-Our PK 0.567 1.558 5.346 22.65 80.03 322.5 -- -- 0.633 1.627 5.453 23.21 82.66 333.6 -- --
-10 LG23 0.161 0.167 0.215 0.531 2.123 8.301 -- -- 4.943 5.178 5.604 12.82 59.17 252.5 -- --
-Our SK 0.03 0.033 0.042 0.073 0.234 1.219 -- -- 2.745 2.883 3.290 6.733 25.12 110.7 -- --
-Our PK 0.498 1.338 4.348 16.35 63.54 260.6 -- -- 0.581 1.435 4.492 17.03 66.77 275.4 -- --
-Our PK 0.677 1.913 6.552 25.11 98.94 404.3 -- -- 0.760 2.010 6.696 25.79 102.2 419.1 -- --
-WAN 3 LG23 4.510 4.516 4.521 4.979 6.295 11.87 17.98 $\times$ 11.87 13.04 15.00 18.33 28.49 85.85 247.6 $\times$
-Our SK 2.413 2.570 2.580 3.402 4.129 5.241 10.91 31.83 8.591 9.731 12.81 16.63 25.56 56.39 192.6 826.3
-Our PK 4.559 4.952 5.787 7.481 13.91 37.62 131.90 505.93 5.300 5.698 6.694 8.420 15.22 39.86 138.5 538.4
-Our PK 4.571 4.662 5.643 7.796 14.82 42.08 150.5 593.6 5.312 5.408 6.550 8.735 16.13 44.33 157.1 626.1
-4 LG23 5.789 5.876 6.293 6.956 7.587 13.58 23.32 $\times$ 17.10 19.64 26.64 32.29 49.43 139.9 490.9 $\times$
-Our SK 3.376 3.536 3.553 4.408 5.919 8.448 20.65 64.01 11.81 14.41 20.42 27.29 48.65 132.1 499.2 2283
-Our PK 6.231 7.181 8.962 11.88 23.57 69.76 251.4 -- 7.374 8.327 10.27 13.22 25.31 73.35 264.1 --
-Our PK 6.264 6.436 8.256 12.14 25.92 81.50 294.2 -- 7.407 7.582 9.565 13.48 27.66 85.08 306.9 --
-5 LG23 7.556 7.794 7.902 9.102 10.63 15.18 34.82 -- 23.26 26.46 35.49 47.53 83.94 250.6 969.6 --
-Our SK 4.501 4.583 5.217 5.955 7.936 12.21 32.14 -- 18.11 21.85 30.93 45.68 91.47 278.9 1112 --
-Our PK 8.050 9.745 11.31 16.08 35.07 108.2 388.9 -- 9.593 11.31 13.03 17.84 37.24 112.5 404.4 --
-Our PK 8.072 8.869 10.91 10.84 42.83 129.1 468.9 -- 9.615 10.43 12.63 18.85 45.00 133.3 484.4 --
-7 LG23 9.722 10.44 11.25 11.67 12.71 22.73 68.51 -- 34.35 45.68 65.99 95.61 203.1 707.2 2881 --
-Our SK 7.241 7.414 8.121 9.808 14.52 22.93 69.19 -- 35.65 46.73 62.48 99.11 234.0 817.7 3399 --
-Our PK 11.44 14.77 17.55 25.72 61.13 201.6 745.7 -- 13.79 17.14 20.10 28.30 64.28 209.3 779.2 --
-Our PK 11.54 14.11 15.89 29.37 70.84 249.3 921.8 -- 13.89 16.48 18.43 31.96 73.98 257.0 955.3 --
-9 LG23 11.07 11.41 13.94 14.91 15.86 34.82 -- -- 55.36 74.52 103.4 166.2 413.7 1531 -- --
-Our SK 10.62 10.81 11.75 13.88 21.07 39.73 -- -- 57.91 88.27 113.7 189.5 505.3 1881 -- --
-Our PK 16.30 19.45 24.87 37.99 93.69 327.8 -- -- 19.45 22.64 28.21 41.33 97.85 341.3 -- --
-Our PK 15.65 19.58 22.33 42.38 109.7 396.3 -- -- 18.80 22.76 25.68 45.73 113.8 409.9 -- --
-10 LG23 12.81 12.99 16.29 16.83 17.50 45.20 -- -- 66.79 91.10 137.2 126.4 579.0 2205 -- --
-Our SK 12.56 12.76 13.60 16.27 25.20 49.94 -- -- 70.02 113.8 140.6 250.8 693.0 2644 -- --
-Our PK 19.48 22.71 28.51 43.45 112.5 392.0 -- -- 23.06 26.32 32.28 47.28 117.2 408.5 -- --
-Our PK 18.28 21.67 26.72 50.03 134.4 487.3 -- -- 21.87 25.28 30.49 53.86 139.1 503.8 -- --
-[…]
-m Protocol Set Size n
-Online communication (MB) Total communication (MB)
-$2^{6}$ $2^{8}$ $2^{10}$ $2^{12}$ $2^{14}$ $2^{16}$ $2^{18}$ $2^{20}$ $2^{6}$ $2^{8}$ $2^{10}$ $2^{12}$ $2^{14}$ $2^{16}$ $2^{18}$ $2^{20}$
-3 LG23 0.157 0.284 0.962 3.662 14.43 57.58 229.8 $\times$ 6.311 7.904 13.37 32.58 114.6 474.3 2054 $\times$
-Our SK 0.031 0.111 0.426 1.687 6.778 27.83 112.5 455.0 2.541 3.582 8.529 30.90 132.7 588.7 2615 11533
-Our PK 1.854 2.147 3.293 7.984 26.69 102.4 406.8 1634 2.123 2.489 3.711 8.474 27.30 103.0 408.8 1640
-Our PK 1.815 1.983 2.655 5.418 16.36 60.78 239.3 959.5 2.084 2.325 3.073 5.908 16.92 61.41 241.3 966.1
-4 LG23 0.242 0.449 1.536 5.868 23.15 92.38 368.6 $\times$ 9.591 12.44 23.89 67.25 253.7 1074 4677 $\times$
-Our SK 0.060 0.218 0.848 3.376 13.57 55.44 223.9 904.4 3.943 6.399 17.01 67.12 294.7 1315 5851 25819
-Our PK 2.781 3.212 4.939 11.98 40.04 153.6 610.2 -- 3.185 3.725 5.566 12.71 40.88 154.5 613.2 --
-Our PK 2.723 2.975 3.983 8.130 24.54 91.20 359.0 -- 3.127 3.488 4.610 8.127 25.38 92.12 362.0 --
-5 LG23 0.330 0.630 2.173 8.325 32.86 131.2 523.5 -- 13.12 17.75 36.48 110.7 435.2 1869 8172 --
-Our SK 0.097 0.361 1.411 5.628 22.63 92.18 372.0 -- 5.511 10.22 30.59 125.3 551.2 2455 10875 --
-Our PK 3.708 4.283 6.585 15.97 53.38 204.8 813.6 -- 4.247 4.966 7.421 16.95 54.51 206.0 817.5 --
-Our PK 3.630 3.967 5.311 10.84 32.72 121.6 478.7 -- 4.169 4.650 6.147 11.82 33.84 122.8 482.6 --
-7 LG23 0.519 1.038 3.635 13.99 55.29 220.8 881.3 -- 19.92 29.90 71.59 243.2 997.9 4327 18930 --
-Our SK 0.198 0.750 2.956 11.82 47.55 193.1 778.6 -- 8.853 18.92 65.23 277.2 1233 5505 24402 --
-Our PK 5.563 6.425 9.878 23.95 80.07 307.1 1220 -- 6.371 7.450 11.13 25.42 81.76 309.0 1226 --
-Our PK 5.445 5.950 7.966 16.25 49.07 182.3 718.0 -- 6.253 6.975 9.220 17.72 50.76 184.2 723.9 --
-9 LG23 0.723 1.509 5.347 20.65 81.72 326.3 -- -- 28.15 44.74 115.0 415.3 1742 7609 -- --
-Our SK 0.333 1.278 5.060 20.27 81.54 330.5 -- -- 13.90 33.27 120.9 520.0 2310 10268 -- --
-Our PK 7.417 8.566 13.17 31.94 106.8 409.5 -- -- 8.495 9.933 14.84 33.90 109.0 412.0 -- --
-Our PK 7.26 7.933 10.62 21.67 65.43 243.1 -- -- 8.338 9.300 12.29 23.63 67.68 245.6 -- --
-10 LG23 0.831 1.768 6.296 24.36 96.43 385.1 -- -- 32.32 54.34 148.7 549.4 2314 10081 -- --
-Our SK 0.413 1.594 6.322 25.34 101.9 412.9 -- -- 16.12 40.87 151.2 656.4 2921 12989 -- --
-Our PK 8.344 9.637 14.82 35.93 120.1 460.7 -- -- 9.556 11.18 16.70 38.13 122.6 463.5 -- --
-Our PK 8.168 8.925 11.95 24.38 73.61 273.5 -- -- 9.380 10.46 13.83 26.59 76.14 276.3 -- --
-[…]
-As we can see in the tables, our protocols outperform LG protocol in almost all the case studies.
-Especially, our SK-MPSU protocol performs best in terms of online running time in LAN setting, online communication costs for a relatively small number of parties, and total running time for medium and large sets in LAN setting among the three protocols.
-We observe a 4.99.3×4.9-9.3$\times$4.9 - 9.3 × improvement in online running time and a 1.78.5×1.7-8.5$\times$1.7 - 8.5 × improvement in total running time in the LAN setting.
-It is worth mentioning that our SK-MPSU takes only $3.6$ seconds for 3 parties with sets of $2^{20}$ items each, and $4.9$ seconds for 4 parties with sets of $2^{20}$ items each in online phase, while the implementation of LG protocol fails to run correctly in these cases.
-On the other hand, our PK-MPSU protocol performs best in terms of total running time in WAN setting and total communication costs among the three protocols.
-It shows a 3.036.5×3.0-36.5$\times$3.0 - 36.5 × improvement in overall communication, and a 1.85.4×1.8-5.4$\times$1.8 - 5.4 × improvement in overall running time in the WAN setting, compared to LG protocol.
-For the cases where both the number of parties and the set size are relatively large and the network environment is bandwidth-constrained, LG protocol marginally surpasses ours in terms of online running time, but its speedup is only up to 1.4×1.4$\times$1.4 ×.
-This result aligns with our theoretical analysis in Section 7.2.
-Nevertheless, when each party’s set size increases to sufficiently large ($2^{20}$), their implementation gives wrong results.
+    3 LG 0.050 0.058 0.069 0.143 0.425 1.582 6.219 25.16 0.982 1.016 1.098 1.767 5.650 32.42 153.8 610.8
+    Our SK 0.005 0.007 0.009 0.017 0.050 0.213 1.005 4.352 0.361 0.386 0.395 0.489 1.170 4.534 19.77 96.23
+    3 Our PK 0.092 0.237 0.822 3.176 12.69 51.17 205.7 829.1 0.117 0.273 0.868 3.270 13.08 53.05 215.0 872.6
+    4 LG 0.069 0.076 0.093 0.162 0.478 1.637 6.375 25.79 1.292 1.358 1.514 2.432 8.716 45.37 197.2 762.9
+    Our SK 0.008 0.010 0.013 0.023 0.071 0.286 1.393 5.645 0.639 0.665 0.696 0.917 2.597 10.94 50.25 237.8
+    4 Our PK 0.159 0.465 1.570 6.136 24.60 44.91 398.4 1604 0.196 0.505 1.643 6.299 25.45 48.95 417.5 1694
+    5 LG 0.107 0.109 0.126 0.190 0.541 1.998 7.588 31.33 1.939 1.993 2.202 3.408 12.54 65.04 289.9 1152
+    Our SK 0.012 0.013 0.017 0.030 0.087 0.368 1.714 7.003 0.914 0.964 1.023 1.558 4.551 18.71 85.79 373.4
+    5 Our PK 0.232 0.693 2.487 9.858 39.33 158.1 637.5 2816 0.292 0.796 2.547 10.09 40.25 162.2 655.0 2979
+    7 LG 0.154 0.165 0.174 0.281 0.795 2.894 10.92 -- 3.484 3.583 3.921 5.84 23.42 111.2 484.5 --
+    Our SK 0.019 0.021 0.028 0.048 0.156 0.607 2.817 -- 2.051 2.079 2.214 3.470 12.92 56.50 245.6 --
+    7 Our PK 0.463 1.365 5.030 19.43 77.22 310.0 1247 -- 0.550 1.469 5.158 19.73 78.59 316.3 1276 --
+    9 LG 0.222 0.226 0.234 0.417 1.216 4.449 17.29 -- 5.501 5.612 6.249 10.27 41.69 182.7 793.3 --
+    Our SK 0.027 0.031 0.039 0.075 0.230 0.970 4.293 -- 3.363 3.402 3.895 7.161 30.09 126.3 678.8 --
+    9 Our PK 0.764 2.271 8.074 31.64 126.7 507.0 2039 -- 0.880 2.375 8.238 32.19 128.7 515.5 2080 --
+    10 LG 0.228 0.243 0.276 0.514 1.479 5.467 -- -- 6.736 6.846 8.335 13.56 55.58 238.3 -- --
+    Our SK 0.031 0.036 0.043 0.088 0.286 1.183 -- -- 3.965 4.232 4.821 9.334 41.97 175.9 -- --
+    Time. (s) LAN 10 Our PK 0.865 2.800 9.944 39.09 155.9 622.7 2503 -- 0.970 2.912 10.11 39.58 158.4 634.6 2556 --
+    3 LG 4.502 4.505 4.522 4.914 6.272 8.744 17.78 54.13 12.46 13.59 15.76 19.52 30.73 78.74 282.8 1188
+    Our SK 2.165 2.166 2.332 3.157 3.734 4.444 9.705 33.10 8.403 9.710 12.62 16.30 25.02 56.88 194.7 801.2
+    3 Our PK 4.419 4.555 5.553 7.984 18.21 59.77 226.3 900.3 5.168 5.306 6.472 8.968 19.65 62.73 237.3 946.1
+    4 LG 5.696 5.880 6.540 7.094 7.323 11.36 23.13 78.81 17.94 20.99 27.63 32.69 50.59 145.7 554.7 2167
+    Our SK 2.967 2.969 3.298 3.976 4.618 6.507 17.10 59.21 11.46 13.93 20.21 26.77 47.49 133.2 520.0 2141
+    4 Our PK 5.622 5.899 7.929 12.17 32.40 113.8 440.9 1761 6.773 7.052 9.25 13.53 34.27 118.9 461.3 1857
+    5 LG 7.385 7.708 8.621 9.198 9.687 14.62 32.83 126.3 23.64 26.82 37.39 48.48 88.32 263.3 1053 4394
+    Our SK 3.768 3.733 4.471 4.800 5.521 8.938 25.95 95.40 17.53 21.10 30.28 44.30 88.15 278.9 1119 4820
+    5 Our PK 6.849 7.928 10.50 17.08 49.75 179.8 703.3 2873 8.424 9.483 12.22 18.86 52.07 185.4 724.5 2980
+    7 LG 9.312 9.833 10.55 11.35 12.14 21.29 66.93 -- 34.92 45.69 66.26 94.89 203.4 705.8 2898 --
+    Our SK 5.373 5.381 6.207 6.644 8.164 17.67 56.894 -- 34.00 44.95 61.34 95.03 228.7 817.4 3506 --
+    7 Our PK 9.504 12.32 15.14 29.73 92.66 348.5 1377.4 -- 11.88 14.71 17.72 32.35 95.86 356.5 1409 --
+    9 LG 11.41 12.21 13.34 14.41 15.09 33.84 115.5 -- 56.65 75.24 104.1 169.1 406.0 1503 6387 --
+    Our SK 6.977 7.068 7.830 8.387 12.24 29.20 105.7 -- 58.81 84.80 107.6 182.0 502.5 1915 8137 --
+    9 Our PK 13.67 18.84 22.96 45.54 148.2 570.9 -- -- 16.87 22.04 26.31 48.98 152.3 581.5 2034 --
+    10 LG 11.77 12.24 15.80 16.49 17.48 45.20 -- -- 66.19 92.22 125.1 219.8 582.1 2179 -- --
+    Our SK 7.780 8.032 8.635 9.203 14.54 38.31 -- -- 71.58 109.2 132.7 242.7 684.0 2687 -- --
+    Time. (s) WAN 10 Our PK 16.32 23.05 28.66 59.79 184.8 708.5 2792 -- 19.90 26.66 32.44 63.66 189.3 720.9 2846 --
+    3 LG 0.157 0.284 0.962 3.662 14.43 57.58 229.8 917.0 6.311 7.904 13.37 32.58 114.6 474.3 2052 8973
+    Our SK 0.032 0.111 0.426 1.690 6.788 27.87 112.7 455.6 2.542 3.582 8.529 30.91 132.7 588.8 2614 11529
+    3 Our PK 1.815 1.983 2.655 5.418 16.36 60.78 239.3 959.5 2.084 2.325 3.073 5.908 16.92 61.41 240.0 960.3
+    4 LG 0.242 0.449 1.536 5.868 23.15 92.38 368.6 1471 9.591 12.44 23.89 67.25 253.7 1074 4674 20419
+    Our SK 0.058 0.204 0.791 3.145 12.81 51.69 208.8 843.7 3.941 6.385 16.96 66.89 293.9 1312 5834 25751
+    4 Our PK 2.723 2.975 3.983 8.126 24.54 101.9 359.0 1439 3.127 3.488 4.610 8.861 25.38 102.8 360.1 1440
+    5 LG 0.330 0.630 2.173 8.325 32.86 131.2 523.5 2090 12.92 17.75 36.48 110.7 435.2 1868 8228 35737
+    Our SK 0.090 0.323 1.257 5.007 20.36 82.11 331.5 1339 5.504 10.18 30.44 124.7 548.9 2445 10832 47635
+    5 Our PK 3.630 3.967 5.311 10.84 32.72 121.6 478.7 1919 4.169 4.650 6.147 11.82 33.84 122.8 480.1 1921
+    7 LG 0.519 1.038 3.635 13.99 55.29 220.8 881.3 -- 19.92 29.90 41.59 243.3 997.9 4326 18924 --
+    Our SK 0.172 0.634 2.489 10.03 40.31 162.5 655.4 -- 8.827 18.80 64.76 275.4 1226 5474 24275 --
+    7 Our PK 5.445 5.950 7.966 16.25 49.07 182.3 718.0 -- 6.253 6.98 9.220 17.72 50.76 184.2 720.1 --
+    9 LG 0.723 1.509 5.347 20.65 81.72 326.3 1303 -- 28.15 44.74 115.0 415.33 1742 7609 33360 --
+    Our SK 0.279 1.046 4.122 16.60 66.76 269.0 1084 -- 13.99 33.04 119.9 516.3 2295 10207 45079 --
+    9 Our PK 7.260 7.933 10.62 21.67 65.43 243.1 957.3 -- 8.338 9.300 12.29 23.63 67.68 245.6 960.1 --
+    10 LG 0.831 1.768 6.296 24.36 96.43 385.1 -- -- 32.32 54.34 148.7 549.4 2314 10081 -- --
+    Our SK 0.341 1.288 5.086 20.48 82.39 332.0 -- -- 16.21 40.56 149.9 651.6 2901 12908 -- --
+    Comm. (MB) 10 Our PK 8.168 8.925 11.95 24.38 73.61 273.5 1077 -- 9.380 10.46 13.83 26.59 76.14 276.3 1080 --
+    Table 2: Online and total running time / communication costs of LG and our protocols in LAN and WAN settings.
+$m$ is the number of parties.
+Our SK /PK denotes our SK-MPSU / PK-MPSU protocol.
+Cells with * denotes trials that the original LG code will give wrong results.
+Cells with -- denotes trials that ran out of memory.
+The best protocol within a setting is marked in blue.
+    • Acknowledgements. We would like to thank Jiahui Gao for the clarification of their work.
+  • References
+    • [BA12] Marina Blanton and Everaldo Aguiar.
+Private and oblivious set and multiset operations.
+In 7th ACM Symposium on Information, Compuer and Communications Security, ASIACCS 2012, pages 40–41.
+ACM, 2012.
+    • [BC23] Dung Bui and Geoffroy Couteau.
+Improved private set intersection for sets with small entries.
+In Public-Key Cryptography - PKC 2023 - 26th IACR International Conference on Practice and Theory of Public-Key Cryptography, volume 13941 of Lecture Notes in Computer Science, pages 190–220.
+Springer, 2023.
+    • [BCG+19a] Elette Boyle, Geoffroy Couteau, Niv Gilboa, Yuval Ishai, Lisa Kohl, Peter Rindal, and Peter Scholl.
+Efficient two-round OT extension and silent non-interactive secure computation.
+In Proceedings of the 2019 ACM SIGSAC Conference on Computer and Communications Security, CCS 2019, pages 291–308.
+ACM, 2019.
+    • [BCG+19b] Elette Boyle, Geoffroy Couteau, Niv Gilboa, Yuval Ishai, Lisa Kohl, and Peter Scholl.
+Efficient pseudorandom correlation generators: Silent OT extension and more.
+In Advances in Cryptology - CRYPTO 2019.
+Springer, 2019.
+    • [Bea91] Donald Beaver.
+Efficient multiparty protocols using circuit randomization.
+In Advances in Cryptology - CRYPTO ’91, 11th Annual International Cryptology Conference, volume 576 of Lecture Notes in Computer Science, pages 420–432.
+Springer, 1991.
+    • [BKM+20] Prasad Buddhavarapu, Andrew Knox, Payman Mohassel, Shubho Sengupta, Erik Taubeneck, and Vlad Vlaskin. Private matching for compute. 2020. https://eprint.iacr.org/2020/599.
+    • [BNOP22] Aner Ben-Efraim, Olga Nissenbaum, Eran Omri, and Anat Paskin-Cherniavsky.
+Psimple: Practical multiparty maliciously-secure private set intersection.
+In ASIA CCS ’22, pages 1098–1112.
+ACM, 2022.
+    • [BPSY23] Alexander Bienstock, Sarvar Patel, Joon Young Seo, and Kevin Yeo.
+Near-Optimal oblivious Key-Value stores for efficient PSI, PSU and Volume-Hiding Multi-Maps.
+In USENIX Security 2023, pages 301–318, 2023.
+    • [BS05] Justin Brickell and Vitaly Shmatikov.
+Privacy-preserving graph algorithms in the semi-honest model.
+In Advances in Cryptology - ASIACRYPT 2005, 11th International Conference on the Theory and Application of Cryptology and Information Security, volume 3788 of Lecture Notes in Computer Science, pages 236–252.
+Springer, 2005.
+    • [CDG+21] Nishanth Chandran, Nishka Dasgupta, Divya Gupta, Sai Lakshmi Bhavana Obbattu, Sruthi Sekar, and Akash Shah.
+Efficient linear multiparty PSI and extensions to circuit/quorum PSI.
+In CCS ’21, pages 1182–1204.
+ACM, 2021.
+    • [CGP20] Melissa Chase, Esha Ghosh, and Oxana Poburinnaya.
+Secret-shared shuffle.
+In Advances in Cryptology - ASIACRYPT 2020, volume 12493 of Lecture Notes in Computer Science, pages 342–372.
+Springer, 2020.
+    • [CGS22] Nishanth Chandran, Divya Gupta, and Akash Shah. Circuit-psi with linear complexity via relaxed batch OPPRF. Proc. Priv. Enhancing Technol., 2022(1):353–372, 2022.
+    • [CM20] Melissa Chase and Peihan Miao.
+Private set intersection in the internet setting from lightweight oblivious PRF.
+In Advances in Cryptology - CRYPTO 2020, volume 12172 of Lecture Notes in Computer Science, pages 34–63.
+Springer, 2020.
+    • [CO18] Michele Ciampi and Claudio Orlandi.
+Combining private set-intersection with secure two-party computation.
+In Security and Cryptography for Networks - 11th International Conference, SCN 2018, volume 11035 of Lecture Notes in Computer Science, pages 464–482.
+Springer, 2018.
+    • [CZZ+24a] Yu Chen, Min Zhang, Cong Zhang, Minglang Dong, and Weiran Liu.
+Private set operations from multi-query reverse private membership test.
+In Public-Key Cryptography - PKC 2024.
+Springer, 2024.
+    • [CZZ+24b] Yu Chen, Min Zhang, Cong Zhang, Minglang Dong, and Weiran Liu.
+Private set operations from multi-query reverse private membership test.
+In PKC 2024, 2024.
+https://eprint.iacr.org/2022/652.
+    • [EB22] Saba Eskandarian and Dan Boneh.
+Clarion: Anonymous communication from multiparty shuffling protocols.
+In 29th Annual Network and Distributed System Security Symposium, NDSS 2022.
+The Internet Society, 2022.
+    • [ElG85] Taher ElGamal. A public key cryptosystem and a signature scheme based on discrete logarithms. IEEE Transactions on Information Theory, 31:469–472, 1985.
+    • [FIPR05] Michael J.
+Freedman, Yuval Ishai, Benny Pinkas, and Omer Reingold.
+Keyword search and oblivious pseudorandom functions.
+In Theory of Cryptography, Second Theory of Cryptography Conference, TCC 2005, volume 3378 of Lecture Notes in Computer Science, pages 303–324.
+Springer, 2005.
+    • [Fri07] Keith B.
+Frikken.
+Privacy-preserving set union.
+In Applied Cryptography and Network Security, 5th International Conference, ACNS 2007, volume 4521 of Lecture Notes in Computer Science, pages 237–252.
+Springer, 2007.
+    • [GMR+21] Gayathri Garimella, Payman Mohassel, Mike Rosulek, Saeed Sadeghian, and Jaspal Singh.
+Private set operations from oblivious switching.
+In Public-Key Cryptography - PKC 2021, volume 12711 of Lecture Notes in Computer Science, pages 591–617.
+Springer, 2021.
+    • [GMW87] Oded Goldreich, Silvio Micali, and Avi Wigderson.
+How to play any mental game or A completeness theorem for protocols with honest majority.
+In Proceedings of the 19th Annual ACM Symposium on Theory of Computing, 1987, pages 218–229.
+ACM, 1987.
+    • [GNT23] Jiahui Gao, Son Nguyen, and Ni Trieu.
+Toward a practical multi-party private set union.
+Cryptology ePrint Archive, Paper 2023/1930, 2023.
+Version: 20240316:210303, https://eprint.iacr.org/archive/2023/1930/20240316:210303.
+    • [GPR+21] Gayathri Garimella, Benny Pinkas, Mike Rosulek, Ni Trieu, and Avishay Yanai.
+Oblivious key-value stores and amplification for private set intersection.
+In Advances in Cryptology - CRYPTO 2021, volume 12826 of Lecture Notes in Computer Science, pages 395–425.
+Springer, 2021.
+    • [HLS+16] Kyle Hogan, Noah Luther, Nabil Schear, Emily Shen, David Stott, Sophia Yakoubov, and Arkady Yerukhimovich.
+Secure multiparty computation for cooperative cyber risk assessment.
+In IEEE Cybersecurity Development, 2016, pages 75–76.
+IEEE Computer Society, 2016.
+    • [JSZ+22] Yanxue Jia, Shi-Feng Sun, Hong-Sheng Zhou, Jiajun Du, and Dawu Gu. Shuffle-based private set union: Faster and more secure. In USENIX 2022, 2022.
+    • [JSZG24] Yanxue Jia, Shi-Feng Sun, Hong-Sheng Zhou, and Dawu Gu.
+Scalable private set union, with stronger security.
+In 33rd USENIX Security Symposium, USENIX Security 2024.
+USENIX Association, 2024.
+    • [KKRT16] Vladimir Kolesnikov, Ranjit Kumaresan, Mike Rosulek, and Ni Trieu. Efficient batched oblivious PRF with applications to private set intersection. In CCS 2016, pages 818–829. ACM, 2016.
+    • [KMP+17] Vladimir Kolesnikov, Naor Matania, Benny Pinkas, Mike Rosulek, and Ni Trieu.
+Practical multi-party private set intersection from symmetric-key techniques.
+In Proceedings of the 2017 ACM SIGSAC Conference on Computer and Communications Security, CCS 2017, pages 1257–1272.
+ACM, 2017.
+    • [KRTW19] Vladimir Kolesnikov, Mike Rosulek, Ni Trieu, and Xiao Wang.
+Scalable private set union from symmetric-key techniques.
+In Advances in Cryptology - ASIACRYPT 2019, volume 11922 of Lecture Notes in Computer Science, pages 636–666.
+Springer, 2019.
+    • [KS05] Lea Kissner and Dawn Xiaodong Song.
+Privacy-preserving set operations.
+In Advances in Cryptology - CRYPTO 2005, volume 3621 of Lecture Notes in Computer Science, pages 241–257.
+Springer, 2005.
+    • [LG23] Xiang Liu and Ying Gao. Scalable multi-party private set union from multi-query secret-shared private membership test. In Advances in Cryptology - ASIACRYPT 2023. Springer, 2023.
+    • [LV04] Arjen K.
+Lenstra and Tim Voss.
+Information security risk assessment, aggregation, and mitigation.
+In Information Security and Privacy: 9th Australasian Conference, ACISP 2004., volume 3108 of Lecture Notes in Computer Science, pages 391–401.
+Springer, 2004.
+    • [MS13] Payman Mohassel and Seyed Saeed Sadeghian.
+How to hide circuits in MPC an efficient framework for private function evaluation.
+In Advances in Cryptology - EUROCRYPT 2013, volume 7881 of Lecture Notes in Computer Science, pages 557–574.
+Springer, 2013.
+    • [NTY21] Ofri Nevo, Ni Trieu, and Avishay Yanai.
+Simple, fast malicious multiparty private set intersection.
+In Proceedings of the 2021 ACM SIGSAC Conference on Computer and Communications Security, CCS 2021, pages 1151–1165.
+ACM, 2021.
+    • [PR04] Rasmus Pagh and Flemming Friche Rodler. Cuckoo hashing. Journal of Algorithms, 51(2):122–144, 2004.
+    • [PRTY19] Benny Pinkas, Mike Rosulek, Ni Trieu, and Avishay Yanai.
+Spot-light: Lightweight private set intersection from sparse OT extension.
+In Advances in Cryptology - CRYPTO 2019, volume 11694 of Lecture Notes in Computer Science, pages 401–431.
+Springer, 2019.
+    • [PRTY20] Benny Pinkas, Mike Rosulek, Ni Trieu, and Avishay Yanai.
+PSI from paxos: Fast, malicious private set intersection.
+In Advances in Cryptology - EUROCRYPT 2020, volume 12106 of Lecture Notes in Computer Science, pages 739–767.
+Springer, 2020.
+    • [PSSZ15] Benny Pinkas, Thomas Schneider, Gil Segev, and Michael Zohner.
+Phasing: Private set intersection using permutation-based hashing.
+In 24th USENIX Security Symposium, USENIX Security 15, Washington, D.C., USA, August 12-14, 2015, pages 515–530.
+USENIX Association, 2015.
+    • [PSTY19] Benny Pinkas, Thomas Schneider, Oleksandr Tkachenko, and Avishay Yanai.
+Efficient circuit-based PSI with linear communication.
+In Advances in Cryptology - EUROCRYPT 2019, volume 11478 of Lecture Notes in Computer Science, pages 122–153.
+Springer, 2019.
+    • [PSZ14] Benny Pinkas, Thomas Schneider, and Michael Zohner.
+Faster private set intersection based on OT extension.
+In Proceedings of the 23rd USENIX Security Symposium, 2014, pages 797–812.
+USENIX Association, 2014.
+    • [Rab05] Michael O. Rabin. How to exchange secrets with oblivious transfer. IACR Cryptol. ePrint Arch., page 187, 2005.
+    • [Roy22] Lawrence Roy. Softspokenot: Quieter OT extension from small-field silent VOLE in the minicrypt model. In Advances in Cryptology - CRYPTO 2022. Springer, 2022.
+    • [RR22] Srinivasan Raghuraman and Peter Rindal. Blazing fast PSI from improved OKVS and subfield VOLE. In ACM CCS 2022, 2022.
+    • [RRT23] Srinivasan Raghuraman, Peter Rindal, and Titouan Tanguy.
+Expand-convolute codes for pseudorandom correlation generators from LPN.
+In Advances in Cryptology - CRYPTO 2023, volume 14084 of Lecture Notes in Computer Science, pages 602–632.
+Springer, 2023.
+    • [RS21] Peter Rindal and Phillipp Schoppmann.
+VOLE-PSI: fast OPRF and circuit-psi from vector-ole.
+In Advances in Cryptology - EUROCRYPT 2021, volume 12697 of Lecture Notes in Computer Science, pages 901–930.
+Springer, 2021.
+    • [VCE22] Jelle Vos, Mauro Conti, and Zekeriya Erkin. Fast multi-party private set operations in the star topology from secure ands and ors. IACR Cryptol. ePrint Arch., page 721, 2022.
+    • [ZCL+23] Cong Zhang, Yu Chen, Weiran Liu, Min Zhang, and Dongdai Lin.
+Optimal private set union from multi-query reverse private membership test.
+In USENIX 2023, 2023.
+https://eprint.iacr.org/2022/358.
+    • [ZMS+21] Shengnan Zhao, Ming Ma, Xiangfu Song, Han Jiang, Yunxue Yan, and Qiuliang Xu.
+Lightweight threshold private set intersection via oblivious transfer.
+In Wireless Algorithms, Systems, and Applications - 16th International Conference, WASA 2021, volume 12939 of Lecture Notes in Computer Science, pages 108–116.
+Springer, 2021.
     Appendix 0.A Leakage Analysis of [GNT23]
-Gao et al.
-[GNT23] claimed that their MPSU protocol is secure in the presence of arbitrary colluding participants.
-However, we find that their protocol fails to achieve this security, and also requires the same non-collusion assumption as LG protocol.
-First, we give a brief review of their protocol.
-Apart from MKR-PKE, the protocol of [GNT23] utilizes three primary ingredients: 1) The multi-point OPRF [PRTY19, CM20, RS21, RR22], where the sender $\mathcal{S}$ inputs a PRF key $k$, and the receiver $\mathcal{R}$ inputs $n$ elements $\{x_{1},\cdots,x_{n}\}$ and receives the corresponding PRF evaluations $\{F_{k}(x_{1}),\cdots,F_{k}(x_{n})\}$.
-2) The conditional oblivious pseudorandom function (cOPRF), an extension they develop on the multi-point OPRF [^17] In their paper, the definition of cOPRF is the single-point version.
-For ease of illustration, we use the multi-point version., where the sender $\mathcal{S}$ additionally inputs a set $Y$.
-If $x_{i}\notin Y$, $\mathcal{R}$ receives $F_{k}(x_{i})$, else $\mathcal{R}$ receives a random value sampled by $\mathcal{S}$ [^18] In the latest version of [GNT23], they have made a slight adjustment to the functionality of cOPRF.
-However, this adjustment does not change their insecurity.
-Their latest protocol is still vulnerable to our attack.
-3) The membership Oblivious Transfer (mOT), where $\mathcal{S}$ inputs an element $y$ and two messages $u_{0},u_{1}$, while $\mathcal{R}$ inputs a set $X$ and receives $u$, one of $u_{0},u_{1}$.
-If $y\notin X$, $u=u_{0}$, else $u=u_{1}$.
-To illustrate the insecurity of their protocol, we consider a four-party case where the leader $P_{1}$ possesses a set $X_{1}$ of items while the remaining parties $P_{2},P_{3},P_{4}$, each possesses a single item $X_{i}$ = $\{x_{i}\}$ ($i=2,3,4$).
-We assume that $x_{2},x_{3},x_{4}\notin X_{1}$ and $x_{2}\neq x_{3}=x_{4}$ [^19] In their paper, to illustrate the high-level idea of the protocol, they give a four-party case as well, assuming that $x_{2}=x_{3}\notin X_{1}$ and $x_{4}\in X_{1}$, see [GNT23]..
-[…]
-Figure 14: Illustration of the mPSU protocol in [GNT23] for $m=4$ parties.
-$P_{1}$ possesses a set $X_{1}=\{x_{1}^{1},\cdots,x_{1}^{n}\}$ of $n$ items while $P_{i}$ has input set of only one item $x_{i}$ for $i=2,3,4$.
-We assume that $x_{2},x_{3},x_{4}\notin X_{1}$ and $x_{2}\neq x_{3}=x_{4}$.
-$P_{1}$ possesses two set $S,E$ and updates them in each round.
-In the $i$-th round, $P_{1}$ and $P_{i}$ invoke mpOPRF and mOT in sequence.
-$F_{k}(X_{1})$ denotes the set of PRF values on $X_{1}$, i.e., $\{F_{k}(x_{1}^{1}),\cdots,F_{k}(x_{1}^{n})\}$.
-$P_{i}$ invokes cOPRF and mOT with $P_{i+1},\cdots,P_{m}$ concurrently.
-At the end of $i$-th round, $S$ contains the multi-key PRF values on $X_{1}\cup\cdots\cup X_{i+1}$ using keys $k_{2},\cdots,k_{i+1}$, $E$ contains the MKR-PKE ciphertexts of $X_{1}\cup\cdots\cup X_{i+1}\setminus X_{1}$.
-Finally, all parties cooperate in a Shuffle&Decrypt protocol to decrypt $E$ in a random permutation.
-We denote $F_{k_{3}}(F_{k_{2}}(\cdot))$ as $F_{k_{2},k_{3}}(\cdot)$, more precisely, it is in the form of $H(k_{3},H(k_{2},\cdot\|1)\|1)$ in the paper.
-We emphasize that the ciphertext $\mathsf{Enc}(pk,x_{3})$ in Round 1 is different from that in Round 2.
-The latter is the rerandomized ciphertext of the former.
-We write them using the same denotation as this does not affect our illustration of the the attack.
-Changes to sets $S$ and $E$ at the end of each round are marked in blue font.
-We demonstrate the execution process of their MPSU in Figure 14 pertaining to the scenario we describe above.
-In light of space constraints, Figure 14 only covers Round 1 and Round 2 and it omits some invocations of cOPRF and mOT since the potential information leakage already occurs during Round 2 when $P_{1}$ and $P_{3}$ invoke the mOT.
-To elaborate, $P_{1}$ receives $F_{k_{2}}(x_{2})$ from the invocation of mOT with $P_{2}$ in Round 1 and $F_{k_{2},k_{3}}(x_{3})$ from the invocation of mOT with $P_{3}$ in Round 2.
-If $P_{1}$ and $P_{3}$ collude, $P_{1}$ can compute $F_{k_{2},k_{3}}(x_{2})$ using $P_{3}$’s key $k_{3}$ and learn the knowledge of $F_{k_{2},k_{3}}(x_{2})\neq F_{k_{2},k_{3}}(x_{3})$, which indicates $x_{2}\neq x_{3}$.
-Recall that $P_{1}$ is the leader to obtain the union $X_{1}\cup\{x_{2},x_{3}=x_{4}\}$ in the end.
-Therefore, the coalition of $P_{1}$ and $P_{3}$ can deduce that the unique element in the union, not present in their joint sets $X_{1}\cup\{x_{3}\}$ (i.e., $x_{2}$), belongs to $P_{2}$.
-Accordingly, they know that $P_{4}$ possesses the duplicate element with either $P_{1}$ or $P_{3}$.
-Note that in the ideal world, the coalition of $P_{1}$ and $P_{3}$ should not acquire knowledge of the source party of the foreign element in the union or the party possessing the duplicate element.
-Therefore, their protocol at least requires the same assumption as LG protocol, assuming that the leader $P_{1}$ is not in collusion with any other parties.
+    The MPSU protocol in [GNT23] is claimed to be secure in the presence of arbitrary colluding participants.
+However, our analysis would suggest that the protocol fails to achieve this security, and also requires the non-collusion assumption as LG.
+First, we give a brief review of the protocol.
+    Apart from MKR-PKE, their protocol utilizes two new ingredients: 1) The conditional oblivious pseudorandom function (cOPRF), an extension they develop on the OPRF, where the sender $\mathcal{S}$ additionally inputs a set $Y$.
+If $x\notin Y$, $\mathcal{R}$ receives $F_{k}(x)$, else $\mathcal{R}$ receives a random value sampled by $\mathcal{S}$.
+2) The membership Oblivious Transfer (mOT), where $\mathcal{S}$ inputs an element $y$ and two messages $u_{0},u_{1}$, while $\mathcal{R}$ inputs a set $X$ and receives $u$, one of $u_{0},u_{1}$.
+If $y\in X$, $u=u_{0}$, else $u=u_{1}$.
+    To illustrate insecurity of their protocol, we consider a three-party case where $P_{1}$ and $P_{3}$ each possess a single item $X_{1}=\{x_{1}\}$ and $X_{3}=\{x_{3}\}$ respectively, while $P_{2}$ possesses a set $X_{2}$.
+We assume that $x_{1}=x_{3}$.
+According to the protocol (cf.
+Figure 8 in their paper), in step 3.(a), $P_{1}$ and $P_{2}$ invoke the OPRF where $P_{1}$ acts as $\mathcal{R}$ inputting $x_{1}$ and $P_{2}$ acts as $\mathcal{S}$ inputting its PRF key $k_{2}$.
+$P_{1}$ receives the PRF value $F_{k_{2}}(x_{1})$.
+Meanwhile, in step 3.(c), $P_{2}$ and $P_{3}$ invoke the cOPRF where $P_{3}$ acts as $\mathcal{R}$ inputting $x_{3}$, and $P_{2}$ acts as $\mathcal{S}$ inputting its PRF key $k_{2}$ and the set $X_{2}$.
+$P_{3}$ receives the output $w$ from the cOPRF.
+By the definition of cOPRF functionality, if $x_{3}\notin X_{2}$, $w=F_{k_{2}}(x_{3})$, otherwise $w$ is a random value.
+    If $P_{1}$ and $P_{3}$ collude, they can distinguish the cases where $x_{3}\in X_{2}$ and $x_{3}\notin X_{2}$ by comparing $P_{1}$’s output $F_{k_{2}}(x_{1})$ from the OPRF and $P_{3}$’s output $w$ from the cOPRF for equality.
+To elaborate, we recall that $x_{1}=x_{3}$, so $F_{k_{2}}(x_{1})=F_{k_{2}}(x_{3})$.
+If $F_{k_{2}}(x_{1})=w$, it implies that $P_{3}$ receives $F_{k_{2}}(x_{3})$ from the cOPRF, so the coalition learns that $x_{3}\notin X_{2}$; On the contrary, if $F_{k_{2}}(x_{1})\neq w$, it implies that $P_{3}$’s output from the cOPRF is not $F_{k_{2}}(x_{3})$, so it is a random value, then the coalition learns that $x_{3}\in X_{2}$.
+More generally, as long as $P_{1}$ and $P_{3}$ collude, they can identify whether each element $x\in X_{1}\cap X_{3}$ belongs to $X_{2}$ or not, by comparing the PRF value $F_{k}(x)$ from the OPRF between $P_{1}$ and $P_{2}$ and the cPRF value (whose condition depends on $x\in X_{2}$ or not) from the cOPRF between $P_{2}$ and $P_{3}$.
+This acquired knowledge is information leakage in MPSU.
+Therefore, the non-collusion assumption is required.
     Appendix 0.B Public-Key Encryption
@@
     $\{\mathsf{Enc}(pk,{m}_{1}^{\prime}),\cdots,\mathsf{Enc}(pk,{m}_{q}^{\prime}):(pk,sk)\leftarrow\mathsf{Gen}(1^{\lambda})\}$
-Appendix 0.C Construction of Multi-Party Secret-Shared Shuffle
-The functionality of share correlation generated in the offline phase is depicted in Figure 15. The online details of multi-party secret-shared shuffle in [EB22] are given in Figure 16.
-  • Parameters. $m$ parties $P_{1},\cdots P_{m}$. The dimension of vector $n$. The item length $l$.
-  • Functionality.
-On input $\pi_{i}:[n]\to[n]$ from each $P_{i}$ ($1\leq i\leq m$), sample $\vec{a}_{i}^{\prime},\vec{b}_{i}\leftarrow(\{0,1\}^{l})^{n}$ for $11\leq i\leq m-$ 1, and $\vec{a}_{i},\vec{\Delta}_{m}\leftarrow(\{0,1\}^{l})^{n}$ for $2\leq i\leq m$, which satisfy
-[…]
-    Give $\vec{a}_{1}^{\prime},\vec{b}_{1}$ to $P_{1}$, $\vec{a}_{i}^{\prime},\vec{a}_{i},\vec{b}_{i}$ to $P_{i}$ ($1<i<m$), and $\vec{a}_{m},\vec{\Delta}_{m}$ to $P_{m}$.
-Figure 15: Share Correlation Functionality $\mathcal{F}_{\mathsf{sc}}$
-  • Parameters: $m$ parties $P_{1},\cdots,P_{m}$. Ideal functionality $\mathcal{F}_{\mathsf{sc}}$ in Figure 15. The dimension of vector $n$. The item length $l$.
-  • Inputs: Each party $P_{i}$ has input $\vec{x}_{i}={(x_{i}^{1},\cdots,x_{i}^{n})}$.
-  • Protocol:
-    1. 1.
-      Each party $P_{i}$ samples a random permutation $\pi_{i}:[n]\to[n]$ and invokes $\mathcal{F}_{\mathsf{sc}}$ with input $\pi_{i}$.
-$P_{1}$ receives $\vec{a}_{1}^{\prime},\vec{b}_{1}\in(\{0,1\}^{l})^{n}$; For $2\leq j<m$, $P_{j}$ receives $\vec{a}_{j}^{\prime},\vec{a}_{j},\vec{b}_{j}\in(\{0,1\}^{l})^{n}$; $P_{m}$ receives $\vec{a}_{m},\vec{\Delta}_{m}\in(\{0,1\}^{l})^{n}$.
-    2. 2.
-      For $2\leq j\leq m$, $P_{j}$ computes $\vec{c}_{j}=\vec{x}_{j}\oplus\vec{a}_{j}$ and sends $\vec{c}_{j}$ to $P_{1}$.
-    3. 3.
-      $P_{1}$ computes $\vec{c^{\prime}}_{1}=\pi_{1}(\bigoplus_{j=2}^{m}\vec{c}_{j}\oplus\vec{x}_{1})\oplus\vec{a}_{1}^{\prime}$ and send it to $P_{2}$. $P_{1}$ outputs $\vec{b}_{1}$.
-    4. 4.
-      For $2\leq j<m$, $P_{j}$ computes $\vec{c^{\prime}}_{j}=\pi_{j}(\vec{c^{\prime}}_{j-1})\oplus\vec{a}_{j}^{\prime}$ and send it to $P_{j+1}$. $P_{j}$ outputs $\vec{b}_{j}$.
-    5. 5.
-      $P_{m}$ outputs $\pi_{m}(\vec{c^{\prime}}_{m-1})\oplus\vec{\Delta}_{m}$.
-Figure 16: Multi-Party Secret-Shared Shuffle Protocol $\Pi_{\mathsf{ms}}$
-Appendix 0.D Construction of MKR-PKE from ElGamal
-The MKR-PKE is instantiated using ElGamal encryption as follows:
+    Appendix 0.C Construction of MKR-PKE from ElGamal
+    MKR-PKE is instantiated with ElGamal encryption below:
     • The key-generation algorithm $\mathsf{Gen}$ takes as input the security parameter $1^{\lambda}$ and generates $(\mathbb{G},g,p)$, where $\mathbb{G}$ is a cyclic group, $g$ is the generator and $q$ is the order.
 Outputs $sk$ and $pk=g^{sk}$.
-  • The randomized encryption algorithm $\mathsf{Enc}$ takes as input a public key $pk$ and a plaintext message $x\in\mathbb{G}$, samples $r\leftarrow\mathbb{Z}_{q}$, and outputs a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})=(g^{r},x\cdot pk^{r})$.
+    • The randomized encryption algorithm $\mathsf{Enc}$ takes as input a public key $pk$ and a plaintext message $x\in\mathbb{G}$, samples $r\leftarrow\mathbb{Z}_{q}$, and outputs $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})=(g^{r},x\cdot pk^{r})$.
     • The partial decryption algorithm $\mathsf{ParDec}$ takes as input a secret key share $sk$ and a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})$, and outputs $\mathsf{ct}^{\prime}=(\mathsf{ct}_{1},\mathsf{ct}_{2}\cdot\mathsf{ct}_{1}^{-sk})$.
@@
     • The rerandomization algorithm $\mathsf{ReRand}$ takes as input $pk$ and a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})$, samples $r\leftarrow\mathbb{Z}_{q}$, and outputs $\mathsf{ct}^{\prime}=(\mathsf{ct}_{1}\cdot g^{r^{\prime}},\mathsf{ct}_{2}\cdot pk^{r^{\prime}})$.
-Appendix 0.E Missing Security Proofs
-0.E.1 The Proof of Theorem 5.1
-  • Parameters: $m$ parties $P_{1},\cdots,P_{m}$. The message length $l$.
-  • Inputs: $P_{1}$ has input $b_{1}\in\{0,1\}$ and $P_{m}$ has input $b_{m}\in\{0,1\}$.
-  • Protocol:
-    1. 1.
-      $P_{1}$ and $P_{m}$ invoke $\mathcal{F}_{\mathsf{rot}}$ twice: First, $P_{1}$ acts as receiver with input $b_{1}$ and $P_{m}$ acts as sender.
-$P_{1}$ receives $r_{m,1}^{b_{1}}\in\{0,1\}^{l}$.
-$P_{m}$ receives $r_{m,1}^{0},r_{m,1}^{1}\in\{0,1\}^{l}$; Second, $P_{m}$ acts as receiver with input $b_{m}$ and $P_{1}$ acts as sender.
-$P_{m}$ receives $r_{1,m}^{b_{m}}\in\{0,1\}^{l}$.
-$P_{1}$ receives $r_{1,m}^{0},r_{1,m}^{1}\in\{0,1\}^{l}$.
-    2. 2.
-      For $1<j<m$: $P_{1}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{1}$ acts as receiver with input $b_{1}$ and $P_{j}$ as sender without input.
-$P_{1}$ receives $r_{j,1}^{b_{1}}\in\{0,1\}^{l}$.
-$P_{j}$ receives $r_{j,1}^{0},r_{j,1}^{1}\in\{0,1\}^{l}$.
-    3. 3.
-      For $1<j<m$: $P_{m}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{m}$ acts as receiver with input $b_{m}$ and $P_{j}$ as sender.
-$P_{m}$ receives $r_{j,m}^{b_{m}}\in\{0,1\}^{l}$.
-$P_{j}$ receives $r_{j,m}^{0},r_{j,m}^{1}\in\{0,1\}^{l}$.
-    4. 4.
-      For $1<j<m$: $P_{j}$ samples $\Delta_{j}\leftarrow\{0,1\}^{l}$ and computes $r_{j}=r_{j,1}^{0}\oplus r_{j,m}^{0}$.
-$P_{j}$ sends $u_{j,1}=\Delta_{j}\oplus r_{j,1}^{0}\oplus r_{j,1}^{1}$ to $P_{1}$, and $u_{j,m}=\Delta_{j}\oplus r_{j,m}^{0}\oplus r_{j,m}^{1}$ to $P_{m}$, then outputs $(r_{j},\Delta_{j})$.
-    5. 5.
-      $P_{1}$ computes $\Delta_{1}=r_{1,m}^{0}\oplus r_{1,m}^{1}$ and $r_{1}=\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1}$ ($\cdot$ denotes bitwise-AND between the repetition code of $b_{1}$ and $u_{j,1}$, which are both strings of length $l$.
-Similarly hereinafter), then outputs $(r_{1},\Delta_{1})$.
-$P_{m}$ computes $\Delta_{m}=r_{m,1}^{0}\oplus r_{m,1}^{1}$ and $r_{m}=\bigoplus_{j=2}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus r_{1,m}^{b_{m}}\oplus b_{m}\cdot\Delta_{m}$, then outputs $(r_{m},\Delta_{m})$.
-Figure 17: Multi-Party Secret-Shared Random OT $\Pi_{\mathsf{mss\text{-}rot}}$
-In order to facilitate the proof, we assume that $\text{ch}_{0}=1$ and $\text{ch}_{1}=m$, and $P_{1}$ and $P_{m}$ provide inputs $b_{1}\in\{0,1\}$ and $b_{m}\in\{0,1\}$ respectively.
-We turn to proving the correctness and security of the protocol in Figure 17.
-Note that when $\text{ch}_{0}$ and $\text{ch}_{1}$ are assigned different values, the proof is essentially the same.
+    Appendix 0.D Missing Security Proofs
+    0.D.1 The Proof of Theorem 5.1
+    We consider the case that $\text{ch}_{0}=1$ and $\text{ch}_{1}=m$, and $J=\{1,\cdots,m\}$.
+Let $P_{1}$ and $P_{m}$ input $b_{1}\in\{0,1\}$ and $b_{m}\in\{0,1\}$, and $P_{i}$ ($1\leq i\leq m$) input $\Delta_{i}$ respectively.
+We turn to proving the correctness and security of the protocol in Figure 10 in this particular case.
+Note that in the different cases, the proof is essentially the same.
     • Correctness. From the description of the protocol, we have the following equations:
-    $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}=(\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1})$
-    $\oplus(r_{j1}^{0}\oplus r_{j,m}^{0})\oplus(\bigoplus_{j=2}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus r_{m,1}^{b_{m}}\oplus b_{m}\cdot\Delta_{m})$ (1)
+      $r_{1}=\bigoplus_{j=2}^{m}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus b_{1}\cdot\Delta_{1},$ (1)
+      $r_{i}=r_{i,1}^{0}\oplus r_{i,m}^{0},1<i<m,$ (2)
+      $r_{m}=\bigoplus_{j=1}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus b_{m}\cdot\Delta_{m},$ (3)
       $u_{j,1}=\Delta_{j}\oplus r_{j,1}^{0}\oplus r_{j,1}^{1},u_{j,m}=\Delta_{j}\oplus r_{j,m}^{0}\oplus r_{j,m}^{1}$ (4)
       From the definition of Random OT functionality (Figure 5), we have the following equations:
-    $r_{j,1}^{b_{1}}=r_{j,1}^{0}\oplus b_{1}\cdot(r_{j,1}^{0}\oplus r_{j,1}^{1}),r_{j,m}^{b_{m}}=r_{j,m}^{0}\oplus b_{m}\cdot(r_{j,m}^{0}\oplus r_{j,m}^{1}),$ (3)
-    $r_{m,1}^{b_{1}}=r_{m,1}^{0}\oplus b_{1}\cdot(r_{j,1}^{0}\oplus r_{j,1}^{1})=r_{m,1}^{0}\oplus b_{1}\cdot\Delta_{1},$ (4)
-    $r_{1,m}^{b_{m}}=r_{1,m}^{0}\oplus b_{m}\cdot(r_{1,m}^{0}\oplus r_{1,m}^{1})=r_{1,m}^{0}\oplus b_{m}\cdot\Delta_{m}$ (5)
+      $r_{j,1}^{b_{1}}=r_{j,1}^{0}\oplus b_{1}\cdot(r_{j,1}^{0}\oplus r_{j,1}^{1}),$ (5)
+      $r_{j,m}^{b_{m}}=r_{j,m}^{0}\oplus b_{m}\cdot(r_{j,m}^{0}\oplus r_{j,m}^{1}),$ (6)
       Substitute Equation 456 into Equation 122 and cancel out the same terms, we obtain:
-    $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}$
-    $=(b_{1}\oplus b_{m})\cdot\Delta_{1}\oplus(\bigoplus_{j=2}^{m-1}(b_{1}\oplus b_{m})\cdot\Delta_{j})\oplus(b_{1}\oplus b_{m})\cdot\Delta_{m}$
-    $=\bigoplus_{i=1}^{m}(b_{1}\oplus b_{m})\cdot\Delta_{i}$
+      $r_{1}=\bigoplus_{j=2}^{m}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{j})\oplus r_{1,m}^{0}\oplus b_{1}\cdot\Delta_{1},$ (7)
+      $r_{m}=\bigoplus_{j=1}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot\Delta_{j})\oplus r_{m,1}^{0}\oplus b_{m}\cdot\Delta_{m},$ (8)
+      Substitute Equation 278 into $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}$, we have $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}=\bigoplus_{i=1}^{m}(b_{1}\oplus b_{m})\cdot\Delta_{i}$.
 Then we can summarize that if $b_{1}\oplus b_{m}=0$, $r_{1}=\bigoplus_{j=2}^{m}r_{j}$, else $r_{1}=\Delta_{1}\oplus(\bigoplus_{j=2}^{m}(r_{j}\oplus\Delta_{j}))$.
@@
 The simulator can easily simulate these outputs from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages by generating random values, which are independent of honest parties’ inputs.
-    To elaborate, in the case that $P_{1}\notin\mathcal{C}orr$ and $P_{m}\notin\mathcal{C}orr$, simulator receives all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$ and needs to emulate each $P_{c}$’s view, including its private $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$.
-The simulator for corrupted $P_{c}$ runs the protocol honestly except that it sets $P_{c}$’s random tape to be the output $\Delta_{c}$, and simulates uniform outputs from $\mathcal{F}_{\mathsf{rot}}$ under the constraint that $r_{c,1}^{0}\oplus r_{c,m}^{0}=r_{c}$.
-Clearly, the joint distribution of all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
+      To elaborate, in the case that $P_{1}\notin\mathcal{C}orr$ and $P_{m}\notin\mathcal{C}orr$, simulator receives all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$ and needs to emulate each $P_{c}$’s view, including its private input $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$.
+The simulator for corrupted $P_{c}$ runs the protocol honestly except that it simulates uniform outputs from $\mathcal{F}_{\mathsf{rot}}$ under the constraint that $r_{c,1}^{0}\oplus r_{c,m}^{0}=r_{c}$.
+Clearly, the joint distribution of all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
       In the case that $P_{1}\in\mathcal{C}orr$ or $P_{m}\in\mathcal{C}orr$, since the protocol is symmetric with respect to the roles of $P_{1}$ and $P_{m}$, we focus on the case of corrupted $P_{1}$.
-The simulator receives $b_{1}$ in addition to all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$.
-For $P_{1}$, its view consists of its private $\Delta_{1}$, outputs $r_{m,1}^{b_{1}}$, $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j<m}$ from $P_{j}$.
-For each $P_{c}(c\neq 1)$, its view consists of its private $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$.
-    The simulator sets each corrupted party $P_{c}$’s random tape to be its output $\Delta_{c}$.
-Then for $P_{1}$’s view, it runs the protocol honestly except that it simulates uniform outputs $r_{m,1}^{b_{1}}$, $(r_{1,m}^{0},r_{1,m}^{1})$, $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$.
+The simulator receives all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$.
+For $P_{1}$, its view consists of the choice bit $b_{1}$, its private input $\Delta_{1}$, outputs $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j\leq m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j\leq m}$ from $P_{j}$.
+For each $P_{c}(c\neq 1)$, its view consists of its private input $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$.
+      For $P_{1}$’s view, the simulator runs the protocol honestly except that it simulates uniform outputs $(r_{1,m}^{0},r_{1,m}^{1})$, $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$.
 For other corrupted parties’ view, it runs the protocol honestly except that it sets $r_{c,m}^{0}=r_{c,1}^{0}\oplus r_{c}$ and simulates uniform output $r_{c,m}^{1}$ from $\mathcal{F}_{\mathsf{rot}}$.
@@
 Therefore, $u_{i,1}$ is uniformly at random from the perspective of $P_{1}$.
-Clearly, the joint distribution of $b_{1}$ and all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
-    In the case that $P_{1}\in\mathcal{C}orr$ and $P_{m}\in\mathcal{C}orr$, the simulator receives $b_{1},b_{m}$ and all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$.
-For $P_{1}$, its view consists of its private $\Delta_{1}$, outputs $r_{m,1}^{b_{1}}$, $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j<m}$ from $P_{j}$.
-For each $P_{c}(c\neq 1,c\neq m)$, its view consists of its private $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$.
-For $P_{m}$, its view consists of its private $\Delta_{m}$, outputs $r_{1,m}^{b_{m}}$, $(r_{m,1}^{0},r_{m,1}^{1})$, $\{r_{j,1}^{b_{m}}\}_{1<j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,m}\}_{1<j<m}$ from $P_{j}$.
-    The simulator sets each corrupted party $P_{c}$’s random tape to be its output $\Delta_{c}$.
-Then for $P_{1}$’s view, it runs the protocol honestly except that it simulates uniform outputs $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$.
+Clearly, the joint distribution of all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
+      In the case that $P_{1}\in\mathcal{C}orr$ and $P_{m}\in\mathcal{C}orr$, the simulator receives all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$.
+For $P_{1}$, its view consists of the choice bit $b_{1}$, its private input $\Delta_{1}$, outputs $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j\leq m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j\leq m}$ from $P_{j}$.
+For each $P_{c}(c\neq 1,c\neq m)$, its view consists of its private input $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$.
+For $P_{m}$, its view consists of its private input $\Delta_{m}$, outputs $(r_{m,1}^{0},r_{m,1}^{1})$, $\{r_{j,m}^{b_{m}}\}_{1\leq j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,m}\}_{1\leq j<m}$ from $P_{j}$.
+      For $P_{1}$’s view, the simulator runs the protocol honestly except that it simulates uniform outputs $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$.
 For the view of $P_{c}(c\neq 1,c\neq m)$, it runs the protocol honestly except that it sets $r_{c,m}^{0}=r_{c,1}^{0}\oplus r_{c}$ and simulates uniform output $r_{c,m}^{1}$ from $\mathcal{F}_{\mathsf{rot}}$.
 For $P_{m}$’s view, it runs the protocol honestly with the following changes:
-    • It simulates uniform outputs $r_{i,m}^{b_{m}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniform messages $u_{i,m}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus r_{1,m}^{b_{m}}\oplus b_{m}\cdot\Delta_{m}=r_{m}$, where $P_{i}\in\mathcal{H}$.
+      • It simulates uniform outputs $r_{i,m}^{b_{m}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniform messages $u_{i,m}$ from $P_{i}$ under the constraint $\bigoplus_{j=1}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus b_{m}\cdot\Delta_{m}=r_{m}$, where $P_{i}\in\mathcal{H}$.
       • It sets the output $r_{c,m}^{b_{m}}$ from $\mathcal{F}_{\mathsf{rot}}$ to be consistent with the partial view $(r_{c,m}^{0},r_{c,m}^{1})$ of each corrupted $P_{c}$ in preceding simulation, where $c\neq 1$ and $c\neq m$.
-    Clearly, the joint distribution of $b_{1},b_{m}$ and all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
-0.E.2 The Proof of Theorem 5.2
+      Clearly, the joint distribution of all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
+    0.D.2 The Proof of Theorem 5.2
     Proof
@@
 Specifically, the simulator receives the input $X_{c}$ of $P_{c}\in\mathcal{C}orr$ and the output $\bigcup_{i=1}^{m}X_{i}$ if $P_{1}\in\mathcal{C}orr$.
-For each $P_{c}$, its view consists of its input $X_{c}$, outputs from $\mathcal{F}_{\mathsf{bssPMT}}$, $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, output $\vec{sh}_{c}^{\prime}$ from $\mathcal{F}_{\mathsf{ms}}$ as its share, and $m-1$ sets of shares $\{\vec{sh}_{i}^{\prime}\}_{1<i\leq m}$($P_{i}$’s output from $\mathcal{F}_{\mathsf{ms}}$) from $P_{i}$ if $c=1$.
+    For each $P_{c}$, its view consists of its input $X_{c}$, outputs from $\mathcal{F}_{\mathsf{bssPMT}}$, $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, output $\vec{sh}_{c}^{\prime}$ from $\mathcal{F}_{\mathsf{ms}}$ as its share, sampled values as shares of $\Delta$ for $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ and $m-1$ sets of shares $\{\vec{sh}_{i}^{\prime}\}_{1<i\leq m}$($P_{i}$’s output from $\mathcal{F}_{\mathsf{ms}}$) from $P_{i}$ if $c=1$.
 The simulator emulates each $P_{c}$’s view by running the protocol honestly with the following changes:
     • In step 2, it simulates uniform outputs $\{e_{c,j}^{b}\}_{c<j\leq m}$ and $\{e_{c,i}^{b}\}_{1\leq i<c}$ from $\mathcal{F}_{\mathsf{bssPMT}}$, on condition that $P_{i},P_{j}\in\mathcal{H}$.
-  • In step 3, it simulates uniform outputs $\{r_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j},\{\Delta_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, on condition that $\exists 1\leq d\leq j,P_{d}\in\mathcal{H}$.
+    • In step 3, it simulates uniform outputs $\{r_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, on condition that $\exists min(2,i)\leq d\leq j,P_{d}\in\mathcal{H}$.
+If $c\neq 1$, it simulates uniform $\{\Delta_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ as $P_{c}$’s random tapes.
     • In step 4, it simulates uniformly output $\vec{sh}_{c}^{\prime}$ from $\mathcal{F}_{\mathsf{ms}}$.
@@
 By the definition of $\mathcal{F}_{\mathsf{bssPMT}}$, each output $e_{c,j}^{b}$ and $e_{c,i}^{b}$ from $\mathcal{F}_{\mathsf{bssPMT}}$ is uniformly distributed as a secret-share between $P_{c}$ and $P_{j}$, or $P_{i}$ and $P_{c}$, where $P_{i},P_{j}\in\mathcal{H}$.
-By the definition of $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, each output $r_{c,ji}^{b}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ is a secret-share of 00 among $P_{1},\cdots,P_{j}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=0$, or a secret-share of $\bigoplus_{d=1}^{j}\Delta_{d,ji}^{b}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=1$.
-Therefore, even if $P_{c}$ colludes with others, $r_{c,ji}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{d}\in\mathcal{H}(1\leq d\leq j)$ holding one share.
-Besides, the outputs $\{\Delta_{c,ji}^{b}\}_{c<j\leq m,1\leq i<j}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ are uniformly distributed.
+By the definition of $\mathcal{F}_{\mathsf{mss\text{-}rot}}$, each output $r_{c,ji}^{b}$ from $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ is a secret-share of 00 among $P_{min(2,i)},\cdots,P_{j}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=0$, or a secret-share of $\bigoplus_{d=2}^{j}\Delta_{d,ji}^{b}$ if $e_{i,j}^{b}\oplus e_{j,i}^{b}=1$.
+Therefore, even if $P_{c}$ colludes with others, $r_{c,ji}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{d}\in\mathcal{H}(min(2,i)\leq d\leq j)$ holding one share.
     It remains to demonstrate that the output $\vec{sh}_{c}^{\prime}$ from $\mathcal{F}_{\mathsf{ms}}$($P_{1}\notin\mathcal{C}orr$) or all outputs $\{\vec{sh}_{i}^{\prime}\}_{1\leq i\leq m}$ from $\mathcal{F}_{\mathsf{ms}}$($P_{1}\in\mathcal{C}orr$) does not leak any other information except for the union.
@@
 $x\in X_{i}$, then for all $1\leq i<j$, $e_{i,j}^{y}\oplus e_{j,i}^{y}=0$.
-By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 10, each $r_{d,ji}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=1}^{j}r_{d,ji}^{b}=0$.
-From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=1}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)$, namely, they are additive shares of $x\|\mathsf{H}(x)$ among all parties.
+By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 9, each $r_{d,ji}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=min(2,i)}^{j}r_{d,ji}^{b}=0$.
+From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform in $\{0,1\}^{l+\kappa}$ conditioned on $\bigoplus_{d=min(2,i)}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)$, namely, they are additive shares of $x\|\mathsf{H}(x)$ among all parties.
 This is exactly identical to the simulation.
@@
 $x\in X_{i}$, then $e_{i,j}^{b}\oplus e_{j,i}^{b}=1$.
-By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 10, each $r_{d,ji}^{b}$ is uniform conditioned on $\bigoplus_{d=1}^{j}r_{d,ji}^{b}=\bigoplus_{d=1}^{j}\Delta_{d,ji}^{b}$, where each $\Delta_{d,ji}^{b}$ is uniformly held by $P_{d}$.
-From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform conditioned on $\bigoplus_{d=1}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)\oplus\bigoplus_{d=1}^{j}\Delta_{d,ji}^{b}\oplus r$, where $r$ is the sum of remaining terms.
-Then, even if $P_{1}$ colludes with others, $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{d}\in\mathcal{H}(1\leq d\leq j)$ holding one uniform $\Delta_{d,ji}^{b}$ and independent of all honest parties’ inputs.
+By the $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ functionality in Figure 9, each $r_{d,ji}^{b}$ is uniform conditioned on $\bigoplus_{d=min(2,i)}^{j}r_{d,ji}^{b}=\bigoplus_{j^{\prime}=2}^{j}\Delta_{j^{\prime},ji}^{b}$, where each $\Delta_{j^{\prime},ji}^{b}$ is uniformly held by $P_{j^{\prime}}$ ($1<j^{\prime}\leq j$).
+From the descriptions of the protocol, We derive that each $u_{d,j}^{b}$ is uniform conditioned on $\bigoplus_{d=min(2,i)}^{j}u_{d,j}^{b}=x\|\mathsf{H}(x)\oplus\bigoplus_{j^{\prime}=2}^{j}\Delta_{j^{\prime},ji}^{b}\oplus r$, where $r$ is the sum of remaining terms.
+Then, even if $P_{1}$ colludes with others, $\bigoplus_{d=min(2,i)}^{j}u_{d,j}^{b}$ is still uniformly random from the perspective of adversary, since there always exists a party $P_{j^{\prime}}\in\mathcal{H}(1<j^{\prime}\leq j)$ holding one uniform $\Delta_{j^{\prime},ji}^{b}$ and independent of all honest parties’ inputs.
 For all empty bins, $u_{d,j}^{b}$ is chosen uniformly random, so the corresponding $\bigoplus_{d=1}^{j}u_{d,j}^{b}$ is also uniformly at random, which is identical to the simulation.
@@
 We conclude that all outputs $\{\vec{sh}_{i}^{\prime}\}_{1\leq i\leq m}$ from $\mathcal{F}_{\mathsf{ms}}$ distribute identically between the real and ideal executions.
-0.E.3 The Proof of Theorem 6.1
+    0.D.3 The Proof of Theorem 6.1
     Proof
@@
 To avoid repetition, we omit the analysis here.
-Appendix 0.F Complete Complexity Analyses of Protocols
-0.F.1 Theoretical Analysis of Protocol $\Pi_{\mathsf{bssPMT}}$
+    Appendix 0.E Complexity Analyses and Comparisons
+    0.E.1 Complete Analysis of LG
+    The costs of each stage in LG are calculated as follows.
+    • mq-ssPMT. The cost of this stage consists of three parts:
+      • SKE encryption: The computation complexity of encrypting $10,1,\cdots,n-$ 1 is $O(n)$.
+      • OKVS: The computation complexities of Encode and Decode algorithms are both $O(n)$.
+The size of OKVS is $1.28\kappa n$ bits, where $\kappa$ is the size of a SKE ciphertext with the same range as before.
+      • ssVODM: The ssVODM protocol requires a total of $(T+l-\log n-1)n$ AND gates, where $T$ is the number of AND gates in the SKE decryption circuit.
+      For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke mq-ssPMT of size $B$. Overall, each party $P_{j}$ engages in $(m-1)$ instances of mq-ssPMT.
+      • Offline phase.
+The computation complexity per party is $O((T+l-\log n)mn\log((T+l-\log n)n))$.
+The communication complexity per party is $O(t\lambda m\log(((T+l-\log n)n)/t))$.
+The round complexity is $O(1)$.
+      • Online phase. The computation complexity per party is $O((T+l-\log n)mn)$. The communication complexity per party is $O((T+l+\kappa-\log n)mn)$. The round complexity is $O(\log(l-\log n))$.
+    • Random oblivious transfers. For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke ROT extension of size $B$. Overall, each party $P_{j}$ engages in $(m-1)$ instances of ROT extension of size $B$.
+      • Offline phase. The computation complexity per party is $O(mn\log n)$. The communication complexity per party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
+      • Online phase. The computation complexity per party is $O(mn)$. The communication complexity per party is $O(mn)$. The round complexity is $O(1)$.
+      The costs of the remaining two steps, multi-party secret-shared shuffle and output reconstruction, are exactly the same as those in our SK-MPSU.(cf. Appendix  0.E.3).
+    • Total costs.
+      • Offline phase.
+The computation complexity per party is $O((T+l-\log n)mn\log((T+l-\log n)n)+m^{2}n\log mn)$.
+The communication complexity per party is $O(t\lambda m\log(((T+l-\log n)n)/t)+\lambda m^{2}n(\log m+\log n))$.
+The round complexity is $O(1)$.
+      • Online phase.
+The computation complexity of $P_{1}$ is $O((T+l-\log n)mn+m^{2}n)$.
+The communication complexity of $P_{1}$ is $O((T+l+\kappa-\log n)mn+(l+\kappa)m^{2}n)$.
+The computation complexity of $P_{j}$ is $O((T+l-\log n)mn)$.
+The communication complexity of $P_{j}$ is $O((T+l+\kappa-\log n)mn)$.
+The round complexity is $O(\log(l-\log n)+m)$.
+    0.E.2 Theoretical Analysis of Batch ssPMT
+    0.E.2.1 Complete Analysis
     The costs of each stage in $\Pi_{\mathsf{bssPMT}}$ (Figure 8) are calculated as follows.
@@
         For $B=O(n)$ instances of OPRF, we require performing a subfield-VOLE of size $B$ in the offline phase, and sending $B$ derandomization messages of length $l$ in the online phase.
-We resort to Dual LPN with Fixed Weight and Regular Noise [BCG+19a, BCG+19b] to improve the efficiency of subfield-VOLE, and instantiate the code family with Expand-Convolute codes presented by [RRT23], which enables near-linear time syndrome computation.
+We resort to Dual LPN with Fixed Weight and Regular Noise [BCG+19a, BCG+19b] to improve efficiency of subfield-VOLE, and instantiate the code family with Expand-Convolute codes [RRT23], which enables near-linear time syndrome computation.
 The computation complexity is $O(n\log n)$.
@@
 The round complexity is $O(1)$.
+        We denote the output length of OPPRF as $\gamma$.
+The lower bound of $\gamma$ is relevant to the total number of the batch OPPRF invocations.
+In our SK-MPSU and PK-MPSU, for $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke batch OPPRF.
+Overall, there are $21+2+\cdots+(m-1)=(m^{2}-m)/$ 2 invocations of batch OPPRF.
+Considering all these invocations, we set $\gamma\geq\sigma+\log((m^{2}-m)/2)+\log_{2}B$, so that the probability of any $t_{i}\neq s_{i}$ occurring if $x\notin X_{i}$, which is $((m^{2}-m)/2))B\cdot 2^{-\gamma}$, is less than or equal to $2^{-\sigma}$.
       • OKVS: The size of key-value pairs encoded into OKVS is $\lvert K_{1}\rvert+\cdots+\lvert K_{B}\rvert$.
@@
 We use the OKVS construction of [RR22], and the computation complexities of $\mathsf{Encode_{H}}$ and $\mathsf{Decode_{H}}$ algorithms are both $O(n)$.
-As we employ their $w=3$ scheme with a cluster size of $2^{14}$, the size of OKVS is $1.28\gamma\cdot(3n)$ bits, where $\gamma$ is the output length of OPPRF and is set as $\sigma+\log_{2}B$ to avoid wrong matching.
-  • Offline phase. The computation complexity of each party is $O(n\log n)$. The communication complexity of each party is $O(t\lambda\log(n/t))$. The round complexity is $O(1)$.
-  • Online phase. The computation complexity of each party is $O(n)$. The communication complexity of each party is $O(\gamma n)$. The round complexity is $O(1)$.
+As we employ their $w=3$ scheme with a cluster size of $2^{14}$, the size of OKVS is $1.28\gamma\cdot(3n)$ bits.
+    • Offline phase. The computation complexity per party is $O(n\log n)$. The communication complexity per party is $O(t\lambda\log(n/t))$. The round complexity is $O(1)$.
+    • Online phase. The computation complexity per party is $O(n)$. The communication complexity per party is $O(\gamma n)$. The round complexity is $O(1)$.
   • Secret-shared private equality tests.
@@
 Therefore, we count the online computation complexity of ssPEQT as $O(1)$.
-    • Offline phase. The computation complexity of each party is $O(\gamma n\log n)$. The communication complexity of each party is $O(t\lambda\log(\gamma n/t))$. The round complexity is $O(1)$.
-    • Online phase. The computation complexity of each party is $O(n)$. The communication complexity of each party is $O(\gamma n)$. The round complexity is $O(\log\gamma)$.
-  • 0.F.2 Theoretical Analysis of Protocol $\Pi_{\mathsf{SK\text{-}MPSU}}$
+    • Offline phase. The computation complexity per party is $O(\gamma n\log n)$. The communication complexity per party is $O(t\lambda\log(\gamma n/t))$. The round complexity is $O(1)$.
+    • Online phase. The computation complexity per party is $O(n)$. The communication complexity per party is $O(\gamma n)$. The round complexity is $O(\log\gamma)$.
+  • 0.E.2.2 Comparison with mq-ssPMT
+    Liu and Gao adapt the SKE-based mq-RPMT in [ZCL+23] to construct mq-ssPMT.
+The most expensive part of their construction is secret-shared oblivious decryption-then-matching (ssVODM) [ZCL+23], which is to implement a decryption circuit and a comparison circuit by the GMW protocol [GMW87].
+In total, the ssVODM requires $(T+l-\log n-1)n$ AND gates, where $l$ is the bit length of set elements, and $T$ is the number of AND gates in the SKE decryption circuit and is set to be considerably large ($\approx 600$) according to their paper.
+    The costs of batch ssPMT of size $B=1.27n$ [^22] We use stash-less Cuckoo hashing [PSTY19] with 3 hash functions, where $B=1.27n$.
+consist of the costs of batch OPPRF of size $1.27n$ and the costs of $1.27n$ instances of ssPEQT, where batch OPPRF can be instantiated by extremely fast specialized protocol and ssPEQT is implemented by the GMW protocol.
+Moreover, the state-of-the-art batch OPPRF construction [CGS22, RS21, RR22] can achieve linear computation and communication with respect to $n$ In ssPEQT, the parties engage in $1.27(\gamma-1)n$ AND gates, where $\gamma$ is the output length of OPPRF.
+In the typical setting where $n\leq 2^{24},l\leq 128,\gamma\leq 64$, we have $(T+l-\log n-1)n\gg 1.27(\gamma-1)n$, which means that the number of AND gates desired involved in batch ssPMT is far smaller than mq-ssPMT.
+Therefore, the construction built on batch ssPMT greatly reduces the dependency on general 2PC and significantly decreases both computation and communication complexity by a considerable factor, compared to mq-ssPMT.
+    0.E.3 Theoretical Analysis of Our SK-MPSU
+    0.E.3.1 Complete Analysis
     The costs of each stage in $\Pi_{\mathsf{SK\text{-}MPSU}}$ (Figure 11) are calculated as follows.
@@
 Overall, each party $P_{j}$ engages in $m-1$ invocations of batch ssPMT, acting as $\mathcal{R}$ in the first $j-1$ invocations, and acting as $\mathcal{S}$ in the last $m-j$ invocations.
-    • Offline phase. The computation complexity of each party is $O(\gamma mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
-    • Online phase. The computation complexity of each party is $O(mn)$. The communication complexity of each party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
+    • Offline phase. The computation complexity per party is $O(\gamma mn\log n)$. The communication complexity per party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
+    • Online phase. The computation complexity per party is $O(mn)$. The communication complexity per party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
   • Multi-party secret-shared random oblivious transfers.
-Each invocation of mss-ROT involves pairwise executions of two-party OT.
-If $P_{i}$ and $P_{j}$ hold the two choice bits, then $P_{i}$ and $P_{j}$ invoke two instances of OT and they separately invoke one instance of OT with the remaining parties.
-    Each OT execution consists of two parts: In the offline phase, the parties engages in a random-choice-bit ROT, then $\mathcal{S}$ sends one messages of length $l+\kappa$ ($\kappa=\sigma+\log(m-1)+\log n$) to $\mathcal{R}$ [^22] The OT execution between two choice-bit-holder does not include the sending step.
-This does not affect the total complexity, so we ignore this.; In the online phase, $\mathcal{R}$ sends a 1-bit derandomization message to transform ROT into a chosen-input version.
-    For $1\leq i<j\leq m$, $P_{1},\cdots,P_{j}$ engage in $B$ instances of mss-ROT, where $P_{i}$ and $P_{j}$ hold the two choice bits.
-Considering one instance of mss-ROT and fixing $j$, there are overall $j-1$ invocations of mss-ROT.
-For $1\leq d<j$, each $P_{d}$ has one chance to act as the choice-bit-holder (who invokes two instances of OT with another and one with $j-2$ remaining parties separately in one invocation of mss-ROT) and acts as a nomal party (who invokes one instances of OT with the two choice-bit-holders separately in one invocation of mss-ROT) in the rest $j-2$ invocations.
-$P_{j}$ act as the choice-bit-holder in $j-1$ invocations (who invokes $2+(j-2)=j$ instances of OT).
-    Overall, each party $P_{j^{\prime}}$ invokes $[{j^{\prime}}({j^{\prime}}-1)+\sum_{j={j^{\prime}}+1}^{m}(j+2(j-2))]B=(\frac{3m^{2}-{j^{\prime}}^{2}-5m+3{j^{\prime}}}{2})B$ instances of two-party OT.
-    • Offline phase. The computation complexity of each party is $O(m^{2}n\log n)$. The communication complexity of each party is $O(t\lambda m^{2}\log(n/t))$. The round complexity is $O(1)$.
-    • Online phase. The computation complexity of each party is $O(m^{2}n)$. The communication complexity of each party is $O(m^{2}n)$. The round complexity is $O(1)$.
+Each invocation of mss-ROT involves pairwise executions of two-party OT, where each OT execution consists of two parts: In the offline phase, the parties invoke random-choice-bit ROT, then $\mathcal{S}$ sends one messages of length $l+\kappa$ ($\kappa=\sigma+\log(m-1)+\log n$) to $\mathcal{R}$; In the online phase, $\mathcal{R}$ sends a 1-bit derandomization message to transform ROT into a chosen-input version.
+If $P_{i}$ and $P_{j}$ hold choice bits, then $P_{i}$ and $P_{j}$ invoke OT with the parties in $\{P_{2},\cdots,P_{j}\}$ except itself.
+    For $1\leq i<j\leq m$, $P_{min(2,i)},\cdots,P_{j}$ engage in $B$ instances of mss-ROT, where $P_{i}$ and $P_{j}$ hold the two choice bits.
+Considering the overall invocations of mss-ROT: $P_{1}$ invokes $[\sum_{j=2}^{m}(j-1)]B=\frac{m^{2}B-mB}{2}$ instances of two-party OT; For $1<i\leq m$, each party $P_{i}$ invokes $[(i-1)+\sum_{j=i+1}^{m}3(j-2)]B=\frac{3m^{2}B-3i^{2}B-9mB+11iB-2B}{2}$ instances of two-party OT.
+    • Offline phase. The computation complexity per party is $O(m^{2}n\log n)$. The communication complexity per party is $O(t\lambda m^{2}\log(n/t))$. The round complexity is $O(1)$.
+    • Online phase. The computation complexity per party is $O(m^{2}n)$. The communication complexity per party is $O(m^{2}n)$. The round complexity is $O(1)$.
   • Multi-party secret-shared shuffle. We use the construction in [EB22]. In the offline phase, each pair of parties runs a Share Translation protocol [CGP20] of size $(m-1)B$ and length $l+\kappa$.
-    • Offline phase. The computation complexity of each party is $O(m^{2}n\log(mn))$. The communication complexity of each party is $O(\lambda m^{2}n\log(mn))$. The round complexity is $O(1)$.
+    • Offline phase. The computation complexity per party is $O(m^{2}n\log(mn))$. The communication complexity per party is $O(\lambda m^{2}n\log(mn))$. The round complexity is $O(1)$.
     • Online phase.
@@
 The online round complexity is $O(\log\gamma+m)$.
-  • 0.F.3 Theoretical Analysis of Protocol $\Pi_{\mathsf{PK\text{-}MPSU}}$
+  • 0.E.3.2 Comparison with LG
+    We present a comparison of the theoretical computation and communication complexity for each party in both offline and online phases between LG and our SK-MPSU.
+    Computation Communication
+    Offline Online Offline Online
+    Leader Client Leader Client
+     [LG23] $(T+l+m)mn\log n$ $(T+l+m)mn$ $(T+l)mn$ $\lambda m^{2}n\log n$ $(T+l)mn+lm^{2}n$ $(T+l)mn$
+    Ours $(\gamma+m)mn\log n$ $m^{2}n$ $\lambda m^{2}n\log n$ $\gamma mn+lm^{2}n$ $(\gamma+l+m)mn$
+    Table 3: Asymptotic communication (bits) and computation costs of LG and our SK-MPSU in the offline and online phases.
+$n$ is the set size.
+$m$ is the number of participants.
+$\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$.
+$T$ is the number of AND gate in an SKE decryption circuit, $T\approx 600$.
+$l$ is the bit length of input elements.
+$\gamma$ is the output length of OPPRF.
+$t$ is the noise weight in dual LPN, $t\approx 128$.
+    Computation Communication
+    Offline Online Offline Online
+     [GNT23] $\gamma mn\log n(\log n/\log\log n)$ $mn(\log n/\log\log n)$ $t\lambda m\log n(\log n/\log\log n)$ $(\gamma+\lambda)mn(\log n/\log\log n)$
+    Ours $\gamma mn\log n$ $mn$ $t\lambda m\log n$ $(\gamma+\lambda)mn$
+    Table 4: Asymptotic communication (bits) and computation costs of [GNT23] and our PK-MPSU in the offline and online phases.
+In the offline phase, the computation is composed of symmetric-key operations; In the online phase, the computation is composed of public-key operations since we ignore symmetric-key operations.
+$n$ is the set size.
+$m$ is the number of participants.
+$\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$.
+$l$ is the bit length of input elements.
+$\gamma$ is the output length of OPPRF.
+$t$ is the noise weight in dual LPN, $t\approx 128$.
+    0.E.4 Theoretical Analysis of Our PK-MPSU
+    0.E.4.1 Complete Analysis
     The costs of each stage in $\Pi_{\mathsf{PK\text{-}MPSU}}$ (Figure 12) are calculated as follows.
-  • Batch ssPMT The cost of this stage is the same as that in $\Pi_{\mathsf{SK\text{-}MPSU}}$ (cf. Appendix 0.F.3).
-    • Offline phase. The computation complexity of each party is $O(\gamma mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
-    • Online phase. The computation complexity of each party is $O(mn)$. The communication complexity of each party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
+  • Batch ssPMT The cost of this stage is the same as that in $\Pi_{\mathsf{SK\text{-}MPSU}}$ (cf. Appendix 0.E.4).
+    • Offline phase. The computation complexity per party is $O(\gamma mn\log n)$. The communication complexity per party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
+    • Online phase. The computation complexity per party is $O(mn)$. The communication complexity per party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
   • Random oblivious transfers and messages rerandomization.
@@
 If $i\neq 1$, $P_{i}$ sends one ciphertext to $P_{j}$ and $P_{j}$ executes one rerandomization as well.
-    • Offline phase. The computation complexity of each party is $O(mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
-    • Online phase. The computation complexity of each party is $O(mn)$ public-key operations. The communication complexity of each party is $O(\lambda mn)$. The round complexity is $O(1)$.
+    • Offline phase. The computation complexity per party is $O(mn\log n)$. The communication complexity per party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
+    • Online phase. The computation complexity per party is $O(mn)$ public-key operations. The communication complexity per party is $O(\lambda mn)$. The round complexity is $O(1)$.
   • Messages decryptions and shufflings. Each party shuffles $(m-1)B$ ciphertexts and executes $(m-1)B$ partial decryptions before sending them to the next party.
-    • Online phase. The computation complexity of each party is $O(mn)$ public-key operations. The communication complexity of each party is $O(\lambda mn)$. The round complexity is $O(m)$.
+    • Online phase. The computation complexity per party is $O(mn)$ public-key operations. The communication complexity per party is $O(\lambda mn)$. The round complexity is $O(m)$.
   • Total costs.
-    • Offline phase. The computation complexity of each party is $O(\gamma mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(\gamma n/t)$. The round complexity is $O(1)$.
+    • Offline phase. The computation complexity per party is $O(\gamma mn\log n)$. The communication complexity per party is $O(t\lambda m\log(\gamma n/t)$. The round complexity is $O(1)$.
     • Online phase.
-The computation complexity of each party is $O(mn)$ symmetric-key operations and $O(mn)$ public-key operations.
-The communication complexity of each party is $O((\gamma+\lambda)mn)$.
+The computation complexity per party is $O(mn)$ symmetric-key operations and $O(mn)$ public-key operations.
+The communication complexity per party is $O((\gamma+\lambda)mn)$.
 The round complexity is $O(\log\gamma+m)$.
-  • 0.F.4 The Complete Analysis of LG Protocol
-    The costs of each stage in LG protocol are calculated as follows.
-    • mq-ssPMT. The cost of this stage mainly consists of three parts:
-      • SKE encryption: The computation complexity of encrypting $10,1,\cdots,n-$ 1 is $O(n)$.
-      • OKVS: The computation complexities of Encode and Decode algorithms are both $O(n)$.
-The size of OKVS is $1.28\kappa n$ bits, where $\kappa$ is the size of a SKE ciphertext with the same range as before.
-      • ssVODM: The ssVODM protocol requires a total of $(T+l-\log n-1)n$ AND gates, where $T$ is the number of AND gates in the SKE decryption circuit.
-      For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke mq-ssPMT of size $B$. Overall, each party $P_{j}$ engages in $(m-1)$ instances of mq-ssPMT.
-      • Offline phase.
-The computation complexity of each party is $O((T+l-\log n)mn\log((T+l-\log n)n))$.
-The communication complexity of each party is $O(t\lambda m\log(((T+l-\log n)n)/t))$.
-The round complexity is $O(1)$.
-      • Online phase.
-The computation complexity of each party is $O((T+l-\log n)mn)$.
-The communication complexity of each party is $O((T+l+\kappa-\log n)mn)$.
-The round complexity is $O(\log(l-\log n))$.
-    • Random oblivious transfers.
-For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke ROT extension of size $B$.
-Overall, each party $P_{j}$ engages in $(m-1)$ independent instances of ROT extension of size $B$.
-      • Offline phase. The computation complexity of each party is $O(mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
-      • Online phase. The computation complexity of each party is $O(mn)$. The communication complexity of each party is $O(mn)$. The round complexity is $O(1)$.
-      The costs of the remaining two steps, multi-party secret-shared shuffle and output reconstruction, are exactly the same as those in our SK-MPSU protocol.(cf. Appendix  0.F.2).
-    • Total costs.
-      • Offline phase.
-The computation complexity of each party is $O((T+l-\log n)mn\log((T+l-\log n)n)+m^{2}n\log mn)$.
-The communication complexity of each party is $O(t\lambda m\log(((T+l-\log n)n)/t)+\lambda m^{2}n(\log m+\log n))$.
-The round complexity is $O(1)$.
-      • Online phase.
-The computation complexity of $P_{1}$ is $O((T+l-\log n)mn+m^{2}n)$.
-The communication complexity of $P_{1}$ is $O((T+l+\kappa-\log n)mn+(l+\kappa)m^{2}n)$.
-The computation complexity of $P_{j}$ is $O((T+l-\log n)mn)$.
-The communication complexity of $P_{j}$ is $O((T+l+\kappa-\log n)mn)$.
-The round complexity is $O(\log(l-\log n)+m)$.
-    Appendix 0.G Multi-Party Private-ID
-    We generalize the two-party private-ID [BKM+20] to multi-party setting.
+  • 0.E.4.2 Comparison with [GNT23]
+    In Table 4, we present a comparison of the theoretical computation and communication complexity for each party in both offline and online phases between [GNT23] and our PK-MPSU.
+    We conclude that the complexity of our PK-MPSU surpasses theirs in all respects, including the computation and communication complexity of the leader and clients in the offline and online phases, as depicted in the Table 4.
+    Appendix 0.F Multi-Party Private-ID
+    It is natural to generalize the two-party private-ID [BKM+20] to multi-party setting.
 Suppose there are $m$ parties, each possessing a set of $n$ elements.
-The multi-party private-ID functionality assigns a unique random identifier to each element across all input sets, ensuring that identical elements in different sets obtain the same identifier.
-Each party receives identifiers associated with its own input set, as well as identifiers associated with the union of all parties’ input sets.
+The multi-party private-ID functionality assigns a unique random identifier to each element across all input sets, ensuring that identical elements in different sets obtain the same identifiers.
+Each party receives identifiers associated with its own input elements, as well as identifiers associated with the union of all parties’ input sets.
 With multi-party private-ID, the parties can sort their private sets based on a global set of identifiers and perform desired private computations item by item, ensuring alignment of identical elements across their sets.
@@
 Hash function $\mathsf{H}(x):\{0,1\}^{l}\to\mathbb{G}$.
-    • Inputs. Each party $P_{i}$ has input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$.
+    • Inputs. Each $P_{i}$ has input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$.
     • Protocol.
@@
       2. 2.
-        For $1\leq i\leq m$, $P_{i}$ receives $\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}}$ for $j\in[n]$, then it computes $(\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}})^{-a_{i}k_{i}}=\mathsf{H}(x_{i}^{j})^{k_{1}\cdots k_{m}}$.
-We denote the set $\{\mathsf{H}(x_{i}^{1})^{k_{1}\cdots k_{m}},\mathsf{H}(x_{i}^{n})^{k_{1}\cdots k_{m}}\}$ as $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$, where each $r_{i}^{j}\in\mathbb{G}$.
+        For $1\leq i\leq m$, $P_{i}$ receives $\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}}$, $j\in[n]$, and computes $(\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}})^{-a_{i}k_{i}}=\mathsf{H}(x_{i}^{j})^{k_{1}\cdots k_{m}}$.
+We denote $\{\mathsf{H}(x_{i}^{1})^{k_{1}\cdots k_{m}},\mathsf{H}(x_{i}^{n})^{k_{1}\cdots k_{m}}\}$ as $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$, where each $r_{i}^{j}\in\mathbb{G}$.
       3. 3.
-        The parties invoke $\Pi_{\mathsf{PK\text{-}MPSU}}$ where $P_{i}$ inputs $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$.
+        The parties invoke $\mathcal{F}_{\mathsf{mpsu}}$ where $P_{i}$ inputs $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$.
 $P_{1}$ receives the union $R^{*}=\bigcup_{i=1}^{m}Y_{i}$, and sends it to other parties.
@@
     Figure 14: DH-based Multi-Party Private ID $\Pi_{\mathsf{MPID}}$
+    Theorem 0.F.1
+    Protocol $\Pi_{\mathsf{MPID}}$ securely implements $\mathcal{F}_{\mathsf{MPID}}$ against any semi-honest adversary corrupting $t<m$ parties in the $\mathcal{F}_{\mathsf{mpsu}}$-hybrid model, assuming the DDH assumption and $\mathsf{H}$ is a random oracle.
     • Correctness.
-The first two steps essentially realize a multi-party distributed OPRF protocol, where each party $P_{i}$ inputs a set $\{x_{i}^{1},\cdots,x_{i}^{n}\}$ and receives its own PRF key $k_{i}$ and the PRF values computed on its input set using all parties’ keys $k_{1},\cdots,k_{m}$, denoted as $\{\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{1}),\cdots,\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{n})\}$.
-In this case, $\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x)=\mathsf{H}(x)^{k_{1}\cdots k_{m}}$.
-Note that even if $m-1$ parties collude, there exist one exponent private to the adversary, ensuring that the PRF values remain pseudorandom (The proof is analogous to that of the DH-based OPRF, so we omit it here).
-    • Security. Assuming $\mathsf{H}$ is a random oracle, the security of protocol follows immediately from the DDH assumption and the security of our PK-MPSU protocol.
+The first two steps essentially realize a multi-party distributed OPRF protocol, where each party $P_{i}$ inputs a set $\{x_{i}^{1},\cdots,x_{i}^{n}\}$ and receives its own PRF key $k_{i}$ along with the PRF values computed on its input set using all parties’ keys $k_{1},\cdots,k_{m}$, denoted as $\{\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{1}),\cdots,\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{n})\}$.
+In Figure 14, each party $P_{i}$ first masks the hash value of each element by raising it to the power of $a_{i}$.
+Then these masked values are then cycled among the parties, with each party raising them to the power of its key.
+Finally, $P_{i}$ unmasks the values and raises them to the power of its won key $k_{i}$, resulting in the distributed PRF values $\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x)=\mathsf{H}(x)^{k_{1}\cdots k_{m}}$.
+Even if $m-1$ parties collude, there still exists one exponent private to the adversary, ensuring that the PRF values remain pseudorandom according to the DDH assumption.
+      In the step 3, the parties invoke an MPSU protocol on their distributed PRF values, so than they can securely receive all PRF values on the union of input sets.
+    • Security. Assuming $\mathsf{H}$ is a random oracle, the security of protocol follows immediately from the DDH assumption and the security of our PK-MPSU.
+      Modern DH-based cryptosystems often use elliptic curves for the underlying cyclic group, due to their compact size and better efficiency at a given level of security.
+In this context, the first two steps of the protocol have relatively low communication costs, meanwhile, the element space of the MPU protocol in step 3 is EC points, which aligns perfectly with the element space of our PK-MPSU.
+This alignment allows us to instantiate the MPSU functionality using our PK-MPSU, leveraging its lowest communication costs.
+As a result, the entire multi-party private-ID protocol achieves extremely low communication costs, making it well-suited for deployment in bandwidth-restricted environments.
+      Another benefit of instantiating the MPSU functionality with our PK-MPSU is to achieve linear computation and communication complexity.
+Since the multi-party distributed OPRF phase has linear computation and communication, by integrating our PK-MPSU, the entire multi-party private-ID protocol can maintain the linear complexities throughout.